### v7

Abstract
Data from national accounts show no effect of change in net saving or consumption, in ratio to market-value capital, on change in growth rate of market-value capital (capital acceleration). Thus it appears that capital growth and acceleration arrive without help from net saving or consumption restraint. We explore ways in which this is possible, and discuss implications for economic teaching and public policy.
keywords:
[…]
1 Introduction and overview
Many economists over the centuries have reasoned that net saving, or equivalently net investment [^1] As reported in national accounts; they differ only by statistical discrepancy. Also see Appendix A., should tend to give equal capital growth. Economists since the early nineteenth century have added the proviso that net saving cannot safely outpace innovation; more capital must mean capital redesigned for greater productivity if economies are to escape risk of capital glut and diminishing returns (West (1815), Ricardo (1815), Malthus (1815)). Roy Harrod (1939) described that limit for safe net saving, meaning the rate of imagining and developing new ideas for more productive forms of capital, as the “warranted rate”. Harrod, and many other economists of his time and since, have focused on growth of output rather than of capital, but have modeled growth of output by first assuming the equivalence of net saving and capital growth, within the warranted rate, and then looking for effects of that capital growth on later output growth.
Some other economists, including John Rae (1834) and John Stuart Mill (1848) (see Appendix C), argued that capital growth might also be explained by a rise in productivity of capital and labor already extant. Ways might found for existing factors to produce more, that is, and so to allow more consumption, or more capital growth, or any mix of the two, without inputs of net saving. Robert Solow (1957) allowed that possibility for “disembodied” growth, where plant and products already existing are repurposed or redeployed in more productive ways.
We test between those two explanations of capital growth, by net saving or by increase in productivity of capital and labor already in existence, by comparing net saving to concurrent change in market-value capital in 86 countries. As changes in net saving are expected to be associated with opposite changes in consumption, we also compare change in consumption to concurrent change in capital growth (capital acceleration). All data are drawn from national accounts of those countries as collated on the free website World Inequality Database.
Those test results from data for net saving/capital ratio and for consumption/capital ratio are easiest to reconcile in an interpretation that change in neither affects change in capital growth rate. [^2] Another interpretation allows that net saving may account for 6.4% of capital growth. See Section 4 below. This finding would support the views of Rae and Mill, and of Solow as to disembodied growth. It would suggest that capital growth, even in acceleration, arrives without help from net saving or consumption restraint. Net saving, if so, raises the physical quantity of capital, but not the aggregate value, and so reduces the value per unit.
The data are most easily explained by the present value principle, and by production efficiencies enabled through innovation. Value is created in the mind of the market at the moment when prospective cash flows are discounted. It is created only if the market sees a path, step by step, from the start, to practical realization of those prospective cash flows. Then capital growth arrives when the market first evaluates prospective cash flows, and is realized eventually in physical outcomes insofar as the market has predicted correctly. That is to say that capital growth, in the mind of the market, precedes and pays for any needed investment in new plant and equipment.
Meanwhile the innovator acquires materials and plant capacity and labor skills at market prices determined by their uses in current technology, but applies them more productively until competition catches up. It is that temporary market advantage to the innovator which explains capital growth without net saving in a practical and mechanical sense, while the present value principle gives the explanation in terms of market valuation. This idea will be called “free growth theory” for easy reference.
It predicts only at the largest scales, and only for the private sector. Individuals and groups and even small economies can grow through investment from outside. That possibility is foreclosed only at the scale of all capital and all economies together. The public sector, meanwhile, responds to political rather than market choices, and grows or shrinks accordingly.
If free growth theory is right, tax policy and other policy to encourage saving over consumption should be reviewed. These policies include the higher tax on ordinary income than on capital gains, and the double tax on corporate dividends.
Inferences for economic teaching include the obvious ones for growth theory and for net saving in general. They include others as well. One of the central doctrines of the marginalist revolution has held that market realization converges to producer cost, when that cost includes imputed interest on assets owned. Net saving gives producer cost, and falls short of market realization in the presence of technological growth from new ideas. Meanwhile the doctrine that net income equals consumption plus net saving is put into question by evidence offered here suggesting that net saving increases the physical quantity of capital, but not the aggregate value. In general, economics might consider relying less on book value and net investment, and more on market value and on the power of ideas.
2 Net saving, capital growth and capital acceleration
This study compares net saving $S_{net}$ to concurrent growth in market-value capital $K$ from data in national accounts. Capital growth ${\Delta K}_{i}$ in each year $i$ for each reporting country is found as $K_{i}-K_{i-1},$ and compared to $S_{net,i}$ reported for that year and country. As we will be testing for differences between $\Delta K$ and $S_{net,i}$, we begin by writing
$\Delta K_{i}=z_{i}S_{net,i}+Q_{i}\ .$ (1)
Here "net saving efficiency" $z$ is the fraction of $S_{net}$ actually realized in growth of capital value $K$. $Q$ means the sum of market noise, which may prove positive or negative or zero, plus any part of capital growth explained by concurrent productivity gain as described by Rae and Mill, and by Solow as to disembodied growth. We call this sum of noise and productivity gain “free growth” in that it costs no net saving.
The object of testing is to find effects of $S_{net}$ on concurrent $\Delta K$, and so to help evaluate historic and current teachings as to those effects. We submitted above that most teaching, with exceptions noted as to Rae, Mill, and Solow, and within the warranted rate, predicts net saving to differ from concurrent growth in market-value capital only by market noise which tends to balance out over scale and time. Free growth $Q$ in Eq. (1), in that case, will give the market noise converging to zero. That consensus prediction, which we will challenge, will here be called “thrift theory”; $S_{net}$, in thrift theory, is expected to converge to $\Delta K$ if held within the warranted rate. That is,
$E(z)=1\ ,\quad E(\Delta K)=S_{net}\quad\text{and}\quad E(Q)=0,\quad\text{if}\quad\frac{\Sigma S_{net}}{\Sigma K}\leq u\ ,\quad\text{in thrift theory,}$, (2)
where:
  1. 1.
    $\Sigma S_{net}$ is collective net saving over the economy
  2. 2.
    $\Sigma K$ is collective capital over the economy before current $\Sigma S_{net},$ and
  3. 3.
    $u$ is the warranted rate.
Free growth theory makes the opposite prediction that net saving adds no capital value, and that capital growth, at the largest scales in the private sector, is explained wholly by increased productivity of skills and materials already present. Thus, referring to Eq. (1), it predicts:
$E(z)=0\quad\text{and}\quad E(\Delta K)=Q\quad\text{in free growth theory.}$. (3)
$E(z)$, $E(\Delta K)$ and $E(Q)$ here give the expected values of $z$, $\Delta K$ and $Q$ respectively. Expected value means predicted average of outcomes over all observations. We will test those predictions against data for saving and capital growth taken from national accounts. As secular economic growth has tended to make later stocks and flows larger than earlier ones, we first divide by $K$ (normalize) to avoid overweighting of more recent years in finding that average. Division of Eq. (1) by $K_{i-1}$ gives
$\frac{(\Delta K)_{i}}{K_{i-1}}=\frac{z_{i}S_{net,i}}{K_{i-1}}+\frac{Q_{i}}{K_{i-1}}\ .$ (4)
The first term in Eq. (4) gives capital growth rate $g(K)$. The fraction in the second term is a variant of the Keynesian net saving rate $s_{net}$ where capital rather than output becomes the denominator. This flow will show as $s^{*}$, with the subscript “net” left implicit, and with the understanding that the denominator shows market-value capital rather than output. The third term in Eq. (4) will be called free growth rate and shown as $q_{s}$. That is,
$g{(K)}_{i}=\frac{(\Delta K)_{i}}{K_{i-1}}\ ,\quad s^{*}_{i}=\frac{S_{net,i}}{K_{i-1}}\ ,\quad\text{and}\quad q_{i}=\frac{Q_{i}}{K_{i-1}}\ ,$
so that Eq. (4) can be shown more compactly as
$g(K)_{i}=z_{i}s^{*}_{i}+q_{i}\ .$ (5)
By the definition $q=\frac{Q}{K}$, an expected value $E(Q)=0$ implies $E(q)=0$. Application of Eq. (2) to Eq. (5) now gives
$E(g(K))=s^{*}\quad\text{and}\quad E(q)\ ,\quad\text{under thrift assumptions,}$, (6)
where “thrift assumptions” are that thrift theory is correct and that the warranted rate is not exceeded. Eqs. (2) and (4), meanwhile, give
$E(g(K))=q\quad\text{in free growth theory.}$. (7)
Note here that free growth theory makes no prediction as to the quantity $s^{*}_{i}$. Rather it predicts $E(z)=0$ in Eq. (5), so that $s^{*}_{i}$ takes no part in calculation of $\Delta g(K)_{i}$.
The prediction in Eq. (6) can be tested directly from data for market-value capital and net saving in national accounts. We define the "auxiliary thrift index" by
θ=sg(K),that is, and test the prediction
[…]
$E(\theta^{*})=1\quad\text{under thrift assumptions.}$.
Results show in Fig. 1 and Tables 1 and 2. They show $\theta^{*}$ as averaging something near $\frac{1}{2}$, and with wide variance from country to country and from year to year. Regression of $s^{*}$ on $g(K)$ is found at 0.2103. Free growth theory, again, makes no prediction as to relationships between net saving and capital growth.
The concepts of correlation, covariance and regression are founded on the principle that causal relationships among variables tend to be revealed in simultaneous changes. Thus the two theories are best tested in their implications for changes in capital growth, which we call capital acceleration and show as $\Delta g(K)$. First note that Eq. (5) allows
$(\Delta g(K))_{i}=(\Delta(zs^{*}))_{i}+(\Delta g)_{i}\ ,\quad\text{where}$ (8)
$(\Delta g{(K)})_{i}=(g{(K)})_{i}-(g(K))_{i-1}\ ,\quad(\Delta(zs^{*}))_{i}=z_{i}s^{*}_{i}-z_{i-1}s^{*}_{i-1}\ ,\quad\text{and}\quad(\Delta q)_{i}=q_{i}-q_{i-1}\ .$
For any variables $a$ and $b$, we may reason $E(a-b)=E(a)-E(b)$. By this and by Eqs. (7) and (8), then,
$E(g(K))=\Delta s^{*}\quad\text{and}\quad E(\Delta q)=0\ ,\quad\text{under thrift assumptions.}$. (9)
Division of Eq. (8) by capital acceleration, and rearrangement, gives
$\frac{(\Delta(zs^{*}))_{i}}{(\Delta g{(K)})_{i}}+\frac{(\Delta q)_{i}}{(\Delta g{(K)})_{i}}=1\ .$ (10)
Fig. 1: $\theta^{*}$ (86 countries)
Table 1: Regression of $s^{*}$ on $g(K)$ (Screen = 0.01). $H_{0}$ per thrift theory: $g(K)\cong s^{*}$ $\overline{\theta^{*}}\cong 1$.
$s^{*}$
Regression of $s^{*}$ on $g(K)$ 0.1857∗∗∗
Observations 1,826
R2 0.77165
Within R2 0.13962
Year fixed effects $\checkmark$
Country fixed effects $\checkmark$
Define $\theta_{s,i}=\frac{(\Delta zs^{*})_{i}}{(\Delta g{(K)})_{i}}$ and $\varphi_{s,i}=\frac{(\Delta q)_{i}}{(\Delta g{(K)})_{i}}$ to restate Eq. (10) as
$\theta_{s,i}+\varphi_{s,i}=1\ .$ (11)
Next define $\theta_{s}=E(\theta_{s,i})$ and $\varphi_{s}=E(\varphi_{s,i})$. By Eq. (11), then,
[…]
$\theta_{s}=1\quad\text{and}\quad\varphi_{s}=0\ ,\quad\text{under thrift assumptions.}$. (13)
Eq. (11) enables measurement of effects of thrift and free growth in a given period $i$, that is, while Eq. (12) does the same for all periods together.
$\theta$ and $\varphi$ will be called the “thrift index” and “free growth index” respectively. $\theta_{s}$ and $\varphi_{s}$ give their values as found from data for change in net saving. Expected values, again, are predicted averages of outcomes. Thus Eq. (13) and thrift theory can be tested by finding average values of $\theta_{s,i}$ and comparing findings to the expected value $\theta=1$. First we find
$\overline{\theta_{s,i}}=\frac{1}{m}\sum^{m}\theta_{s,i}\quad\text{and}\quad\overline{\varphi_{s,i}}=\frac{1}{m}\sum^{m}\varphi_{s,i}\ ,$
Table 2: Average $\theta^{*}$ in 86 countries (screen = 0.01). Number of years clearing screen shown in ()
Country Period $\overline{\theta^{*}}$ Armenia 1996 - 2020 (23) -0.63 Aruba 1996 - 2001 (6) 1.25 Australia 1980 - 2018 (33) 0.24 Austria 1996 - 2021 (23) 0.83 Azerbaijan 1996 - 2020 (23) 1.57 Bahrain 2009 - 2013 (5) -2.43 Belgium 1996 - 2021 (21) 0.61 Bolivia 1998 - 2015 (17) 0.40 Botswana 1996 - 1999 (4) 3.54 Brazil 1996 - 2019 (21) 0.25 British Virgin Islands 1996 - 1999 (4) 2.45 Bulgaria 1996 - 2016 (16) 0.14 Burkina Faso 2000 - 2019 (20) 0.23 Cameroon 1996 - 2019 (24) 1.10 Canada 1980 - 2022 (38) 0.37 Cape Verde 2008 - 2017 (9) 0.53 Chile 1997 - 2021 (19) 0.84 China 1992 - 2020 (29) 0.81 Colombia 1996 - 2022 (27) 0.60 Costa Rica 2013 - 2020 (6) 0.21 Croatia 1996 - 2021 (24) 0.43 Curaçao 2001 - 2016 (14) 1.26 Cyprus 1996 - 2021 (25) 0.07 Czechia 1994 - 2020 (20) 0.17 Côte d’Ivoire 1996 - 2000 (5) 0.29 Denmark 1996 - 2022 (26) 0.22 Dominican Republic 1996 - 2016 (10) 1.52 Ecuador 2008 - 2020 (13) 0.53 Egypt 1997 - 2015 (19) 0.92 El Salvador 2015 - 2019 (5) -0.62 Estonia 1996 - 2021 (24) 0.45 Finland 1996 - 2021 (23) 0.17 France 1980 - 2021 (35) 0.26 Germany 1980 - 2022 (38) 0.70 Greece 1996 - 2021 (26) 0.17 Guatemala 2002 - 2021 (20) -0.72 Guinea 2004 - 2010 (5) 0.72 Honduras 2001 - 2015 (14) -0.57 Hong Kong 1997 - 2011 (15) 0.90 Hungary 1996 - 2021 (21) 0.14 Iceland 2001 - 2014 (14) 0.22 India 1999 - 2019 (21) 0.57 Indonesia 2017 - 2019 (3) 0.94 Iran 1997 - 2017 (20) 4.96 Ireland 1996 - 2021 (22) 0.26 Israel 1996 - 2019 (24) 0.46 Country Period $\overline{\theta^{*}}$ Italy 1980 - 2022 (36) 0.13 Japan 1980 - 2021 (29) 0.19 Kazakhstan 1997 - 2022 (25) 0.60 Kuwait 2003 - 2017 (14) 1.11 Kyrgyzstan 1996 - 2021 (24) 0.46 Latvia 1996 - 2021 (24) -0.23 Lithuania 1996 - 2021 (24) 0.13 Luxembourg 1996 - 2021 (22) 0.03 Malaysia 2007 - 2015 (9) 2.52 Malta 1996 - 2021 (23) 0.16 Mauritius 2014 - 2018 (5) 0.04 Mexico 1996 - 2021 (22) 0.15 Moldova 1996 - 2019 (22) -0.91 Mongolia 2006 - 2020 (15) 0.10 Morocco 1999 - 2021 (23) 1.13 Netherlands 1996 - 2021 (23) 0.51 New Zealand 1996 - 2019 (21) 0.78 Nicaragua 2007 - 2018 (12) 0.08 Niger 1996 - 2019 (24) 0.90 Norway 1981 - 2021 (38) 0.95 Peru 2008 - 2021 (14) 0.76 Philippines 1996 - 2022 (27) 1.76 Poland 1996 - 2021 (25) 0.71 Portugal 1996 - 2022 (24) 0.00 Qatar 2002 - 2017 (14) 1.89 Romania 1996 - 2020 (22) 0.03 Russia 1996 - 2017 (13) -0.19 Saudi Arabia 2003 - 2009 (7) 2.65 Senegal 2015 - 2021 (7) 0.24 Serbia 1998 - 2021 (23) 0.61 Slovakia 1996 - 2022 (24) 0.15 Slovenia 1996 - 2021 (20) 0.28 South Africa 1996 - 2022 (24) 0.28 South Korea 1996 - 2020 (24) 0.63 Spain 1995 - 2021 (25) 0.32 Sweden 1980 - 2021 (37) 0.41 Switzerland 1993 - 2021 (25) 0.69 Tunisia 1996 - 2011 (16) 0.29 Turkey 2010 - 2017 (8) 1.32 USA 1980 - 2021 (38) 0.31 Ukraine 1996 - 2019 (21) 0.17 United Kingdom 1981 - 2021 (36) 0.16 Uruguay 2016 - 2017 (2) 1.16 Uzbekistan 2011 - 2021 (10) 0.21 Vanuatu 2002 - 2007 (6) 0.29 Venezuela 1998 - 2019 (20) 0.55
Note: Thrift theory predicts $\overline{\theta^{*}}\cong 1$. Free growth theory makes no prediction for these data.
where $m$ is the number of observed values of $\theta_{s,i}$ and $\varphi_{s,i}$, and test the predictions
$\overline{\theta_{s,i}}\cong 1\quad\text{and}\quad\overline{\varphi_{s,i}}\cong 0\ ,\quad\text{under thrift assumptions.}$.
Calculations of $\overline{\theta_{s,i}}$ and $\overline{\varphi_{s,i}}$ are not expected to show 1 and 0 exactly, under thrift assumptions, so that $\overline{\theta_{s,i}}$ and $\overline{\varphi_{s,i}}$ are not exactly equal to $\theta_{s}$ and $\varphi_{s}$, because the number of samples $m$ is finite.
Free growth theory, maenwhile, referring to Eq. (13), does not predict the exact opposite findings $\theta_{s}=0$ and $\varphi_{s}=1$. It rather expects a value of $\varphi_{s}$ somewhat less than unity for reasons shown in Section 4.
Fig. 2 shows average values of $\theta_{s,i}$ and $\varphi_{s,i}$ for 86 countries, both unweighted and weighted to GDP, over the period 1980-2022. Table 3 shows associated regressions of $\Delta s^{*}$ and $\Delta g_{s}$ respectively on $\Delta g(K)$ at 0.1000 and 0.9000. Average values of $\varphi_{s}$ for the 86 individual countries show in Table 4. To control distortions brought by small absolute denominators, years were screened out where $|\Delta g(K)|$ was found at less than 0.01 (see Section 13). Results show $\overline{\varphi_{s,i}}=0.936$ and $\overline{\theta_{s,i}}=0.064$ for data weighted to GDP. These findings largely support free growth theory, but could suggest a small role (6.4%) for thrift as well. This possibility will be discussed in Section 4. We will first test thrift theory from a different approach.
Fig. 2: Thrift and free growth indexes derived from net saving (86 countries).
Table 3: Regression of $\Delta s^{*}$ and $\Delta q_{s}$ on $\Delta g(K)$ (Screen = 0.01). $H_{0}\ \text{per thrift theory:}\ \Delta g(K)\cong\Delta s^{*}\ \&\ \Delta q_{s}\cong 0$
$\Delta s^{*}$ $\Delta q_{s}$
Regression of value shown on $\Delta g(K)$ 0.1000∗∗∗ 0.9000∗∗∗
(0.0145) (0.0145)
Observations 1,574 1,574
R2 0.32867 0.96233
Within R2 0.22713 0.95965
Year fixed effects $\checkmark$ $\checkmark$
Country fixed effects $\checkmark$ $\checkmark$
3 Consumption and capital growth
The concept of cash flow is little used in economics, aside from finance and in present value calculations. It is defined implicitly in the equation
$P=\Delta K+F\ ,$ (14)
where $P$ and $F$ mean net profit and cash flow respectively [^3] We have found no standard notation for cash flow.. Cash flow means the net outflow of value, in cash or kind, from an asset to its owner for consumption $C$ or for reinvestment in other assets. This flow of spending is augmented by pay $\Pi$ [^4] Compensation of employees plus a share of mixed income. We avoid the standard notation $W$, which is also used for work, as we prefer to distinguish pay from work.. At the scale of the closed economy, meaning of all assets together, reinvestment cancels to zero. At this scale, then, the sum of spending $F+\Pi$ is spent on consumption alone. That is,
$\Pi+F=C$ (15)
at the collective scale. Eq. (14), meanwhile, may be arranged as $F=P-\Delta K$. Substitution of this in Eq. (15), with rearrangement, gives
$\Delta K=\Pi+P-C\ .$ (16)
Thrift theory, which expects $\Delta K=S_{net}$, writes this as
$S_{net}=\Pi+P-C\ ,\quad\text{under thrift assumptions.}$. (17)
Pay, net profit [^5] Jordà et al. (2019), The rate of return on everything. and consumption can all be measured, to an extent, as can market-value capital, so that this prediction of thrift theory could be measured directly. Comparison between Eqs. (16) and (17), however, shows that such a test, given perfect data in national accounts, would only repeat that of Eq. (6). As before, the two theories are tested most revealingly in capital acceleration. Given the definitions [^6] The notation $c^{*}$, as with $s^{*}$, shows an asterisk to distinguish its meaning $c^{*}=\frac{C}{K}$ from the Keynesean $c=\frac{C}{Y}$
$\pi_{i}=\frac{\Pi_{i}}{K_{i-1}}\ ,\quad p_{i}=\frac{P_{i}}{K_{i-1}}\quad\text{and}\quad c^{*}_{i}=\frac{C_{i}}{K_{i-1}}\ ,$
Eq. (16) can be divided by capital to reach
$g(K)_{i}=\pi_{i}+p_{i}-c^{*}_{i}\ ,\quad\text{implying}$ (18)
$(\Delta g(K))_{i}=\Delta\pi_{i}+\Delta p_{i}-\Delta c^{*}_{i}\ ,\quad\text{where}$ (19)
$\Delta\pi_{i}=\pi_{i}-\pi_{i-1}\ ,\quad\Delta p_{i}=p_{i}-p_{i-1}\quad\text{and}\quad\Delta c^{*}_{i}=c^{*}_{i}-c^{*}_{i-1}\ ,$
as before. Thrift theory expects $\Delta g(K)=\Delta s^{*}=\Delta\pi+\Delta p-\Delta c$ by Eqs. (9) and (17). It further expects changes in productivity, as realized in the pay/capital and profit/capital ratios $\pi$ and $p$, to converge to zero as market noise, so that the only source of net saving is in consumption restraint $-\Delta c^{*}$. Thus
$E(\Delta g(k))=-\Delta c^{*}\quad\text{and}\quad E(\Delta\pi)=E(\Delta p)=0\ ,\quad\text{under thrift assumptions.}$. (20)
Free growth theory, conversely, predicts that net saving contributes nothing to capital growth, and hence that net saving through consumption restraint, at the collective scale and in the private sector, is sacrifice with nothing in return. Free growth theory expects rather that costless increases in the ratios of pay and net profit to capital, enabled by greater productivity of human and physical capital already present, account for all capital growth and acceleration. Referring to Eq. (19), however, it predicts
[…]
for reasons shown in Section 4. As in Eq. (10), we prepare to test thrift theory by first dividing Eq. (20) by capital acceleration to reach
$E(\frac{-\Delta c^{*}}{\Delta g(K)})=1\ ,\quad\text{under thrift assumptions.}$. (22)
Define $\theta_{c,i}=\frac{-\Delta c^{*}_{i}}{\Delta g(K)_{i}}$ and $\varphi_{c,i}=1-\theta_{c,i}$ to re-express Eq. (22) as
$\theta_{c,i}+\varphi_{c,i}=1\ .$ (23)
By Eqs. (20) through (23), further,
$E\left(\theta_{c}\right)=1\quad\text{and}\quad E\left(\varphi_{c}\right)=0\ ,\quad\text{under thrift assumptions.}$. (24)
Define $\theta_{c}=E\left(\theta_{c}\right)$ and $\varphi_{c}=E\left(\varphi_{c}\right)$ to re-express Eqs. (23) and (24) respectively as
$\theta_{c}+\varphi_{c}=1\ ,\quad\text{and}$ (25)
$\theta_{c}=1\quad\text{and}\quad\varphi_{c}=0\ ,\quad\text{under thrift assumptions.}$. (26)
We infer $\theta_{c}\cong\overline{\theta_{c,i}}$ and $\varphi_{c}\cong\overline{\varphi_{c,i}}$ as before, and test thrift theory by comparing average yearly values of $\theta_{c,i}$ and $\varphi_{c,i}$ to its predictions $\overline{\theta_{c,i}}\cong 1$ and $\overline{\varphi_{c,i}}\cong 0$.
[…]
Fig. 3: Thrift and free growth indexes derived from consumption (86 countries).
Table 4: Average $\varphi_{s,i}$ and $\varphi_{c,i}$ in 86 countries (screen = 0.01). Number of years clearing screen shown in ()
Country Period $\overline{\varphi_{s,i}}$ $\overline{\varphi_{c,i}}$ Armenia 1997 - 2020 (20) 0.88 1.18 Aruba 1997 - 2001 (5) -0.22 8.44 Australia 1974 - 2019 (35) 0.97 1.08 Austria 1997 - 2021 (16) 0.90 1.13 Azerbaijan 1997 - 2020 (23) 0.25 2.20 Bahrain 2010 - 2013 (4) 0.59 1.32 Belgium 1997 - 2021 (19) 0.95 1.06 Bolivia 1999 - 2014 (12) 0.96 1.12 Botswana 1997 - 2000 (4) 0.76 1.27 Brazil 1998 - 2020 (22) 0.90 1.24 British Virgin Islands 1997 - 1998 (2) -1.24 -2.20 Bulgaria 1997 - 2017 (15) 1.03 1.19 Burkina Faso 2001 - 2019 (16) 0.96 1.28 Cameroon 1997 - 2019 (19) 0.82 2.71 Canada 1974 - 2022 (41) 0.91 1.08 Cape Verde 2009 - 2016 (8) 0.73 2.02 Chile 1998 - 2021 (22) 0.94 1.48 China 1993 - 2020 (23) 0.91 1.05 Colombia 1997 - 2022 (18) 1.03 2.39 Costa Rica 2014 - 2020 (6) 0.96 1.25 Croatia 1997 - 2021 (22) 0.81 2.09 Curaçao 2002 - 2016 (14) 0.59 0.51 Cyprus 1997 - 2021 (23) 1.06 1.16 Czechia 1995 - 2021 (19) 1.03 1.08 Côte d’Ivoire 1997 - 2000 (4) 1.03 1.05 Denmark 1997 - 2022 (22) 0.95 1.14 Dominican Republic 2007 - 2015 (9) 0.85 1.41 Ecuador 2009 - 2020 (12) 0.83 1.91 Egypt 1998 - 2015 (16) 0.87 2.14 El Salvador 2016 - 2019 (4) 0.02 -2.32 Estonia 1997 - 2021 (19) 0.94 1.11 Finland 1997 - 2022 (21) 1.02 1.20 France 1972 - 2021 (33) 0.91 1.16 Germany 1972 - 2022 (31) 0.95 1.14 Greece 1997 - 2021 (24) 0.76 1.47 Guatemala 2003 - 2021 (11) 0.85 1.32 Guinea 2005 - 2010 (6) 0.92 1.23 Honduras 2003 - 2015 (10) 0.60 0.85 Hong Kong 1997 - 2011 (13) 0.96 1.25 Hungary 1997 - 2021 (21) 0.93 1.10 Iceland 2002 - 2014 (12) 0.90 1.12 India 2000 - 2019 (14) 0.81 0.99 Indonesia 2019 - 2019 (1) 1.41 0.98 Iran 1998 - 2018 (19) -0.09 1.09 Ireland 1997 - 2021 (23) 0.98 1.23 Israel 1997 - 2017 (18) 0.93 1.72 Country Period $\overline{\varphi_{s,i}}$ $\overline{\varphi_{c,i}}$ Italy 1981 - 2021 (22) 0.94 1.12 Japan 1981 - 2021 (30) 0.95 1.13 Kazakhstan 1997 - 2022 (23) 0.92 1.16 Kuwait 2005 - 2017 (11) 0.96 1.13 Kyrgyzstan 1998 - 2021 (22) 0.77 1.29 Latvia 1998 - 2021 (20) 0.99 1.15 Lithuania 1997 - 2021 (18) 0.90 1.19 Luxembourg 1997 - 2021 (25) 1.00 1.02 Malaysia 2008 - 2015 (7) 0.99 1.19 Malta 1997 - 2021 (23) 0.92 1.04 Mauritius 2015 - 2018 (4) 0.98 0.99 Mexico 1997 - 2021 (22) 0.94 1.16 Moldova 1997 - 2019 (21) 0.89 1.27 Mongolia 2008 - 2021 (11) 0.90 1.06 Morocco 2000 - 2021 (17) 0.81 1.27 Netherlands 1997 - 2021 (20) 0.97 1.08 New Zealand 1997 - 2019 (18) 0.86 1.02 Nicaragua 2008 - 2019 (12) 0.94 1.13 Niger 1997 - 2019 (19) 0.78 1.87 Norway 1982 - 2021 (34) 0.88 1.08 Peru 2009 - 2021 (11) 0.96 1.30 Philippines 1997 - 2022 (20) 0.63 1.20 Poland 1997 - 2021 (15) 0.80 1.07 Portugal 1997 - 2022 (17) 0.98 1.08 Qatar 2004 - 2018 (12) 0.71 0.86 Romania 1997 - 2020 (14) 0.98 1.16 Russia 1997 - 2018 (14) 0.91 1.11 Saudi Arabia 2005 - 2009 (3) -1.23 0.02 Senegal 2018 - 2021 (3) 0.81 2.31 Serbia 1999 - 2021 (22) 0.78 1.67 Slovakia 1997 - 2021 (18) 0.91 1.18 Slovenia 1997 - 2021 (20) 0.90 1.07 South Africa 1997 - 2022 (21) 0.90 1.13 South Korea 1997 - 2020 (15) 0.96 1.05 Spain 1996 - 2021 (23) 0.96 1.11 Sweden 1974 - 2022 (41) 0.98 1.15 Switzerland 1993 - 2020 (24) 0.92 1.09 Tunisia 1997 - 2011 (12) 0.68 1.17 Turkey 2011 - 2015 (5) 1.16 1.54 USA 1972 - 2021 (44) 0.98 1.13 Ukraine 1997 - 2020 (23) 0.95 1.18 United Kingdom 1972 - 2021 (42) 0.98 1.21 Uruguay 2017 - 2017 (1) 0.44 0.98 Uzbekistan 2012 - 2021 (9) 0.95 1.14 Vanuatu 2003 - 2007 (3) 0.85 1.19 Venezuela 1999 - 2019 (20) 0.69 1.13
Note: Thrift theory predicts $\overline{\varphi_{s,i}}\cong\overline{\varphi_{c,i}}\cong 0$. Free growth theory predicts $\overline{\varphi_{s,i}}\cong\overline{\varphi_{c,i}}\cong 1$.
Table 5: Regression of $-\Delta c^{*}$ and $\Delta q_{c}$ on $\Delta g(K)$ (Screen = 0.01). $H_{0}\text{ per thrift theory: }\Delta g(K)\cong-\Delta c^{*}\&\Delta q_{c}\cong 0$
$-\Delta c^{*}$ $\Delta q_{c}$
Regression of value shown on $\Delta g(K)$0.3760∗∗∗ 1.376∗∗∗
(0.1055) (0.1055)
Observations 1,548 1,548
R2 0.53993 0.88294
Within R2 0.31853 0.86226
Year fixed effects $\checkmark$ $\checkmark$
Country fixed effects $\checkmark$ $\checkmark$
4 Differences in test results
Fig. 2 and Table 3 show $\theta_{s}$ to average 0.064, and regression of $\Delta s^{*}$ on $\Delta g(K)$ to average $0.100$, across all years and countries GDP-weighted. Fig. 3 and Table 5, meanwhile, showed these overall averages as $-0.142$ for $\theta_{c}$, and as $-0.376$ for regression of $-\Delta c^{*}$ on $\Delta g(K)$. These differences might be explained by systematic mismeasurements of $C$, $K$ and $S_{net}$ in national accounts. In that case, the evidence would leave open the possibility of some role for net saving (6.4%) in capital growth.
There is another possible interpretation. Regression or correlation or covariance among variables such as net saving and capital growth suggests causal relationship among them, but does not in itself reveal the direction of causality. Thrift theory accepts the natural inference that net saving is the cause, and capital growth the effect. Logic and experience, however, also suggest that adverse shocks such as war or depression might lower capital values despite all saving and investment, and might then motivate negative net saving (dissaving), meaning suspension of replacement saving, in order to maintain consumption. When peace and prosperity return, higher incomes will allow restoration of replacement saving to offset depreciation. Capital growth need not depend on positive net saving at any point, but rather allows an upward normalization of net saving from a negative value to zero. In this interpretation, change in capital growth becomes the cause, rather than the effect, of change in net saving.
This possibility could help explain differences between the measurements of GDP-weighted $\varphi_{s}(0.936)$ and $\varphi_{c}(1.142)$, without need to assume systematic mismeasurements of $S_{net}$, $K$ and $C$. Note here that the finding $\varphi>1$ is possible in principle when gain in productivity allows growth in capital and consumption at the same time (see the quote from Mill in Appendix C.) We welcome further study to resolve these uncertainties.
5 Mechanics of free growth
Some growth is capital widening, where structures and implements increase in number but do not change in design. Capital widening, however, is practical only so far before glut and diminishing returns set in. Further growth from that point must come from capital deepening, meaning improvements in the design of capital. Solow (1956) noted a kind of middle ground between capital widening and capital deepening in the disembodied growth mentioned earlier; ships carrying coals to Newcastle can raise prospective cash flows, and hence present value, by reversing the business plan. But Solow, who came to conclusions similar to ours from different evidence, puzzled as to how capital growth without net saving could be possible for capital deepening through “embodied” growth, where products of new design are made from plant of new design. [^7] The terms capital deepening, capital widening, embodied growth and disembodied growth are all Solow’s.
The solution, we suggest, is that embodied growth is disembodied growth on a finer scale. At each step toward realization of the new plant and products, raw materials and products and labor skills and plant capacity currently available on the market are adapted to new uses. The innovator pays for these inputs at a market price determined by their value in established productive uses, but applies them innovatively to realize higher prospective cash flows, and hence higher present values, to the innovator (Marshall (1890), Schumpeter and Opie (1934)). This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.888i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns. The present value of additional cash flow enabled by this advantage in return quantifies the innovator’s reserve and equivalently the non-random component of free growth. The innovator’s reserve quantifies the part of free growth explained by productivity gain as distinct from random market noise. As such, it is the quantity added to depreciation saving to enable embodied growth, so that net saving is never needed.
Our findings support those of Piketty and Zucman (2014) and Kurz (2023) as to the market power of innovators to explain capital growth beyond net saving. Again, we go farther by questioning the assumption that net saving contributes even a part of capital growth.
6 Optimum investment policy
Data and arguments adduced, with uncertainties discussed in Section 4, suggest that the optimum amount of saving, at the global scale, is replacement saving to offset depreciation, and nothing more. That would not mean book depreciation, as this study has stressed differences between book and market values. Up to a point, it should be possible to analyze the composition of market capital, and to model depreciation of the whole. A better plan, as Solow (1956) wrote in response to Harrod’s knife edge argument (Harrod, 1939), is to trust the market to maximize rate of return, and to sense the point where glut begins and returns fall. [^9] Harrod had argued that saving must hit the warranted rate exactly or risk positive feedback through the operation of the output/capital ratio (accelerator). Markets do so imperfectly when tax and other public policy reward saving over distributions and consumption. Findings in this paper suggest review of such policies. These include the double tax on dividends, and the greater tax rate on ordinary income than on capital gains. Effects of removing the double tax, and removing the difference between tax rates on ordinary income and on capital gains, could be revenue-neutral and non-partisan if the corporate tax were raised to match, if the tax rates on ordinary income and on capital gains met somewhere between, and if thoughtful grandfathering eased the transition.
7 Inferences for national accounts and macroeconomics
Eq. (16) was presented as a tautology or logical certainty. Cash flow $F$ from all physical capital $K$ together at the collective scale, that is, means outflow of value from $K$ net of any inflow into $K$ from any source, meaning from pay $\Pi$ as well as plowback of cash flow. Thus the sum $F+\Pi$ is spent on consumption $C$ alone by exhaust of alternatives. Meanwhile cash flow is net profit $P$ less capital growth $\Delta K$ [^10] Or equivalently gross profit less capital growth less depreciation. This reasoning gave $F+\Pi=C$, $F=P-\Delta K$, and consequently $\Pi+P=C+\Delta K$ (Eq. (16)).
Kuznets and the other founders of national accounts, under the influence of Keynes, seem to have reasoned alike. National accounts are double-entry $T$ accounts where net product $C+\Delta K$ is equated to net "factor payments" $\Pi+P$. National accounts, however, equate capital growth to net saving/investment, and effectively write Eq. (16) as $\Pi+P=C+S_{net}$. We have attempted to show that net saving does not tend to equate to capital growth, and indeed contributes no part of capital growth. If our argument proves correct, then national accounts, and the teachings of macroeconomics (macro) which they express, are constructed on a fundamental misperception. We would therefore welcome any critiques of our findings as a matter of academic and public interest.
8 Inferences for microeconomics
Microeconomics ("micro") studies the mutual effects of supply, demand and price. It teaches that producer cost, including interest as the cost of time, tends to equilibrate to market value realized. We have proposed an exception in the innovator’s reserve; the innovator pays the going rates for skills and materials needed, but applies them to make products of greater market value.
9 Inferences for inflation
If free growth theory is correct, then net investment adds a superfluous cost to capital growth, while decreasing market value per unit through capital glut. Both effects would tend to contribute directly to capital inflation, and indirectly to inflation of consumer goods by adding depreciation costs. This consideration adds to the case for urgent attention to our argument, and for more study to confirm or refute it.
10 Net output
The concept of net output is found in our discussions, but not in our equations. Net output means value added. Appendix B will show an argument that to equate value added to the sum of consumption and capital growth, and neither less nor more, would miss two components, one negative and one positive, in value added to human capital, and that both of these components are unmeasurable by known means. Since we work here from measurable data for consumption and net saving and market value capital, we did our best to reason from those three quantities alone.
11 Data sources
All our data are drawn from Distributional National Accounts (DINA) from the free online database World Inequality Database (WID). This source collates data from national accounts and tax data of 105 countries in constant currency units, and adjusts them where needed to conform to current standards of the System of National Accounts (SNA) published by the United Nations. We show results for the 86 of those countries which report all three of the factors, namely net saving, consumption and market-value capital, needed for deriving the thrift and free growth indexes. WID’s source for these data is national accounts.
Consumption $C$ in our text and equations is reproduced from the sum of Government Final Consumption Expenditure (mcongo) [^11] WID code and Private Expenditures of Households and NPISH (mconhn). This sums government expenditure GCE and personal consumption expenditure PCE. Net saving $S_{net}$ and market-value $K$ are taken from net national saving (msavin) and market-value Capital Wealth (mnweal) respectively. GDP, which we use only for weighting purposes in Figs. 1 and 2, is reproduced from GDP (mgdpro).
12 Accessing our results and methods
Tables and other displays of our findings for each country, and showing our methods of calculation, can be accessed at the web appendix (https://web-appendix.shinyapps.io/Sources_of_Capital_Growth).
13 Displays
Eqs. (12) and (25) show $\theta_{s}+\varphi_{s}=1$ and $\theta_{c}+\varphi_{c}=1$. Some displays here and in the web appendix save space by showing $\varphi_{s}$ and $\varphi_{c}$ only, leaving $\theta_{s}$ and $\theta_{c}$ implicit as their complements to unity. The web appendix includes displays of $\varphi_{s}$ and $\varphi_{c}$ for each year in each country over the report period. These tend to show upward and downward spikes in values of $\varphi_{s}$ and $\varphi_{c}$ in some years. Those spikes tend to be associated with small absolute values of denominators, in these cases $\Delta g(K)$, in those countries and years. Small denominators magnify errors in measurements of numerators. Worse, when denominators are small, small mismeasurements in them might reverse the term in sign.
To maximize reliability of test results, we apply a range of screens to omit years where absolute denominators fall below a given threshold. Some displays show data for all years, regardless of denominator size. Others screen out all years where absolute denominators are less then .01, then .025, then continuing upward in increments of .025 to a maximum screen of .15. $\varphi_{s}$ and $\varphi_{c}$ are plotted for each country unscreened and at each of the seven successive levels of screening. All displays applied a screen of .01. The denominator whose absolute value is screened is capital acceleration $\Delta g(K)$ or capital growth rate $g(K)$ in all displays.
Screening out years where absolute $\Delta g(K)$ or $g(K)$ is small would cost little in informative value even if measurements were exact. In those years, there is little capital acceleration or capital growth, positive or negative, for either thrift theory or free growth theory to explain. Market noise alone might account for $\Delta g(K)$ or $g(K)$ in such years. Screening reduces the number of observations, but increases the reliability and informative value of each.
14 Disclaimers
We accept that capital growth is impossible without capital replacement first, to make up losses to depreciation, and that the only source of capital replacement is replacement investment. Thus we accept the necessity and efficacy of replacement investment. We dispute only the efficacy of net investment in realizing capital growth after replacement investment has made up for depreciation.
15 Discussion and conclusions
Capital glut is the condition warned against by West, Ricardo, Malthus and Harrod. It is loosely defined as oversupply of capital at the current state of technology. We will not attempt a more exact definition here. Findings shown in our displays, anyhow, suggest that net saving raises the physical quantity of capital, say in number of shops, manufacturing plants or finished goods of similar design, without raising aggregate value of capital, and so contributes to capital glut.
These findings challenge the teachings that capital growth is effected by net saving enabled by consumption restraint, and that producer cost, including imputed interest as the opportunity cost of capital, converges to market realization. Evidence showing $\overline{\varphi_{s,i}}\cong 1$ and $\overline{\varphi_{c,i}}\cong 1$ suggests that all capital growth is free, and consequently that market realization, in the presence of innovation, exceeds producer cost by the entirety of capital growth. Meanwhile the same evidence, which indicates that net saving adds no capital value, suggests a review of the teachings that consumption plus net saving gives net income, and that consumption plus net investment gives net output. Appendix B will also question the latter teachings.
Embodied growth is disembodied growth on a finer scale. It redeploys or repurposes existing labor skills, raw materials, and plant capacity, as well as existing finished goods, to achieve higher returns than available from the customary uses which determine their prices. The present value of yields from this advantage in return, or equivalently the innovator’s reserve, defines the non-random component in free growth.
Appendix A.   Equivalence of saving and investment
Any usage which treats saving and investment as equal must deal with the fact that saving held in cash is not investment in the usual sense of spending on productive capital, and contributes nothing to output. Keynes (1936) defined that unspent saving as intended saving, and actual saving as the part so spent.
We suggest that investors seek to maximize return, within risk tolerance, and will sometimes hold saving in cash or equivalents at zero return in recessions and depressions when positive returns cannot be found, and so when investing in the usual sense would tend to reduce capital and output rather than increase them. Thus investment could have the usual meaning in usual times, when prospective returns bring animal spirits, and could include saving in cash when not. It is in this sense that we equate saving to investment.
Appendix B.  Net output with human capital
Human capital is impractical to measure, as it leaves little market record other than for its rental income in pay and investment cost in schooling. Thus national accounts leave it implicit, and allow us to infer what we can from data for pay and schooling. Those accounts are founded on the principle, sound when terms are appropriately redefined, that net output, or value added, is expressed in the sum of capital growth and net outflow from the value-added chain. In national accounts, then, where physical capital is the whole of capital while net outflow of the chain is the whole of consumption, the reasoning is
[…]
It is possible in principle to model a value-added chain which includes human capital, and to compare findings with those shown in Eq. (B.1). Let human capital $H$, in that new model, stand as the last link in the value-added chain. Adapting the classic illustration of the value added principle, say that farms produce wheat, mills convert the wheat to flour, bakeries convert the flour into bread, and humans convert some of the bread, called invested consumption, into human capital. The net outflow from this extended value-added chain is not all of consumption, but only pure consumption, meaning the part remaining after the part invested in human capital (invested consumption) is subtracted [^13] The concepts of invested consumption, pure consumption, self-invested work and human depreciation were introduced in Schultz (1961).. By this reasoning, the principle that net output is expressed in capital growth plus net outflow gives
$Y=\Delta K+\Delta H+C_{p}\ ,\quad\text{allowing human capital,}$, (B.2)
where $C_{p}$ gives pure consumption.
Yoram Ben-Porath (1967) reasoned that growth in human capital equals invested consumption plus self-invested work less human depreciation [^14] Equation 4 in Ben-Porath’s paper, summarizing his first three equations. His terms and notation differ from ours. The concepts of invested consumption were also introduced by Schultz (1961).. Let $C_{s},\ W_{s}$ and $D(H)$ show these flows respectively. Thus the combined arguments of Schultz and Ben-Porath arrive at
$C=C_{s}+C_{p}\quad\text{and}\quad\Delta H=C_{s}+W_{s}-D(H)\ ,\quad\text{allowing human capital.}$. (B.3)
Substitution of these equations into Eq. (B.2) finds
$Y=\Delta K+C_{s}+W_{s}-D(H)+C_{p}\quad\text{and consequently}$
$Y=\Delta K+C+W_{s}-D(H)\ ,\quad\text{allowing human capital,}$, (B.4)
if Schultz and Ben-Porath and the reasoning here are right.
Appendix C.  Mill’s statement of the free growth idea
Mill (1848), book 1, chapter 5, section 4, includes:
If it were said, for instance, that the only way to accelerate the increase of capital is by increase of saving, the idea would probably be suggested of greater abstinence, and increased privation. But it is obvious that whatever increases the productive power of labour creates an additional fund to make savings from, and enables capital to be enlarged not only without additional privation, but concurrently with an increase of personal consumption.
This passage may be the first clear statement of what we call free growth theory. His use of the words "accelerate" and "concurrently" suggest that his path of reasoning was something like ours, and that he meant what we call free growth rather than the alternating phases of higher and lower saving rates pictured in thrift theory.
[…]

### v8

@@
 Some other economists, including John Rae (1834) and John Stuart Mill (1848) (see Appendix C), argued that capital growth might also be explained by a rise in productivity of capital and labor already extant.
-Ways might found for existing factors to produce more, that is, and so to allow more consumption, or more capital growth, or any mix of the two, without inputs of net saving.
+Ways might be found for existing factors to produce more, that is, and so to allow more consumption, or more capital growth, or any mix of the two, without inputs of net saving.
 Robert Solow (1957) allowed that possibility for “disembodied” growth, where plant and products already existing are repurposed or redeployed in more productive ways.
-We test between those two explanations of capital growth, by net saving or by increase in productivity of capital and labor already in existence, by comparing net saving to concurrent change in market-value capital in 86 countries.
-As changes in net saving are expected to be associated with opposite changes in consumption, we also compare change in consumption to concurrent change in capital growth (capital acceleration).
+We test between those two explanations of capital growth, by net saving or by increase in productivity of capital and labor already in existence, by comparing net saving to concurrent change in market-value capital in 92 countries.
 All data are drawn from national accounts of those countries as collated on the free website World Inequality Database.
-Those test results from data for net saving/capital ratio and for consumption/capital ratio are easiest to reconcile in an interpretation that change in neither affects change in capital growth rate.
-[^2] Another interpretation allows that net saving may account for 6.4% of capital growth.
-See Section 4 below.
-This finding would support the views of Rae and Mill, and of Solow as to disembodied growth.
-It would suggest that capital growth, even in acceleration, arrives without help from net saving or consumption restraint.
+Those test results from data for net saving/capital ratio are easiest to reconcile in an interpretation that net saving has no effect on capital growth.
 Net saving, if so, raises the physical quantity of capital, but not the aggregate value, and so reduces the value per unit.
@@
 It predicts only at the largest scales, and only for the private sector.
-Individuals and groups and even small economies can grow through investment from outside.
+Individuals and groups and even small economies can grow or decline through net saving or dissaving.
 That possibility is foreclosed only at the scale of all capital and all economies together.
@@
 2 Net saving, capital growth and capital acceleration
-This study compares net saving $S_{net}$ to concurrent growth in market-value capital $K$ from data in national accounts.
-Capital growth ${\Delta K}_{i}$ in each year $i$ for each reporting country is found as $K_{i}-K_{i-1},$ and compared to $S_{net,i}$ reported for that year and country.
-As we will be testing for differences between $\Delta K$ and $S_{net,i}$, we begin by writing
-$\Delta K_{i}=z_{i}S_{net,i}+Q_{i}\ .$ (1)
-Here "net saving efficiency" $z$ is the fraction of $S_{net}$ actually realized in growth of capital value $K$.
-$Q$ means the sum of market noise, which may prove positive or negative or zero, plus any part of capital growth explained by concurrent productivity gain as described by Rae and Mill, and by Solow as to disembodied growth.
-We call this sum of noise and productivity gain “free growth” in that it costs no net saving.
-The object of testing is to find effects of $S_{net}$ on concurrent $\Delta K$, and so to help evaluate historic and current teachings as to those effects.
-We submitted above that most teaching, with exceptions noted as to Rae, Mill, and Solow, and within the warranted rate, predicts net saving to differ from concurrent growth in market-value capital only by market noise which tends to balance out over scale and time.
-Free growth $Q$ in Eq.
-(1), in that case, will give the market noise converging to zero.
-That consensus prediction, which we will challenge, will here be called “thrift theory”; $S_{net}$, in thrift theory, is expected to converge to $\Delta K$ if held within the warranted rate.
-That is,
-$E(z)=1\ ,\quad E(\Delta K)=S_{net}\quad\text{and}\quad E(Q)=0,\quad\text{if}\quad\frac{\Sigma S_{net}}{\Sigma K}\leq u\ ,\quad\text{in thrift theory,}$, (2)
-where:
-  1. 1.
-    $\Sigma S_{net}$ is collective net saving over the economy
-  2. 2.
-    $\Sigma K$ is collective capital over the economy before current $\Sigma S_{net},$ and
-  3. 3.
-    $u$ is the warranted rate.
-Free growth theory makes the opposite prediction that net saving adds no capital value, and that capital growth, at the largest scales in the private sector, is explained wholly by increased productivity of skills and materials already present.
-Thus, referring to Eq.
-(1), it predicts:
-$E(z)=0\quad\text{and}\quad E(\Delta K)=Q\quad\text{in free growth theory.}$. (3)
-$E(z)$, $E(\Delta K)$ and $E(Q)$ here give the expected values of $z$, $\Delta K$ and $Q$ respectively.
-Expected value means predicted average of outcomes over all observations.
-We will test those predictions against data for saving and capital growth taken from national accounts.
-As secular economic growth has tended to make later stocks and flows larger than earlier ones, we first divide by $K$ (normalize) to avoid overweighting of more recent years in finding that average.
+Thrift theory will mean the combined ideas that all net saving $S_{net}$ is realized in equal growth of capital $K$, if $S_{net}$ holds within the warranted rate, and that there is no other source of capital growth except market noise regressing to zero.
+"Thrift assumptions" will mean those two plus the assumptions that current $S_{net}$ holds within the warranted rate.
+The acronym uta will mean "under thrift assumptions." Then
+$\Delta K=S_{net}\ ,\quad\text{uta}.$ (1)
+Define "thrift" $s^{*}$ by $s^{*}=\frac{S_{net}}{K}$, and divide Eq. (1) by capital to predict capital growth rate $g(K)$ as
+$g(K)=s^{*}\ ,\quad\text{uta,}\quad\text{where }s^{*}=\frac{S_{net}}{K}.\quad\text{ Then}$ (2)
+[…]
+$\Delta g(K)$ and $\Delta s^{*}$ respectively will be called "capital acceleration" and "thrift change," either of which may be positive or negative.
 Division of Eq.
-(1) by $K_{i-1}$ gives
-$\frac{(\Delta K)_{i}}{K_{i-1}}=\frac{z_{i}S_{net,i}}{K_{i-1}}+\frac{Q_{i}}{K_{i-1}}\ .$ (4)
-The first term in Eq.
-(4) gives capital growth rate $g(K)$.
-The fraction in the second term is a variant of the Keynesian net saving rate $s_{net}$ where capital rather than output becomes the denominator.
-This flow will show as $s^{*}$, with the subscript “net” left implicit, and with the understanding that the denominator shows market-value capital rather than output.
-The third term in Eq.
-(4) will be called free growth rate and shown as $q_{s}$.
-That is,
-$g{(K)}_{i}=\frac{(\Delta K)_{i}}{K_{i-1}}\ ,\quad s^{*}_{i}=\frac{S_{net,i}}{K_{i-1}}\ ,\quad\text{and}\quad q_{i}=\frac{Q_{i}}{K_{i-1}}\ ,$
-so that Eq. (4) can be shown more compactly as
-$g(K)_{i}=z_{i}s^{*}_{i}+q_{i}\ .$ (5)
-By the definition $q=\frac{Q}{K}$, an expected value $E(Q)=0$ implies $E(q)=0$. Application of Eq. (2) to Eq. (5) now gives
-$E(g(K))=s^{*}\quad\text{and}\quad E(q)\ ,\quad\text{under thrift assumptions,}$, (6)
-where “thrift assumptions” are that thrift theory is correct and that the warranted rate is not exceeded. Eqs. (2) and (4), meanwhile, give
-$E(g(K))=q\quad\text{in free growth theory.}$. (7)
-Note here that free growth theory makes no prediction as to the quantity $s^{*}_{i}$. Rather it predicts $E(z)=0$ in Eq. (5), so that $s^{*}_{i}$ takes no part in calculation of $\Delta g(K)_{i}$.
-The prediction in Eq. (6) can be tested directly from data for market-value capital and net saving in national accounts. We define the "auxiliary thrift index" by
-θ=sg(K),that is, and test the prediction
-[…]
-[…]
-$E(\theta^{*})=1\quad\text{under thrift assumptions.}$.
-Results show in Fig.
-1 and Tables 1 and 2.
-They show $\theta^{*}$ as averaging something near $\frac{1}{2}$, and with wide variance from country to country and from year to year.
-Regression of $s^{*}$ on $g(K)$ is found at 0.2103.
-Free growth theory, again, makes no prediction as to relationships between net saving and capital growth.
-The concepts of correlation, covariance and regression are founded on the principle that causal relationships among variables tend to be revealed in simultaneous changes.
-Thus the two theories are best tested in their implications for changes in capital growth, which we call capital acceleration and show as $\Delta g(K)$.
-First note that Eq.
-(5) allows
-$(\Delta g(K))_{i}=(\Delta(zs^{*}))_{i}+(\Delta g)_{i}\ ,\quad\text{where}$ (8)
-$(\Delta g{(K)})_{i}=(g{(K)})_{i}-(g(K))_{i-1}\ ,\quad(\Delta(zs^{*}))_{i}=z_{i}s^{*}_{i}-z_{i-1}s^{*}_{i-1}\ ,\quad\text{and}\quad(\Delta q)_{i}=q_{i}-q_{i-1}\ .$
-For any variables $a$ and $b$, we may reason $E(a-b)=E(a)-E(b)$. By this and by Eqs. (7) and (8), then,
-$E(g(K))=\Delta s^{*}\quad\text{and}\quad E(\Delta q)=0\ ,\quad\text{under thrift assumptions.}$. (9)
-Division of Eq. (8) by capital acceleration, and rearrangement, gives
-$\frac{(\Delta(zs^{*}))_{i}}{(\Delta g{(K)})_{i}}+\frac{(\Delta q)_{i}}{(\Delta g{(K)})_{i}}=1\ .$ (10)
-Fig. 1: $\theta^{*}$ (86 countries)
-Table 1: Regression of $s^{*}$ on $g(K)$ (Screen = 0.01). $H_{0}$ per thrift theory: $g(K)\cong s^{*}$ $\overline{\theta^{*}}\cong 1$.
-$s^{*}$
-Regression of $s^{*}$ on $g(K)$ 0.1857∗∗∗
+(3) by capital acceleration, with rearrangement, gives
+$\frac{\Delta s^{*}}{\Delta g(K)}=\frac{\Delta g(K)}{\Delta g(K)}=1\ ,\quad\text{uta}.$ (4)
+The argument has continued from growth in Eq.
+(2) to acceleration in Eqs.
+(3) and (4) on the reasoning that causal relationships among variables, in this case between thrift and capital growth rate, tend to be revealed most reliably in their concurrent changes.
+[^3] The ratio $\frac{\Delta A}{\Delta B}$, that is, compares changes in A to concurrent changes in B, and so suggests a causal relationship between A and B insofar as that ratio holds constant.
+This is the principle on which the concepts of correlation, covariance and regression are founded.
+Eq.
+(4) was derived from Eq.
+(3) so that success of predictions from growth assumptions can be measured against a standard of unity.
+For notational convenience, then, we define the "thrift index" $\theta$ by $\theta=\frac{\Delta s^{*}}{\Delta g(K)}$, and restate Eq.
+(4) as
+$\theta=1\ ,\quad\text{uta,}\quad\text{where}\ \theta=\frac{\Delta s^{*}}{\Delta g(K)}\ .$ (5)
+Test results for the predictions $\frac{s^{*}}{g(K)}=\frac{\Delta s^{*}}{\Delta g(K)}=1$, under thrift assumptions, appear in Fig.
+1 and in Tables 1 and 2.
+Results were found as GDP-weighted averages over all countries and years, from data for net saving and market-value capital taken from national accounts.
+They show $\frac{s^{*}}{g(K)}=0.481$, with regression of $s^{*}$ on $g(K)$ found at 0.0771, and $\frac{\Delta s^{*}}{\Delta g(K)}=0.064$, with regression of $\Delta s^{*}$ on $\Delta g(K)$ at 0.0559.
+Fig. 1: Average $\frac{s^{*}}{g(K)}$ with LOESS smoothing over all countries, GDP-weighted, 1980-2022
+Table 1: Regression of $s^{*}$ on $g(K)$ and $\Delta s^{*}$ on $\Delta g(K)$, GDP-weighted. Screen = 0.01. $H_{0}$ per thrift theory: $\frac{s^{*}}{g(K)}=\frac{\Delta s^{*}}{\Delta g(K)}=1$.
+Regression of $s^{*}$ on $g(K)$ 0.0771∗∗∗
 Observations 1,826
-R2 0.77165
-Within R2 0.13962
+R2 0.84511
+Within R2 0.10669
 Year fixed effects $\checkmark$
 Country fixed effects $\checkmark$
-Define $\theta_{s,i}=\frac{(\Delta zs^{*})_{i}}{(\Delta g{(K)})_{i}}$ and $\varphi_{s,i}=\frac{(\Delta q)_{i}}{(\Delta g{(K)})_{i}}$ to restate Eq. (10) as
-$\theta_{s,i}+\varphi_{s,i}=1\ .$ (11)
-Next define $\theta_{s}=E(\theta_{s,i})$ and $\varphi_{s}=E(\varphi_{s,i})$. By Eq. (11), then,
-[…]
-$\theta_{s}=1\quad\text{and}\quad\varphi_{s}=0\ ,\quad\text{under thrift assumptions.}$. (13)
-Eq. (11) enables measurement of effects of thrift and free growth in a given period $i$, that is, while Eq. (12) does the same for all periods together.
-$\theta$ and $\varphi$ will be called the “thrift index” and “free growth index” respectively.
-$\theta_{s}$ and $\varphi_{s}$ give their values as found from data for change in net saving.
-Expected values, again, are predicted averages of outcomes.
-Thus Eq.
-(13) and thrift theory can be tested by finding average values of $\theta_{s,i}$ and comparing findings to the expected value $\theta=1$.
-First we find
-$\overline{\theta_{s,i}}=\frac{1}{m}\sum^{m}\theta_{s,i}\quad\text{and}\quad\overline{\varphi_{s,i}}=\frac{1}{m}\sum^{m}\varphi_{s,i}\ ,$
-Table 2: Average $\theta^{*}$ in 86 countries (screen = 0.01). Number of years clearing screen shown in ()
-Country Period $\overline{\theta^{*}}$ Armenia 1996 - 2020 (23) -0.63 Aruba 1996 - 2001 (6) 1.25 Australia 1980 - 2018 (33) 0.24 Austria 1996 - 2021 (23) 0.83 Azerbaijan 1996 - 2020 (23) 1.57 Bahrain 2009 - 2013 (5) -2.43 Belgium 1996 - 2021 (21) 0.61 Bolivia 1998 - 2015 (17) 0.40 Botswana 1996 - 1999 (4) 3.54 Brazil 1996 - 2019 (21) 0.25 British Virgin Islands 1996 - 1999 (4) 2.45 Bulgaria 1996 - 2016 (16) 0.14 Burkina Faso 2000 - 2019 (20) 0.23 Cameroon 1996 - 2019 (24) 1.10 Canada 1980 - 2022 (38) 0.37 Cape Verde 2008 - 2017 (9) 0.53 Chile 1997 - 2021 (19) 0.84 China 1992 - 2020 (29) 0.81 Colombia 1996 - 2022 (27) 0.60 Costa Rica 2013 - 2020 (6) 0.21 Croatia 1996 - 2021 (24) 0.43 Curaçao 2001 - 2016 (14) 1.26 Cyprus 1996 - 2021 (25) 0.07 Czechia 1994 - 2020 (20) 0.17 Côte d’Ivoire 1996 - 2000 (5) 0.29 Denmark 1996 - 2022 (26) 0.22 Dominican Republic 1996 - 2016 (10) 1.52 Ecuador 2008 - 2020 (13) 0.53 Egypt 1997 - 2015 (19) 0.92 El Salvador 2015 - 2019 (5) -0.62 Estonia 1996 - 2021 (24) 0.45 Finland 1996 - 2021 (23) 0.17 France 1980 - 2021 (35) 0.26 Germany 1980 - 2022 (38) 0.70 Greece 1996 - 2021 (26) 0.17 Guatemala 2002 - 2021 (20) -0.72 Guinea 2004 - 2010 (5) 0.72 Honduras 2001 - 2015 (14) -0.57 Hong Kong 1997 - 2011 (15) 0.90 Hungary 1996 - 2021 (21) 0.14 Iceland 2001 - 2014 (14) 0.22 India 1999 - 2019 (21) 0.57 Indonesia 2017 - 2019 (3) 0.94 Iran 1997 - 2017 (20) 4.96 Ireland 1996 - 2021 (22) 0.26 Israel 1996 - 2019 (24) 0.46 Country Period $\overline{\theta^{*}}$ Italy 1980 - 2022 (36) 0.13 Japan 1980 - 2021 (29) 0.19 Kazakhstan 1997 - 2022 (25) 0.60 Kuwait 2003 - 2017 (14) 1.11 Kyrgyzstan 1996 - 2021 (24) 0.46 Latvia 1996 - 2021 (24) -0.23 Lithuania 1996 - 2021 (24) 0.13 Luxembourg 1996 - 2021 (22) 0.03 Malaysia 2007 - 2015 (9) 2.52 Malta 1996 - 2021 (23) 0.16 Mauritius 2014 - 2018 (5) 0.04 Mexico 1996 - 2021 (22) 0.15 Moldova 1996 - 2019 (22) -0.91 Mongolia 2006 - 2020 (15) 0.10 Morocco 1999 - 2021 (23) 1.13 Netherlands 1996 - 2021 (23) 0.51 New Zealand 1996 - 2019 (21) 0.78 Nicaragua 2007 - 2018 (12) 0.08 Niger 1996 - 2019 (24) 0.90 Norway 1981 - 2021 (38) 0.95 Peru 2008 - 2021 (14) 0.76 Philippines 1996 - 2022 (27) 1.76 Poland 1996 - 2021 (25) 0.71 Portugal 1996 - 2022 (24) 0.00 Qatar 2002 - 2017 (14) 1.89 Romania 1996 - 2020 (22) 0.03 Russia 1996 - 2017 (13) -0.19 Saudi Arabia 2003 - 2009 (7) 2.65 Senegal 2015 - 2021 (7) 0.24 Serbia 1998 - 2021 (23) 0.61 Slovakia 1996 - 2022 (24) 0.15 Slovenia 1996 - 2021 (20) 0.28 South Africa 1996 - 2022 (24) 0.28 South Korea 1996 - 2020 (24) 0.63 Spain 1995 - 2021 (25) 0.32 Sweden 1980 - 2021 (37) 0.41 Switzerland 1993 - 2021 (25) 0.69 Tunisia 1996 - 2011 (16) 0.29 Turkey 2010 - 2017 (8) 1.32 USA 1980 - 2021 (38) 0.31 Ukraine 1996 - 2019 (21) 0.17 United Kingdom 1981 - 2021 (36) 0.16 Uruguay 2016 - 2017 (2) 1.16 Uzbekistan 2011 - 2021 (10) 0.21 Vanuatu 2002 - 2007 (6) 0.29 Venezuela 1998 - 2019 (20) 0.55
-Note: Thrift theory predicts $\overline{\theta^{*}}\cong 1$. Free growth theory makes no prediction for these data.
-where $m$ is the number of observed values of $\theta_{s,i}$ and $\varphi_{s,i}$, and test the predictions
-$\overline{\theta_{s,i}}\cong 1\quad\text{and}\quad\overline{\varphi_{s,i}}\cong 0\ ,\quad\text{under thrift assumptions.}$.
-Calculations of $\overline{\theta_{s,i}}$ and $\overline{\varphi_{s,i}}$ are not expected to show 1 and 0 exactly, under thrift assumptions, so that $\overline{\theta_{s,i}}$ and $\overline{\varphi_{s,i}}$ are not exactly equal to $\theta_{s}$ and $\varphi_{s}$, because the number of samples $m$ is finite.
-Free growth theory, maenwhile, referring to Eq.
-(13), does not predict the exact opposite findings $\theta_{s}=0$ and $\varphi_{s}=1$.
-It rather expects a value of $\varphi_{s}$ somewhat less than unity for reasons shown in Section 4.
-Fig.
-2 shows average values of $\theta_{s,i}$ and $\varphi_{s,i}$ for 86 countries, both unweighted and weighted to GDP, over the period 1980-2022.
-Table 3 shows associated regressions of $\Delta s^{*}$ and $\Delta g_{s}$ respectively on $\Delta g(K)$ at 0.1000 and 0.9000.
-Average values of $\varphi_{s}$ for the 86 individual countries show in Table 4.
-To control distortions brought by small absolute denominators, years were screened out where $|\Delta g(K)|$ was found at less than 0.01 (see Section 13).
-Results show $\overline{\varphi_{s,i}}=0.936$ and $\overline{\theta_{s,i}}=0.064$ for data weighted to GDP.
-These findings largely support free growth theory, but could suggest a small role (6.4%) for thrift as well.
-This possibility will be discussed in Section 4.
-We will first test thrift theory from a different approach.
-Fig. 2: Thrift and free growth indexes derived from net saving (86 countries).
-Table 3: Regression of $\Delta s^{*}$ and $\Delta q_{s}$ on $\Delta g(K)$ (Screen = 0.01). $H_{0}\ \text{per thrift theory:}\ \Delta g(K)\cong\Delta s^{*}\ \&\ \Delta q_{s}\cong 0$
-$\Delta s^{*}$ $\Delta q_{s}$
-Regression of value shown on $\Delta g(K)$ 0.1000∗∗∗ 0.9000∗∗∗
-(0.0145) (0.0145)
-Observations 1,574 1,574
-R2 0.32867 0.96233
-Within R2 0.22713 0.95965
-Year fixed effects $\checkmark$ $\checkmark$
-Country fixed effects $\checkmark$ $\checkmark$
-3 Consumption and capital growth
-The concept of cash flow is little used in economics, aside from finance and in present value calculations. It is defined implicitly in the equation
-$P=\Delta K+F\ ,$ (14)
-where $P$ and $F$ mean net profit and cash flow respectively [^3] We have found no standard notation for cash flow..
-Cash flow means the net outflow of value, in cash or kind, from an asset to its owner for consumption $C$ or for reinvestment in other assets.
-This flow of spending is augmented by pay $\Pi$ [^4] Compensation of employees plus a share of mixed income.
-We avoid the standard notation $W$, which is also used for work, as we prefer to distinguish pay from work..
-At the scale of the closed economy, meaning of all assets together, reinvestment cancels to zero.
-At this scale, then, the sum of spending $F+\Pi$ is spent on consumption alone.
-That is,
-$\Pi+F=C$ (15)
-at the collective scale. Eq. (14), meanwhile, may be arranged as $F=P-\Delta K$. Substitution of this in Eq. (15), with rearrangement, gives
-$\Delta K=\Pi+P-C\ .$ (16)
-Thrift theory, which expects $\Delta K=S_{net}$, writes this as
-$S_{net}=\Pi+P-C\ ,\quad\text{under thrift assumptions.}$. (17)
-Pay, net profit [^5] Jordà et al.
-(2019), The rate of return on everything.
-and consumption can all be measured, to an extent, as can market-value capital, so that this prediction of thrift theory could be measured directly.
-Comparison between Eqs.
-(16) and (17), however, shows that such a test, given perfect data in national accounts, would only repeat that of Eq.
-(6).
-As before, the two theories are tested most revealingly in capital acceleration.
-Given the definitions [^6] The notation $c^{*}$, as with $s^{*}$, shows an asterisk to distinguish its meaning $c^{*}=\frac{C}{K}$ from the Keynesean $c=\frac{C}{Y}$
-$\pi_{i}=\frac{\Pi_{i}}{K_{i-1}}\ ,\quad p_{i}=\frac{P_{i}}{K_{i-1}}\quad\text{and}\quad c^{*}_{i}=\frac{C_{i}}{K_{i-1}}\ ,$
-Eq. (16) can be divided by capital to reach
-$g(K)_{i}=\pi_{i}+p_{i}-c^{*}_{i}\ ,\quad\text{implying}$ (18)
-$(\Delta g(K))_{i}=\Delta\pi_{i}+\Delta p_{i}-\Delta c^{*}_{i}\ ,\quad\text{where}$ (19)
-$\Delta\pi_{i}=\pi_{i}-\pi_{i-1}\ ,\quad\Delta p_{i}=p_{i}-p_{i-1}\quad\text{and}\quad\Delta c^{*}_{i}=c^{*}_{i}-c^{*}_{i-1}\ ,$
-as before.
-Thrift theory expects $\Delta g(K)=\Delta s^{*}=\Delta\pi+\Delta p-\Delta c$ by Eqs.
-(9) and (17).
-It further expects changes in productivity, as realized in the pay/capital and profit/capital ratios $\pi$ and $p$, to converge to zero as market noise, so that the only source of net saving is in consumption restraint $-\Delta c^{*}$.
-Thus
-$E(\Delta g(k))=-\Delta c^{*}\quad\text{and}\quad E(\Delta\pi)=E(\Delta p)=0\ ,\quad\text{under thrift assumptions.}$. (20)
-Free growth theory, conversely, predicts that net saving contributes nothing to capital growth, and hence that net saving through consumption restraint, at the collective scale and in the private sector, is sacrifice with nothing in return.
-Free growth theory expects rather that costless increases in the ratios of pay and net profit to capital, enabled by greater productivity of human and physical capital already present, account for all capital growth and acceleration.
-Referring to Eq.
-(19), however, it predicts
-[…]
-for reasons shown in Section 4. As in Eq. (10), we prepare to test thrift theory by first dividing Eq. (20) by capital acceleration to reach
-$E(\frac{-\Delta c^{*}}{\Delta g(K)})=1\ ,\quad\text{under thrift assumptions.}$. (22)
-Define $\theta_{c,i}=\frac{-\Delta c^{*}_{i}}{\Delta g(K)_{i}}$ and $\varphi_{c,i}=1-\theta_{c,i}$ to re-express Eq. (22) as
-$\theta_{c,i}+\varphi_{c,i}=1\ .$ (23)
-By Eqs. (20) through (23), further,
-$E\left(\theta_{c}\right)=1\quad\text{and}\quad E\left(\varphi_{c}\right)=0\ ,\quad\text{under thrift assumptions.}$. (24)
-Define $\theta_{c}=E\left(\theta_{c}\right)$ and $\varphi_{c}=E\left(\varphi_{c}\right)$ to re-express Eqs. (23) and (24) respectively as
-$\theta_{c}+\varphi_{c}=1\ ,\quad\text{and}$ (25)
-$\theta_{c}=1\quad\text{and}\quad\varphi_{c}=0\ ,\quad\text{under thrift assumptions.}$. (26)
-We infer $\theta_{c}\cong\overline{\theta_{c,i}}$ and $\varphi_{c}\cong\overline{\varphi_{c,i}}$ as before, and test thrift theory by comparing average yearly values of $\theta_{c,i}$ and $\varphi_{c,i}$ to its predictions $\overline{\theta_{c,i}}\cong 1$ and $\overline{\varphi_{c,i}}\cong 0$.
-[…]
-Fig. 3: Thrift and free growth indexes derived from consumption (86 countries).
-Table 4: Average $\varphi_{s,i}$ and $\varphi_{c,i}$ in 86 countries (screen = 0.01). Number of years clearing screen shown in ()
-Country Period $\overline{\varphi_{s,i}}$ $\overline{\varphi_{c,i}}$ Armenia 1997 - 2020 (20) 0.88 1.18 Aruba 1997 - 2001 (5) -0.22 8.44 Australia 1974 - 2019 (35) 0.97 1.08 Austria 1997 - 2021 (16) 0.90 1.13 Azerbaijan 1997 - 2020 (23) 0.25 2.20 Bahrain 2010 - 2013 (4) 0.59 1.32 Belgium 1997 - 2021 (19) 0.95 1.06 Bolivia 1999 - 2014 (12) 0.96 1.12 Botswana 1997 - 2000 (4) 0.76 1.27 Brazil 1998 - 2020 (22) 0.90 1.24 British Virgin Islands 1997 - 1998 (2) -1.24 -2.20 Bulgaria 1997 - 2017 (15) 1.03 1.19 Burkina Faso 2001 - 2019 (16) 0.96 1.28 Cameroon 1997 - 2019 (19) 0.82 2.71 Canada 1974 - 2022 (41) 0.91 1.08 Cape Verde 2009 - 2016 (8) 0.73 2.02 Chile 1998 - 2021 (22) 0.94 1.48 China 1993 - 2020 (23) 0.91 1.05 Colombia 1997 - 2022 (18) 1.03 2.39 Costa Rica 2014 - 2020 (6) 0.96 1.25 Croatia 1997 - 2021 (22) 0.81 2.09 Curaçao 2002 - 2016 (14) 0.59 0.51 Cyprus 1997 - 2021 (23) 1.06 1.16 Czechia 1995 - 2021 (19) 1.03 1.08 Côte d’Ivoire 1997 - 2000 (4) 1.03 1.05 Denmark 1997 - 2022 (22) 0.95 1.14 Dominican Republic 2007 - 2015 (9) 0.85 1.41 Ecuador 2009 - 2020 (12) 0.83 1.91 Egypt 1998 - 2015 (16) 0.87 2.14 El Salvador 2016 - 2019 (4) 0.02 -2.32 Estonia 1997 - 2021 (19) 0.94 1.11 Finland 1997 - 2022 (21) 1.02 1.20 France 1972 - 2021 (33) 0.91 1.16 Germany 1972 - 2022 (31) 0.95 1.14 Greece 1997 - 2021 (24) 0.76 1.47 Guatemala 2003 - 2021 (11) 0.85 1.32 Guinea 2005 - 2010 (6) 0.92 1.23 Honduras 2003 - 2015 (10) 0.60 0.85 Hong Kong 1997 - 2011 (13) 0.96 1.25 Hungary 1997 - 2021 (21) 0.93 1.10 Iceland 2002 - 2014 (12) 0.90 1.12 India 2000 - 2019 (14) 0.81 0.99 Indonesia 2019 - 2019 (1) 1.41 0.98 Iran 1998 - 2018 (19) -0.09 1.09 Ireland 1997 - 2021 (23) 0.98 1.23 Israel 1997 - 2017 (18) 0.93 1.72 Country Period $\overline{\varphi_{s,i}}$ $\overline{\varphi_{c,i}}$ Italy 1981 - 2021 (22) 0.94 1.12 Japan 1981 - 2021 (30) 0.95 1.13 Kazakhstan 1997 - 2022 (23) 0.92 1.16 Kuwait 2005 - 2017 (11) 0.96 1.13 Kyrgyzstan 1998 - 2021 (22) 0.77 1.29 Latvia 1998 - 2021 (20) 0.99 1.15 Lithuania 1997 - 2021 (18) 0.90 1.19 Luxembourg 1997 - 2021 (25) 1.00 1.02 Malaysia 2008 - 2015 (7) 0.99 1.19 Malta 1997 - 2021 (23) 0.92 1.04 Mauritius 2015 - 2018 (4) 0.98 0.99 Mexico 1997 - 2021 (22) 0.94 1.16 Moldova 1997 - 2019 (21) 0.89 1.27 Mongolia 2008 - 2021 (11) 0.90 1.06 Morocco 2000 - 2021 (17) 0.81 1.27 Netherlands 1997 - 2021 (20) 0.97 1.08 New Zealand 1997 - 2019 (18) 0.86 1.02 Nicaragua 2008 - 2019 (12) 0.94 1.13 Niger 1997 - 2019 (19) 0.78 1.87 Norway 1982 - 2021 (34) 0.88 1.08 Peru 2009 - 2021 (11) 0.96 1.30 Philippines 1997 - 2022 (20) 0.63 1.20 Poland 1997 - 2021 (15) 0.80 1.07 Portugal 1997 - 2022 (17) 0.98 1.08 Qatar 2004 - 2018 (12) 0.71 0.86 Romania 1997 - 2020 (14) 0.98 1.16 Russia 1997 - 2018 (14) 0.91 1.11 Saudi Arabia 2005 - 2009 (3) -1.23 0.02 Senegal 2018 - 2021 (3) 0.81 2.31 Serbia 1999 - 2021 (22) 0.78 1.67 Slovakia 1997 - 2021 (18) 0.91 1.18 Slovenia 1997 - 2021 (20) 0.90 1.07 South Africa 1997 - 2022 (21) 0.90 1.13 South Korea 1997 - 2020 (15) 0.96 1.05 Spain 1996 - 2021 (23) 0.96 1.11 Sweden 1974 - 2022 (41) 0.98 1.15 Switzerland 1993 - 2020 (24) 0.92 1.09 Tunisia 1997 - 2011 (12) 0.68 1.17 Turkey 2011 - 2015 (5) 1.16 1.54 USA 1972 - 2021 (44) 0.98 1.13 Ukraine 1997 - 2020 (23) 0.95 1.18 United Kingdom 1972 - 2021 (42) 0.98 1.21 Uruguay 2017 - 2017 (1) 0.44 0.98 Uzbekistan 2012 - 2021 (9) 0.95 1.14 Vanuatu 2003 - 2007 (3) 0.85 1.19 Venezuela 1999 - 2019 (20) 0.69 1.13
-Note: Thrift theory predicts $\overline{\varphi_{s,i}}\cong\overline{\varphi_{c,i}}\cong 0$. Free growth theory predicts $\overline{\varphi_{s,i}}\cong\overline{\varphi_{c,i}}\cong 1$.
-Table 5: Regression of $-\Delta c^{*}$ and $\Delta q_{c}$ on $\Delta g(K)$ (Screen = 0.01). $H_{0}\text{ per thrift theory: }\Delta g(K)\cong-\Delta c^{*}\&\Delta q_{c}\cong 0$
-$-\Delta c^{*}$ $\Delta q_{c}$
-Regression of value shown on $\Delta g(K)$0.3760∗∗∗ 1.376∗∗∗
-(0.1055) (0.1055)
-Observations 1,548 1,548
-R2 0.53993 0.88294
-Within R2 0.31853 0.86226
-Year fixed effects $\checkmark$ $\checkmark$
-Country fixed effects $\checkmark$ $\checkmark$
-4 Differences in test results
-Fig.
-2 and Table 3 show $\theta_{s}$ to average 0.064, and regression of $\Delta s^{*}$ on $\Delta g(K)$ to average $0.100$, across all years and countries GDP-weighted.
-Fig.
-3 and Table 5, meanwhile, showed these overall averages as $-0.142$ for $\theta_{c}$, and as $-0.376$ for regression of $-\Delta c^{*}$ on $\Delta g(K)$.
-These differences might be explained by systematic mismeasurements of $C$, $K$ and $S_{net}$ in national accounts.
-In that case, the evidence would leave open the possibility of some role for net saving (6.4%) in capital growth.
-There is another possible interpretation.
-Regression or correlation or covariance among variables such as net saving and capital growth suggests causal relationship among them, but does not in itself reveal the direction of causality.
-Thrift theory accepts the natural inference that net saving is the cause, and capital growth the effect.
-Logic and experience, however, also suggest that adverse shocks such as war or depression might lower capital values despite all saving and investment, and might then motivate negative net saving (dissaving), meaning suspension of replacement saving, in order to maintain consumption.
-When peace and prosperity return, higher incomes will allow restoration of replacement saving to offset depreciation.
-Capital growth need not depend on positive net saving at any point, but rather allows an upward normalization of net saving from a negative value to zero.
-In this interpretation, change in capital growth becomes the cause, rather than the effect, of change in net saving.
-This possibility could help explain differences between the measurements of GDP-weighted $\varphi_{s}(0.936)$ and $\varphi_{c}(1.142)$, without need to assume systematic mismeasurements of $S_{net}$, $K$ and $C$.
-Note here that the finding $\varphi>1$ is possible in principle when gain in productivity allows growth in capital and consumption at the same time (see the quote from Mill in Appendix C.) We welcome further study to resolve these uncertainties.
+Regression of $\Delta s^{*}$ on $\Delta g(K)$ 0.0559∗∗∗
+Observations 1,574
+R2 0.37071
+Within R2 0.27233
+Year fixed effects $\checkmark$
+Country fixed effects $\checkmark$
+Table 2: Average $\frac{s^{*}}{g(K)}$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()
+Armenia 1996 - 2020 (23) -0.63 Aruba 1996 - 2001 (6) 1.25 Australia 1980 - 2018 (33) 0.24 Austria 1996 - 2021 (23) 0.83 Azerbaijan 1996 - 2020 (23) 1.57 Bahrain 2009 - 2013 (5) -2.43 Belgium 1996 - 2021 (21) 0.61 Bolivia 1998 - 2015 (17) 0.40 Botswana 1996 - 1999 (4) 3.54 Brazil 1996 - 2019 (21) 0.25 British Virgin Islands 1996 - 1999 (4) 2.45 Bulgaria 1996 - 2016 (16) 0.14 Burkina Faso 2000 - 2019 (20) 0.23 Cameroon 1996 - 2019 (24) 1.10 Canada 1980 - 2022 (38) 0.37 Cape Verde 2008 - 2017 (9) 0.53 Chile 1997 - 2021 (19) 0.84 China 1992 - 2020 (29) 0.81 Colombia 1996 - 2022 (27) 0.60 Costa Rica 2013 - 2020 (6) 0.21 Croatia 1996 - 2021 (24) 0.43 Curaçao 2001 - 2016 (14) 1.26 Cyprus 1996 - 2021 (25) 0.07 Czechia 1994 - 2020 (20) 0.17 Côte d’Ivoire 1996 - 2000 (5) 0.29 Denmark 1996 - 2022 (26) 0.22 Dominican Republic 1996 - 2016 (10) 1.52 Ecuador 2008 - 2020 (13) 0.53 Egypt 1997 - 2015 (19) 0.92 El Salvador 2015 - 2019 (5) -0.62 Estonia 1996 - 2021 (24) 0.45 Finland 1996 - 2021 (23) 0.17 France 1980 - 2021 (35) 0.26 Germany 1980 - 2022 (38) 0.70 Greece 1996 - 2021 (26) 0.17 Guatemala 2002 - 2021 (20) -0.72 Guinea 2004 - 2010 (5) 0.72 Honduras 2001 - 2015 (14) -0.57 Hong Kong 1997 - 2011 (15) 0.90 Hungary 1996 - 2021 (21) 0.14 Iceland 2001 - 2014 (14) 0.22 India 1999 - 2019 (21) 0.57 Indonesia 2017 - 2019 (3) 0.94 Iran 1997 - 2017 (20) 4.96 Ireland 1996 - 2021 (22) 0.26 Israel 1996 - 2019 (24) 0.46 Italy 1980 - 2022 (36) 0.13 Japan 1980 - 2021 (29) 0.19 Kazakhstan 1997 - 2022 (25) 0.60 Kuwait 2003 - 2017 (14) 1.11 Kyrgyzstan 1996 - 2021 (24) 0.46 Latvia 1996 - 2021 (24) -0.23 Lithuania 1996 - 2021 (24) 0.13 Luxembourg 1996 - 2021 (22) 0.03 Malaysia 2007 - 2015 (9) 2.52 Malta 1996 - 2021 (23) 0.16 Mauritius 2014 - 2018 (5) 0.04 Mexico 1996 - 2021 (22) 0.15 Moldova 1996 - 2019 (22) -0.91 Mongolia 2006 - 2020 (15) 0.10 Morocco 1999 - 2021 (23) 1.13 Netherlands 1996 - 2021 (23) 0.51 New Zealand 1996 - 2019 (21) 0.78 Nicaragua 2007 - 2018 (12) 0.08 Niger 1996 - 2019 (24) 0.90 Norway 1981 - 2021 (38) 0.95 Peru 2008 - 2021 (14) 0.76 Philippines 1996 - 2022 (27) 1.76 Poland 1996 - 2021 (25) 0.71 Portugal 1996 - 2022 (24) 0.00 Qatar 2002 - 2017 (14) 1.89 Romania 1996 - 2020 (22) 0.03 Russia 1996 - 2017 (13) -0.19 Saudi Arabia 2003 - 2009 (7) 2.65 Senegal 2015 - 2021 (7) 0.24 Serbia 1998 - 2021 (23) 0.61 Slovakia 1996 - 2022 (24) 0.15 Slovenia 1996 - 2021 (20) 0.28 South Africa 1996 - 2022 (24) 0.28 South Korea 1996 - 2020 (24) 0.63 Spain 1995 - 2021 (25) 0.32 Sweden 1980 - 2021 (37) 0.41 Switzerland 1993 - 2021 (25) 0.69 Tunisia 1996 - 2011 (16) 0.29 Turkey 2010 - 2017 (8) 1.32 USA 1980 - 2021 (38) 0.31 Ukraine 1996 - 2019 (21) 0.17 United Kingdom 1981 - 2021 (36) 0.16 Uruguay 2016 - 2017 (2) 1.16 Uzbekistan 2011 - 2021 (10) 0.21 Vanuatu 2002 - 2007 (6) 0.29 Venezuela 1998 - 2019 (20) 0.55
+Note: Thrift theory predicts $\frac{s^{*}}{g(K)}=1$. Free growth theory makes no prediction for these data.
+Table 3: Average $\theta_{s}^{*}$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()
+Armenia 1997 - 2020 (20) 0.12 Aruba 1997 - 2001 (5) 1.22 Australia 1974 - 2019 (35) 0.03 Austria 1997 - 2021 (16) 0.10 Azerbaijan 1997 - 2020 (23) 0.75 Bahrain 2010 - 2013 (4) 0.41 Belgium 1997 - 2021 (19) 0.05 Bolivia 1999 - 2014 (12) 0.04 Botswana 1997 - 2000 (4) 0.24 Brazil 1998 - 2020 (22) 0.10 British Virgin Islands 1997 - 1998 (2) 2.24 Bulgaria 1997 - 2017 (15) -0.03 Burkina Faso 2001 - 2019 (16) 0.04 Cameroon 1997 - 2019 (19) 0.18 Canada 1974 - 2022 (41) 0.09 Cape Verde 2009 - 2016 (8) 0.27 Chile 1998 - 2021 (22) 0.06 China 1981 - 2020 (33) 0.07 Colombia 1997 - 2022 (18) -0.03 Costa Rica 2014 - 2020 (6) 0.04 Croatia 1997 - 2021 (22) 0.19 Curaçao 2002 - 2016 (14) 0.41 Cyprus 1997 - 2021 (23) -0.06 Czechia 1995 - 2021 (19) -0.03 Côte d’Ivoire 1997 - 2000 (4) -0.03 Denmark 1997 - 2022 (22) 0.05 Dominican Republic 2007 - 2015 (9) 0.15 Ecuador 2009 - 2020 (12) 0.17 Egypt 1998 - 2015 (16) 0.13 El Salvador 2016 - 2019 (4) 0.98 Estonia 1997 - 2021 (19) 0.06 Finland 1997 - 2022 (21) -0.02 France 1972 - 2021 (33) 0.09 Germany 1972 - 2022 (31) 0.05 Greece 1997 - 2021 (24) 0.24 Guatemala 2003 - 2021 (11) 0.15 Guinea 2005 - 2010 (6) 0.08 Honduras 2003 - 2015 (10) 0.40 Hong Kong 1997 - 2011 (13) 0.04 Hungary 1997 - 2021 (21) 0.07 Iceland 2002 - 2014 (12) 0.10 India 1997 - 2019 (16) 0.14 Indonesia 2019 - 2019 (1) -0.41 Iran 1998 - 2018 (19) 1.09 Ireland 1997 - 2021 (23) 0.02 Israel 1997 - 2017 (18) 0.07 Italy 1972 - 2021 (29) 0.05 Japan 1972 - 2021 (38) 0.05 Kazakhstan 1997 - 2022 (23) 0.08 Kuwait 2005 - 2017 (11) 0.04 Kyrgyzstan 1998 - 2021 (22) 0.23 Latvia 1998 - 2021 (20) 0.01 Lithuania 1997 - 2021 (18) 0.10 Luxembourg 1997 - 2021 (25) 0.00 Malaysia 2008 - 2015 (7) 0.01 Malta 1997 - 2021 (23) 0.08 Mauritius 2015 - 2018 (4) 0.02 Mexico 1997 - 2021 (22) 0.06 Moldova 1997 - 2019 (21) 0.11 Mongolia 2008 - 2021 (11) 0.10 Morocco 2000 - 2021 (17) 0.19 Netherlands 1997 - 2021 (20) 0.03 New Zealand 1997 - 2019 (18) 0.14 Nicaragua 2008 - 2019 (12) 0.06 Niger 1997 - 2019 (19) 0.22 Norway 1982 - 2021 (34) 0.12 Peru 2009 - 2021 (11) 0.04 Philippines 1997 - 2022 (20) 0.37 Poland 1997 - 2021 (15) 0.20 Portugal 1997 - 2022 (17) 0.02 Qatar 2004 - 2018 (12) 0.29 Romania 1997 - 2020 (14) 0.02 Russia 1997 - 2018 (14) 0.09 Saudi Arabia 2005 - 2009 (3) 2.23 Senegal 2018 - 2021 (3) 0.19 Serbia 1999 - 2021 (22) 0.22 Slovakia 1997 - 2021 (18) 0.09 Slovenia 1997 - 2021 (20) 0.10 South Africa 1997 - 2022 (21) 0.10 South Korea 1997 - 2020 (15) 0.04 Spain 1972 - 2021 (42) 0.06 Sweden 1974 - 2022 (41) 0.02 Switzerland 1993 - 2020 (24) 0.08 Tunisia 1997 - 2011 (12) 0.32 Turkey 2011 - 2015 (5) -0.16 USA 1972 - 2021 (44) 0.02 Ukraine 1997 - 2020 (23) 0.05 United Kingdom 1972 - 2021 (42) 0.02 Uruguay 2017 - 2017 (1) 0.56 Uzbekistan 2012 - 2021 (9) 0.05 Vanuatu 2003 - 2007 (3) 0.15 Venezuela 1999 - 2019 (20) 0.31
+Note: Thrift theory predicts $\theta_{s}^{*}=1$.
+Fig. 2: Average $\theta_{s}^{*}$ with LOESS smoothing over all countries, GDP-weighted, 1980-2022
+3 Interpretation of test results
+The finding $\frac{s^{*}}{g(K)}=0.481$ reveals average observed proportions between $s^{*}$ and $g(K)$, and does not of itself reveal causal relatedness between those two variables.
+Causal relatedness is shown rather by the degree of constancy in proportions, and thus on regression of $s^{*}$ on $g(K)$ found at 0.0771.
+$\theta_{s}$ or $\frac{\Delta s^{*}}{\Delta g(K)}$, measured at $\theta_{s}^{*}=0.064$, gives a separate measure of causal relatedness between $s^{*}$ and $g(K)$ as explained.
+Regression of $\Delta s^{*}$ on $\Delta g(K)$, found at 0.0559, gives a measure of causal relatedness between $\Delta s^{*}$ and $\Delta g(K)$.
+Might these small but positive findings allow the possibility that some capital growth is explained by net saving?
+Any explanation must account for the chance that replacement saving will sometimes not be enough to offset depreciation.
+This shortfall is likeliest in prolonged downturns, where loss of income motivates households to forego net saving, and then to invade capital by putting off maintenance expense and replacement investment in order to protect consumption.
+Effects will include a positive effect on regression of $s^{*}$ on $g(K)$, and an upward effect on $\frac{\Delta s^{*}}{\Delta g(K)}$, as $s^{*}$ and $g(K)$ decline together in response to downward pressure on income.
+This interpretation could fit the appearance in Fig.
+2 that $\frac{\Delta s^{*}}{\Delta g(K)}$, shown as $\theta^{*}_{s}$, found higher values during the period of the dot com and subprime crises.
+In general, it might be expected that households stabilize consumption by building up reserves through net saving when earnings are higher, and deplete them when earnings are lower, as described in the life cycle theory of saving (Modigliani (1954)) or the permanent income hypothesis (Friedman (1957)).
+Free growth theory explains the rise in earnings as caused by a rise in productivity of current plant and implements and skills, and so predicts the same effects on regressions and $\frac{\Delta s^{*}}{\Delta g(K)}$ as does thrift theory, but from an opposite causality.
+Although these results cannot exclude the possibility that some net saving has effect, data are easiest to reconcile with free growth theory, where all capital growth is explained by productivity gain.
+Pending further study, we will proceed on this interpretation.
 4 Mechanics of free growth
@@
 The innovator pays for these inputs at a market price determined by their value in established productive uses, but applies them innovatively to realize higher prospective cash flows, and hence higher present values, to the innovator (Marshall (1890), Schumpeter and Opie (1934)).
-This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.888i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns.
+This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.555i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns.
 The present value of additional cash flow enabled by this advantage in return quantifies the innovator’s reserve and equivalently the non-random component of free growth.
@@
 5 Optimum investment policy
-Data and arguments adduced, with uncertainties discussed in Section 4, suggest that the optimum amount of saving, at the global scale, is replacement saving to offset depreciation, and nothing more.
+Data and arguments suggest that the optimum amount of saving, at the global scale, is replacement saving to offset depreciation, and nothing more.
 That would not mean book depreciation, as this study has stressed differences between book and market values.
@@
 6 Inferences for national accounts and macroeconomics
-Eq.
-(16) was presented as a tautology or logical certainty.
-Cash flow $F$ from all physical capital $K$ together at the collective scale, that is, means outflow of value from $K$ net of any inflow into $K$ from any source, meaning from pay $\Pi$ as well as plowback of cash flow.
-Thus the sum $F+\Pi$ is spent on consumption $C$ alone by exhaust of alternatives.
-Meanwhile cash flow is net profit $P$ less capital growth $\Delta K$ [^10] Or equivalently gross profit less capital growth less depreciation.
-This reasoning gave $F+\Pi=C$, $F=P-\Delta K$, and consequently $\Pi+P=C+\Delta K$ (Eq.
-(16)).
-Kuznets and the other founders of national accounts, under the influence of Keynes, seem to have reasoned alike.
-National accounts are double-entry $T$ accounts where net product $C+\Delta K$ is equated to net "factor payments" $\Pi+P$.
-National accounts, however, equate capital growth to net saving/investment, and effectively write Eq.
-(16) as $\Pi+P=C+S_{net}$.
 We have attempted to show that net saving does not tend to equate to capital growth, and indeed contributes no part of capital growth.
@@
 This source collates data from national accounts and tax data of 105 countries in constant currency units, and adjusts them where needed to conform to current standards of the System of National Accounts (SNA) published by the United Nations.
-We show results for the 86 of those countries which report all three of the factors, namely net saving, consumption and market-value capital, needed for deriving the thrift and free growth indexes.
+We show results for the 92 of those countries which report both factors, net saving and market-value capital, needed for deriving the thrift indexes.
 WID’s source for these data is national accounts.
-Consumption $C$ in our text and equations is reproduced from the sum of Government Final Consumption Expenditure (mcongo) [^11] WID code and Private Expenditures of Households and NPISH (mconhn).
-This sums government expenditure GCE and personal consumption expenditure PCE.
 Net saving $S_{net}$ and market-value $K$ are taken from net national saving (msavin) and market-value Capital Wealth (mnweal) respectively.
@@
 12 Displays
-Eqs.
-(12) and (25) show $\theta_{s}+\varphi_{s}=1$ and $\theta_{c}+\varphi_{c}=1$.
-Some displays here and in the web appendix save space by showing $\varphi_{s}$ and $\varphi_{c}$ only, leaving $\theta_{s}$ and $\theta_{c}$ implicit as their complements to unity.
-The web appendix includes displays of $\varphi_{s}$ and $\varphi_{c}$ for each year in each country over the report period.
-These tend to show upward and downward spikes in values of $\varphi_{s}$ and $\varphi_{c}$ in some years.
+The web appendix includes displays of $\theta_{s}^{*}$ for each year in each country over the report period.
+These tend to show upward and downward spikes in values of $\theta_{s}^{*}$ in some years.
 Those spikes tend to be associated with small absolute values of denominators, in these cases $\Delta g(K)$, in those countries and years.
@@
 Some displays show data for all years, regardless of denominator size.
-Others screen out all years where absolute denominators are less then .01, then .025, then continuing upward in increments of .025 to a maximum screen of .15.
-$\varphi_{s}$ and $\varphi_{c}$ are plotted for each country unscreened and at each of the seven successive levels of screening.
-All displays applied a screen of .01.
+Others screen out all years where absolute denominators are less then 0.01, then 0.025, then continuing upward in increments of 0.025 to a maximum screen of 0.15.
+$\theta_{s}^{*}$ is plotted for each country unscreened and at each of the seven successive levels of screening.
+All displays applied a screen of 0.01.
 The denominator whose absolute value is screened is capital acceleration $\Delta g(K)$ or capital growth rate $g(K)$ in all displays.
@@
 These findings challenge the teachings that capital growth is effected by net saving enabled by consumption restraint, and that producer cost, including imputed interest as the opportunity cost of capital, converges to market realization.
-Evidence showing $\overline{\varphi_{s,i}}\cong 1$ and $\overline{\varphi_{c,i}}\cong 1$ suggests that all capital growth is free, and consequently that market realization, in the presence of innovation, exceeds producer cost by the entirety of capital growth.
+Evidence suggests that all capital growth is free, and consequently that market realization, in the presence of innovation, exceeds producer cost by the entirety of capital growth.
 Meanwhile the same evidence, which indicates that net saving adds no capital value, suggests a review of the teachings that consumption plus net saving gives net income, and that consumption plus net investment gives net output.
@@
 In national accounts, then, where physical capital is the whole of capital while net outflow of the chain is the whole of consumption, the reasoning is
-[…]
+$Y=\Delta K+C\ ,\quad\text{neglecting human capital.}$. (B.1)
 It is possible in principle to model a value-added chain which includes human capital, and to compare findings with those shown in Eq.
@@
 Yoram Ben-Porath (1967) reasoned that growth in human capital equals invested consumption plus self-invested work less human depreciation [^8] Equation 4 in Ben-Porath’s paper, summarizing his first three equations.
-His terms and notation differ from ours.
-The concepts of invested consumption were also introduced by Schultz (1961)..
+His terms and notation differ from ours..
 Let $C_{s},\ W_{s}$ and $D(H)$ show these flows respectively.

### v9

@@
 Abstract
-Data from national accounts show no effect of change in net saving or consumption, in ratio to market-value capital, on change in growth rate of market-value capital (capital acceleration).
-Thus it appears that capital growth and acceleration arrive without help from net saving or consumption restraint.
+Data from national accounts show no effect of change in net saving, in ratio to market-value capital, on change in growth rate of market-value capital (capital acceleration).
+Thus it appears that capital growth and acceleration arrive without help from net saving.
 We explore ways in which this is possible, and discuss implications for economic teaching and public policy.
@@
 Net saving, if so, raises the physical quantity of capital, but not the aggregate value, and so reduces the value per unit.
-The data are most easily explained by the present value principle, and by production efficiencies enabled through innovation.
+The data are most easily explained by the present value principle, by production efficiencies enabled through innovation, and by investment in the new technology from depreciation of the technology replaced.
 Value is created in the mind of the market at the moment when prospective cash flows are discounted.
@@
 That is to say that capital growth, in the mind of the market, precedes and pays for any needed investment in new plant and equipment.
+[^2] Present value is time-discounted, and affords innovators only enough, at first, for first steps in implementation.
+As implementation proceeds, present value rises and affords more.
 Meanwhile the innovator acquires materials and plant capacity and labor skills at market prices determined by their uses in current technology, but applies them more productively until competition catches up.
@@
 $g(K)=s^{*}\ ,\quad\text{uta,}\quad\text{where }s^{*}=\frac{S_{net}}{K}.\quad\text{ Then}$ (2)
-[…]
+$\Delta g(K)=\Delta s^{*}\ ,\quad\text{uta.}$. (3)
 $\Delta g(K)$ and $\Delta s^{*}$ respectively will be called "capital acceleration" and "thrift change," either of which may be positive or negative.
@@
 (3) and (4) on the reasoning that causal relationships among variables, in this case between thrift and capital growth rate, tend to be revealed most reliably in their concurrent changes.
-[^3] The ratio $\frac{\Delta A}{\Delta B}$, that is, compares changes in A to concurrent changes in B, and so suggests a causal relationship between A and B insofar as that ratio holds constant.
-This is the principle on which the concepts of correlation, covariance and regression are founded.
 Eq.
@@
 $\theta=1\ ,\quad\text{uta,}\quad\text{where}\ \theta=\frac{\Delta s^{*}}{\Delta g(K)}\ .$ (5)
-Test results for the predictions $\frac{s^{*}}{g(K)}=\frac{\Delta s^{*}}{\Delta g(K)}=1$, under thrift assumptions, appear in Fig.
-1 and in Tables 1 and 2.
+Test results for the predictions $\frac{s^{*}}{g(K)}=\theta=1$, under thrift assumptions, appear in Fig.
+1 and in Tables 1, 2 and 3.
+Screens to which they refer are explained in Section 13 below.
 Results were found as GDP-weighted averages over all countries and years, from data for net saving and market-value capital taken from national accounts.
 They show $\frac{s^{*}}{g(K)}=0.481$, with regression of $s^{*}$ on $g(K)$ found at 0.0771, and $\frac{\Delta s^{*}}{\Delta g(K)}=0.064$, with regression of $\Delta s^{*}$ on $\Delta g(K)$ at 0.0559.
+Table 1: Regression of $s^{*}$ on $g(K)$ and $\Delta s^{*}$ on $\Delta g(K)$, GDP-weighted. Screen = 0.01. $H_{0}$ per thrift theory: $\frac{s^{*}}{g(K)}=\frac{\Delta s^{*}}{\Delta g(K)}=1$.
+  Regression of $s^{*}$ on $g(K)$ 0.0771∗∗∗ (0.0144) Observations 1,826 R2 0.84511 Within R2 0.10669 Year fixed effects $\checkmark$ Country fixed effects $\checkmark$ Regression of $\Delta s^{*}$ on $\Delta g(K)$ 0.0559∗∗∗ (0.0057) Observations 1,574 R2 0.37071 Within R2 0.27233 Year fixed effects $\checkmark$ Country fixed effects $\checkmark$
 Fig. 1: Average $\frac{s^{*}}{g(K)}$ with LOESS smoothing over all countries, GDP-weighted, 1980-2022
-Table 1: Regression of $s^{*}$ on $g(K)$ and $\Delta s^{*}$ on $\Delta g(K)$, GDP-weighted. Screen = 0.01. $H_{0}$ per thrift theory: $\frac{s^{*}}{g(K)}=\frac{\Delta s^{*}}{\Delta g(K)}=1$.
-Regression of $s^{*}$ on $g(K)$ 0.0771∗∗∗
-Observations 1,826
-R2 0.84511
-Within R2 0.10669
-Year fixed effects $\checkmark$
-Country fixed effects $\checkmark$
-Regression of $\Delta s^{*}$ on $\Delta g(K)$ 0.0559∗∗∗
-Observations 1,574
-R2 0.37071
-Within R2 0.27233
-Year fixed effects $\checkmark$
-Country fixed effects $\checkmark$
+Fig. 2: Average $\theta_{s}^{*}$ with LOESS smoothing over all countries, GDP-weighted, 1980-2022
 Table 2: Average $\frac{s^{*}}{g(K)}$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()
   Armenia 1996 - 2020 (23) -0.63 Aruba 1996 - 2001 (6) 1.25 Australia 1980 - 2018 (33) 0.24 Austria 1996 - 2021 (23) 0.83 Azerbaijan 1996 - 2020 (23) 1.57 Bahrain 2009 - 2013 (5) -2.43 Belgium 1996 - 2021 (21) 0.61 Bolivia 1998 - 2015 (17) 0.40 Botswana 1996 - 1999 (4) 3.54 Brazil 1996 - 2019 (21) 0.25 British Virgin Islands 1996 - 1999 (4) 2.45 Bulgaria 1996 - 2016 (16) 0.14 Burkina Faso 2000 - 2019 (20) 0.23 Cameroon 1996 - 2019 (24) 1.10 Canada 1980 - 2022 (38) 0.37 Cape Verde 2008 - 2017 (9) 0.53 Chile 1997 - 2021 (19) 0.84 China 1992 - 2020 (29) 0.81 Colombia 1996 - 2022 (27) 0.60 Costa Rica 2013 - 2020 (6) 0.21 Croatia 1996 - 2021 (24) 0.43 Curaçao 2001 - 2016 (14) 1.26 Cyprus 1996 - 2021 (25) 0.07 Czechia 1994 - 2020 (20) 0.17 Côte d’Ivoire 1996 - 2000 (5) 0.29 Denmark 1996 - 2022 (26) 0.22 Dominican Republic 1996 - 2016 (10) 1.52 Ecuador 2008 - 2020 (13) 0.53 Egypt 1997 - 2015 (19) 0.92 El Salvador 2015 - 2019 (5) -0.62 Estonia 1996 - 2021 (24) 0.45 Finland 1996 - 2021 (23) 0.17 France 1980 - 2021 (35) 0.26 Germany 1980 - 2022 (38) 0.70 Greece 1996 - 2021 (26) 0.17 Guatemala 2002 - 2021 (20) -0.72 Guinea 2004 - 2010 (5) 0.72 Honduras 2001 - 2015 (14) -0.57 Hong Kong 1997 - 2011 (15) 0.90 Hungary 1996 - 2021 (21) 0.14 Iceland 2001 - 2014 (14) 0.22 India 1999 - 2019 (21) 0.57 Indonesia 2017 - 2019 (3) 0.94 Iran 1997 - 2017 (20) 4.96 Ireland 1996 - 2021 (22) 0.26 Israel 1996 - 2019 (24) 0.46 Italy 1980 - 2022 (36) 0.13 Japan 1980 - 2021 (29) 0.19 Kazakhstan 1997 - 2022 (25) 0.60 Kuwait 2003 - 2017 (14) 1.11 Kyrgyzstan 1996 - 2021 (24) 0.46 Latvia 1996 - 2021 (24) -0.23 Lithuania 1996 - 2021 (24) 0.13 Luxembourg 1996 - 2021 (22) 0.03 Malaysia 2007 - 2015 (9) 2.52 Malta 1996 - 2021 (23) 0.16 Mauritius 2014 - 2018 (5) 0.04 Mexico 1996 - 2021 (22) 0.15 Moldova 1996 - 2019 (22) -0.91 Mongolia 2006 - 2020 (15) 0.10 Morocco 1999 - 2021 (23) 1.13 Netherlands 1996 - 2021 (23) 0.51 New Zealand 1996 - 2019 (21) 0.78 Nicaragua 2007 - 2018 (12) 0.08 Niger 1996 - 2019 (24) 0.90 Norway 1981 - 2021 (38) 0.95 Peru 2008 - 2021 (14) 0.76 Philippines 1996 - 2022 (27) 1.76 Poland 1996 - 2021 (25) 0.71 Portugal 1996 - 2022 (24) 0.00 Qatar 2002 - 2017 (14) 1.89 Romania 1996 - 2020 (22) 0.03 Russia 1996 - 2017 (13) -0.19 Saudi Arabia 2003 - 2009 (7) 2.65 Senegal 2015 - 2021 (7) 0.24 Serbia 1998 - 2021 (23) 0.61 Slovakia 1996 - 2022 (24) 0.15 Slovenia 1996 - 2021 (20) 0.28 South Africa 1996 - 2022 (24) 0.28 South Korea 1996 - 2020 (24) 0.63 Spain 1995 - 2021 (25) 0.32 Sweden 1980 - 2021 (37) 0.41 Switzerland 1993 - 2021 (25) 0.69 Tunisia 1996 - 2011 (16) 0.29 Turkey 2010 - 2017 (8) 1.32 USA 1980 - 2021 (38) 0.31 Ukraine 1996 - 2019 (21) 0.17 United Kingdom 1981 - 2021 (36) 0.16 Uruguay 2016 - 2017 (2) 1.16 Uzbekistan 2011 - 2021 (10) 0.21 Vanuatu 2002 - 2007 (6) 0.29 Venezuela 1998 - 2019 (20) 0.55
-Note: Thrift theory predicts $\frac{s^{*}}{g(K)}=1$. Free growth theory makes no prediction for these data.
-Table 3: Average $\theta_{s}^{*}$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()
+Note: $s^{*}$ is defined as $\frac{S_{net}}{K}$. Thrift theory predicts $\frac{s^{*}}{g(K)}=1$. Free growth theory makes no prediction for these data.
+Table 3: Average $\theta$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()
   Armenia 1997 - 2020 (20) 0.12 Aruba 1997 - 2001 (5) 1.22 Australia 1974 - 2019 (35) 0.03 Austria 1997 - 2021 (16) 0.10 Azerbaijan 1997 - 2020 (23) 0.75 Bahrain 2010 - 2013 (4) 0.41 Belgium 1997 - 2021 (19) 0.05 Bolivia 1999 - 2014 (12) 0.04 Botswana 1997 - 2000 (4) 0.24 Brazil 1998 - 2020 (22) 0.10 British Virgin Islands 1997 - 1998 (2) 2.24 Bulgaria 1997 - 2017 (15) -0.03 Burkina Faso 2001 - 2019 (16) 0.04 Cameroon 1997 - 2019 (19) 0.18 Canada 1974 - 2022 (41) 0.09 Cape Verde 2009 - 2016 (8) 0.27 Chile 1998 - 2021 (22) 0.06 China 1981 - 2020 (33) 0.07 Colombia 1997 - 2022 (18) -0.03 Costa Rica 2014 - 2020 (6) 0.04 Croatia 1997 - 2021 (22) 0.19 Curaçao 2002 - 2016 (14) 0.41 Cyprus 1997 - 2021 (23) -0.06 Czechia 1995 - 2021 (19) -0.03 Côte d’Ivoire 1997 - 2000 (4) -0.03 Denmark 1997 - 2022 (22) 0.05 Dominican Republic 2007 - 2015 (9) 0.15 Ecuador 2009 - 2020 (12) 0.17 Egypt 1998 - 2015 (16) 0.13 El Salvador 2016 - 2019 (4) 0.98 Estonia 1997 - 2021 (19) 0.06 Finland 1997 - 2022 (21) -0.02 France 1972 - 2021 (33) 0.09 Germany 1972 - 2022 (31) 0.05 Greece 1997 - 2021 (24) 0.24 Guatemala 2003 - 2021 (11) 0.15 Guinea 2005 - 2010 (6) 0.08 Honduras 2003 - 2015 (10) 0.40 Hong Kong 1997 - 2011 (13) 0.04 Hungary 1997 - 2021 (21) 0.07 Iceland 2002 - 2014 (12) 0.10 India 1997 - 2019 (16) 0.14 Indonesia 2019 - 2019 (1) -0.41 Iran 1998 - 2018 (19) 1.09 Ireland 1997 - 2021 (23) 0.02 Israel 1997 - 2017 (18) 0.07 Italy 1972 - 2021 (29) 0.05 Japan 1972 - 2021 (38) 0.05 Kazakhstan 1997 - 2022 (23) 0.08 Kuwait 2005 - 2017 (11) 0.04 Kyrgyzstan 1998 - 2021 (22) 0.23 Latvia 1998 - 2021 (20) 0.01 Lithuania 1997 - 2021 (18) 0.10 Luxembourg 1997 - 2021 (25) 0.00 Malaysia 2008 - 2015 (7) 0.01 Malta 1997 - 2021 (23) 0.08 Mauritius 2015 - 2018 (4) 0.02 Mexico 1997 - 2021 (22) 0.06 Moldova 1997 - 2019 (21) 0.11 Mongolia 2008 - 2021 (11) 0.10 Morocco 2000 - 2021 (17) 0.19 Netherlands 1997 - 2021 (20) 0.03 New Zealand 1997 - 2019 (18) 0.14 Nicaragua 2008 - 2019 (12) 0.06 Niger 1997 - 2019 (19) 0.22 Norway 1982 - 2021 (34) 0.12 Peru 2009 - 2021 (11) 0.04 Philippines 1997 - 2022 (20) 0.37 Poland 1997 - 2021 (15) 0.20 Portugal 1997 - 2022 (17) 0.02 Qatar 2004 - 2018 (12) 0.29 Romania 1997 - 2020 (14) 0.02 Russia 1997 - 2018 (14) 0.09 Saudi Arabia 2005 - 2009 (3) 2.23 Senegal 2018 - 2021 (3) 0.19 Serbia 1999 - 2021 (22) 0.22 Slovakia 1997 - 2021 (18) 0.09 Slovenia 1997 - 2021 (20) 0.10 South Africa 1997 - 2022 (21) 0.10 South Korea 1997 - 2020 (15) 0.04 Spain 1972 - 2021 (42) 0.06 Sweden 1974 - 2022 (41) 0.02 Switzerland 1993 - 2020 (24) 0.08 Tunisia 1997 - 2011 (12) 0.32 Turkey 2011 - 2015 (5) -0.16 USA 1972 - 2021 (44) 0.02 Ukraine 1997 - 2020 (23) 0.05 United Kingdom 1972 - 2021 (42) 0.02 Uruguay 2017 - 2017 (1) 0.56 Uzbekistan 2012 - 2021 (9) 0.05 Vanuatu 2003 - 2007 (3) 0.15 Venezuela 1999 - 2019 (20) 0.31
-Note: Thrift theory predicts $\theta_{s}^{*}=1$.
-Fig. 2: Average $\theta_{s}^{*}$ with LOESS smoothing over all countries, GDP-weighted, 1980-2022
+Note: $\theta$ is defined as $\frac{\Delta s^{*}}{\Delta g(K)}$. Thrift theory predicts $\theta=1$.
 3 Interpretation of test results
@@
 Causal relatedness is shown rather by the degree of constancy in proportions, and thus on regression of $s^{*}$ on $g(K)$ found at 0.0771.
-$\theta_{s}$ or $\frac{\Delta s^{*}}{\Delta g(K)}$, measured at $\theta_{s}^{*}=0.064$, gives a separate measure of causal relatedness between $s^{*}$ and $g(K)$ as explained.
+$\theta$ or $\frac{\Delta s^{*}}{\Delta g(K)}$, measured at $0.064$, gives a separate measure of causal relatedness between $s^{*}$ and $g(K)$ as explained.
 Regression of $\Delta s^{*}$ on $\Delta g(K)$, found at 0.0559, gives a measure of causal relatedness between $\Delta s^{*}$ and $\Delta g(K)$.
@@
 This interpretation could fit the appearance in Fig.
-2 that $\frac{\Delta s^{*}}{\Delta g(K)}$, shown as $\theta^{*}_{s}$, found higher values during the period of the dot com and subprime crises.
+2 that $\theta$ found higher values during the period of the Iraq-Kuwait, dot com and subprime crises.
 In general, it might be expected that households stabilize consumption by building up reserves through net saving when earnings are higher, and deplete them when earnings are lower, as described in the life cycle theory of saving (Modigliani (1954)) or the permanent income hypothesis (Friedman (1957)).
-Free growth theory explains the rise in earnings as caused by a rise in productivity of current plant and implements and skills, and so predicts the same effects on regressions and $\frac{\Delta s^{*}}{\Delta g(K)}$ as does thrift theory, but from an opposite causality.
+Free growth theory explains the rise in earnings as caused by a rise in productivity of current plant and implements and skills, and so predicts the same effects on regressions and $\theta$ as does thrift theory, but from an opposite causality.
 Although these results cannot exclude the possibility that some net saving has effect, data are easiest to reconcile with free growth theory, where all capital growth is explained by productivity gain.
 Pending further study, we will proceed on this interpretation.
+4 Inferences for the creative destruction idea
+Creative destruction (Schumpeter (1912) [^3] Published by Schumpeter in 1912, and Translated by Opie in 1934.) is obsolescence of existing capital through innovation of new and preferred technology, as with horse-drawn carriages and automobiles.
+Granting the obsolescence of old ways by new, it does not follow that innovation reduces values of assets already present.
+Data cited in this paper appear to show that capital growth is financed wholly by depreciation investment.
+A natural source of that investment is depreciation of assets the new technology will replace, and a natural market for the new technology may be owners of the assets to be replaced.
+Carriage owners were better able to afford cars when the carriages in which they had invested before lived out their economic lives as originally expected, so that the depreciation component in the cash flows originally expected can pay for cars in replacement just as it might have paid for new carriages.
+This analysis suggests that market forces favor a kind through orderly obsolescence through replacement without premature displacement as the new technology awaits its turn.
+It also shows that the new technology as a separate entity grows by net investment financed from disinvestment from the technology replaced.
+It is only at the scale of the closed economy, meaning of all assets together, that net saving and investment appear to be unproductive.
 5 Mechanics of free growth
@@
 Net output means value added.
-Appendix B will show an argument that to equate value added to the sum of consumption and capital growth, and neither less nor more, would miss two components, one negative and one positive, in value added to human capital, and that both of these components are unmeasurable by known means.
-Since we work here from measurable data for consumption and net saving and market value capital, we did our best to reason from those three quantities alone.
+Appendix B will show an argument that to interpret value added as the sum of consumption and capital growth, and neither less nor more, would miss two components, one negative (human depreciation) and one positive (self-invested work), in value added to human capital, and that both of these components are unmeasurable by known means.
+Since we work here from measurable data for net saving and market value capital, we did our best to reason from those two quantities alone.
 11 Data sources
@@
 13 Displays
-The web appendix includes displays of $\theta_{s}^{*}$ for each year in each country over the report period.
-These tend to show upward and downward spikes in values of $\theta_{s}^{*}$ in some years.
+The web appendix includes displays of $\theta$ for each year in each country over the report period.
+These tend to show upward and downward spikes in values of $\theta$ in some years.
 Those spikes tend to be associated with small absolute values of denominators, in these cases $\Delta g(K)$, in those countries and years.
@@
 To maximize reliability of test results, we apply a range of screens to omit years where absolute denominators fall below a given threshold.
-Some displays show data for all years, regardless of denominator size.
+Some displays in the web appendix show data for all years, regardless of denominator size.
 Others screen out all years where absolute denominators are less then 0.01, then 0.025, then continuing upward in increments of 0.025 to a maximum screen of 0.15.
-$\theta_{s}^{*}$ is plotted for each country unscreened and at each of the seven successive levels of screening.
-All displays applied a screen of 0.01.
+$\theta$ is plotted for each country unscreened and at each of the seven successive levels of screening.
+All displays in the main body of this paper apply a screen of 0.01.
 The denominator whose absolute value is screened is capital acceleration $\Delta g(K)$ or capital growth rate $g(K)$ in all displays.
@@
 14 Disclaimers
-We accept that capital growth is impossible without capital replacement first, to make up losses to depreciation, and that the only source of capital replacement is replacement investment.
+We accept that capital growth is impossible without capital replacement first, to make up losses to depreciation, and that the expected source of capital replacement is replacement investment.
 Thus we accept the necessity and efficacy of replacement investment.
@@
 Appendix A.   Equivalence of saving and investment
-Any usage which treats saving and investment as equal must deal with the fact that saving held in cash is not investment in the usual sense of spending on productive capital, and contributes nothing to output.
+Any usage which treats saving and investment as equal must deal with the fact that saving held in cash is not investment in the usual sense of spending on new productive capital, and contributes nothing to output.
 Keynes (1936) defined that unspent saving as intended saving, and actual saving as the part so spent.
@@
 Thus investment could have the usual meaning in usual times, when prospective returns bring animal spirits, and could include saving in cash when not.
-It is in this sense that we equate saving to investment.
+It is in this sense, differing from Keynes by equating intended (cash) investment to actual investment when firms and households see nothing better, that we equate saving to investment.
 Appendix B.  Net output with human capital

### v10

@@
 It is only at the scale of the closed economy, meaning of all assets together, that net saving and investment appear to be unproductive.
-5 Mechanics of free growth
+5 Solow’s puzzle
 Some growth is capital widening, where structures and implements increase in number but do not change in design.
@@
 But Solow, who came to conclusions similar to ours from different evidence, puzzled as to how capital growth without net saving could be possible for capital deepening through “embodied” growth, where products of new design are made from plant of new design.
-[^4] The terms capital deepening, capital widening, embodied growth and disembodied growth are all Solow’s.
-The solution, we suggest, is that embodied growth is disembodied growth on a finer scale.
+Solow’s Nobel Prize acceptance speech in mentions the embodiment problem: [^4] Solow (1988). The terms capital deepening, capital widening, embodied growth and disembodied growth are all Solow’s.
+"… much technological progress, maybe most of it, could find its way into actual production only with the use of new and different capital equipment.
+Therefore the effectiveness of innovation in increasing output would be paced by the rate of gross investment.
+A policy to increase investment would therefore lead to … a faster transfer of new technology into actual production … By 1958 I was able to produce a model that allowed for the embodiment effect.
+If common sense was right, the embodiment model should have fit the facts better than the earlier one.
+But it did not… I do not know if that finding should be described as a paradox, but it was at least a puzzle."
+Solow was right in that much innovation requires months or years of gross investment to pay for negative cash flows in building new plant before the new products roll out from the factory doors.
+What he left unsaid is that this investment would not have been forthcoming unless positive cash flows at the end of the tunnel were predicted to be sufficient to create positive present value from the start.
+Data in our displays suggest that this initial creation of value, plus depreciation investment, are enough.
+6 Mechanics of free growth
+The present value argument just given offers an accounting explanation of embodied growth in terms of sources of funds.
+The step-by-step mechanical explanation, we suggest, in terms of things rather than dollars, is that embodied growth is disembodied growth on a finer scale.
 At each step toward realization of the new plant and products, raw materials and products and labor skills and plant capacity currently available on the market are adapted to new uses.
@@
 The present value of yields from this advantage in return, or equivalently the innovator’s reserve, defines the non-random component in free growth.
+Solow’s puzzle runs deep.
+How can thoughts be things?
+How can assessments of present value pay, at least in part, for their own realization?
+Science and philosophy push our minds into areas where they are not at home.
+Neither relativity nor quantum mechanics, which explore the scale of great speeds and minute distances, sink in easily.
+Economics, too, invites surprises as it looks farther from the scale of the household and community and firm to which we are accustomed.
+Free growth theory predicts only at the universal scale, and accepts that thrift explains a greater part of growth at progressively smaller scales.
+Our findings, with that allowance, are awkward for much of economics as generally taught.
+We encourage others to check our data and methods and inferences, and improve on them wherever they can.
 Appendix A.   Equivalence of saving and investment

### v11

@@
 Economists since the early nineteenth century have added the proviso that net saving cannot safely outpace innovation; more capital must mean capital redesigned for greater productivity if economies are to escape risk of capital glut and diminishing returns (West (1815), Ricardo (1815), Malthus (1815)).
-Roy Harrod (1939) described that limit for safe net saving, meaning the rate of imagining and developing new ideas for more productive forms of capital, as the “warranted rate”.
+Roy Harrod (1939)) described that limit for safe net saving, meaning the rate of imagining and developing new ideas for more productive forms of capital, as the “warranted rate”.
 Harrod, and many other economists of his time and since, have focused on growth of output rather than of capital, but have modeled growth of output by first assuming the equivalence of net saving and capital growth, within the warranted rate, and then looking for effects of that capital growth on later output growth.
@@
 Those test results from data for net saving/capital ratio are easiest to reconcile in an interpretation that net saving has no effect on capital growth.
+[^2] Assuming that gross saving is enough to offset depreciation (see Section 3).
+We will assume this coverage unless stated otherwise.
 Net saving, if so, raises the physical quantity of capital, but not the aggregate value, and so reduces the value per unit.
-The data are most easily explained by the present value principle, by production efficiencies enabled through innovation, and by investment in the new technology from depreciation of the technology replaced.
+The data are most easily explained by the present value principle, and by production efficiencies enabled through innovation.
 Value is created in the mind of the market at the moment when prospective cash flows are discounted.
@@
 As implementation proceeds, present value rises and affords more.
+Early investors borrow from the future, in effect, and pay the interest from increments in present value as positive cash flows approach with time.
 Meanwhile the innovator acquires materials and plant capacity and labor skills at market prices determined by their uses in current technology, but applies them more productively until competition catches up.
-It is that temporary market advantage to the innovator which explains capital growth without net saving in a practical and mechanical sense, while the present value principle gives the explanation in terms of market valuation.
+It is that temporary market advantage to the innovator which explains capital growth without net saving in a practical and mechanical sense, while the present value principle gives the source of investment.
 This idea will be called “free growth theory” for easy reference.
 It predicts only at the largest scales, and only for the private sector.
-Individuals and groups and even small economies can grow or decline through net saving or dissaving.
+Individuals and groups and even small economies can grow in capital through saving if others dissave equally.
 That possibility is foreclosed only at the scale of all capital and all economies together.
@@
 Meanwhile the doctrine that net income equals consumption plus net saving is put into question by evidence offered here suggesting that net saving increases the physical quantity of capital, but not the aggregate value.
-In general, economics might consider relying less on book value and net investment, and more on market value and on the power of ideas.
 2 Net saving, capital growth and capital acceleration
@@
 $g(K)=s^{*}\ ,\quad\text{uta,}\quad\text{where }s^{*}=\frac{S_{net}}{K}.\quad\text{ Then}$ (2)
-$\Delta g(K)=\Delta s^{*}\ ,\quad\text{uta.}$. (3)
-$\Delta g(K)$ and $\Delta s^{*}$ respectively will be called "capital acceleration" and "thrift change," either of which may be positive or negative.
+$\Delta g(K)=\Delta s^{*}\ ,\quad\text{uta.}$ (3)
+$\Delta g(K)$ and $\Delta s^{*}$ respectively will be called "capital acceleration" and "thrift change," either of which may be positive or negative or zero.
 Division of Eq.
@@
 Results were found as GDP-weighted averages over all countries and years, from data for net saving and market-value capital taken from national accounts.
-They show $\frac{s^{*}}{g(K)}=0.481$, with regression of $s^{*}$ on $g(K)$ found at 0.0771, and $\frac{\Delta s^{*}}{\Delta g(K)}=0.064$, with regression of $\Delta s^{*}$ on $\Delta g(K)$ at 0.0559.
+Figures 1 and 2 show $\frac{s^{*}}{g(K)}$ and $\frac{\Delta s^{*}}{\Delta g(K)}$ as 0.481 and 0.064 respectively.
+Table 1 shows regression of $s^{*}$ on $g(K)$ as 0.0771, and regression of $\Delta s^{*}$ on $\Delta g(K)$ as 0.0559.
 Table 1: Regression of $s^{*}$ on $g(K)$ and $\Delta s^{*}$ on $\Delta g(K)$, GDP-weighted. Screen = 0.01. $H_{0}$ per thrift theory: $\frac{s^{*}}{g(K)}=\frac{\Delta s^{*}}{\Delta g(K)}=1$.
-  Regression of $s^{*}$ on $g(K)$ 0.0771∗∗∗ (0.0144) Observations 1,826 R2 0.84511 Within R2 0.10669 Year fixed effects $\checkmark$ Country fixed effects $\checkmark$ Regression of $\Delta s^{*}$ on $\Delta g(K)$ 0.0559∗∗∗ (0.0057) Observations 1,574 R2 0.37071 Within R2 0.27233 Year fixed effects $\checkmark$ Country fixed effects $\checkmark$
+Regression of $s^{*}$ on $g(K)$ 0.0771∗∗∗
+Observations 1,826
+R2 0.84511
+Within R2 0.10669
+Year fixed effects $\checkmark$
+Country fixed effects $\checkmark$
+Regression of $\Delta s^{*}$ on $\Delta g(K)$ 0.0559∗∗∗
+Observations 1,574
+R2 0.37071
+Within R2 0.27233
+Year fixed effects $\checkmark$
+Country fixed effects $\checkmark$
 Fig. 1: Average $\frac{s^{*}}{g(K)}$ with LOESS smoothing over all countries, GDP-weighted, 1980-2022
@@
 Armenia 1997 - 2020 (20) 0.12 Aruba 1997 - 2001 (5) 1.22 Australia 1974 - 2019 (35) 0.03 Austria 1997 - 2021 (16) 0.10 Azerbaijan 1997 - 2020 (23) 0.75 Bahrain 2010 - 2013 (4) 0.41 Belgium 1997 - 2021 (19) 0.05 Bolivia 1999 - 2014 (12) 0.04 Botswana 1997 - 2000 (4) 0.24 Brazil 1998 - 2020 (22) 0.10 British Virgin Islands 1997 - 1998 (2) 2.24 Bulgaria 1997 - 2017 (15) -0.03 Burkina Faso 2001 - 2019 (16) 0.04 Cameroon 1997 - 2019 (19) 0.18 Canada 1974 - 2022 (41) 0.09 Cape Verde 2009 - 2016 (8) 0.27 Chile 1998 - 2021 (22) 0.06 China 1981 - 2020 (33) 0.07 Colombia 1997 - 2022 (18) -0.03 Costa Rica 2014 - 2020 (6) 0.04 Croatia 1997 - 2021 (22) 0.19 Curaçao 2002 - 2016 (14) 0.41 Cyprus 1997 - 2021 (23) -0.06 Czechia 1995 - 2021 (19) -0.03 Côte d’Ivoire 1997 - 2000 (4) -0.03 Denmark 1997 - 2022 (22) 0.05 Dominican Republic 2007 - 2015 (9) 0.15 Ecuador 2009 - 2020 (12) 0.17 Egypt 1998 - 2015 (16) 0.13 El Salvador 2016 - 2019 (4) 0.98 Estonia 1997 - 2021 (19) 0.06 Finland 1997 - 2022 (21) -0.02 France 1972 - 2021 (33) 0.09 Germany 1972 - 2022 (31) 0.05 Greece 1997 - 2021 (24) 0.24 Guatemala 2003 - 2021 (11) 0.15 Guinea 2005 - 2010 (6) 0.08 Honduras 2003 - 2015 (10) 0.40 Hong Kong 1997 - 2011 (13) 0.04 Hungary 1997 - 2021 (21) 0.07 Iceland 2002 - 2014 (12) 0.10 India 1997 - 2019 (16) 0.14 Indonesia 2019 - 2019 (1) -0.41 Iran 1998 - 2018 (19) 1.09 Ireland 1997 - 2021 (23) 0.02 Israel 1997 - 2017 (18) 0.07 Italy 1972 - 2021 (29) 0.05 Japan 1972 - 2021 (38) 0.05 Kazakhstan 1997 - 2022 (23) 0.08 Kuwait 2005 - 2017 (11) 0.04 Kyrgyzstan 1998 - 2021 (22) 0.23 Latvia 1998 - 2021 (20) 0.01 Lithuania 1997 - 2021 (18) 0.10 Luxembourg 1997 - 2021 (25) 0.00 Malaysia 2008 - 2015 (7) 0.01 Malta 1997 - 2021 (23) 0.08 Mauritius 2015 - 2018 (4) 0.02 Mexico 1997 - 2021 (22) 0.06 Moldova 1997 - 2019 (21) 0.11 Mongolia 2008 - 2021 (11) 0.10 Morocco 2000 - 2021 (17) 0.19 Netherlands 1997 - 2021 (20) 0.03 New Zealand 1997 - 2019 (18) 0.14 Nicaragua 2008 - 2019 (12) 0.06 Niger 1997 - 2019 (19) 0.22 Norway 1982 - 2021 (34) 0.12 Peru 2009 - 2021 (11) 0.04 Philippines 1997 - 2022 (20) 0.37 Poland 1997 - 2021 (15) 0.20 Portugal 1997 - 2022 (17) 0.02 Qatar 2004 - 2018 (12) 0.29 Romania 1997 - 2020 (14) 0.02 Russia 1997 - 2018 (14) 0.09 Saudi Arabia 2005 - 2009 (3) 2.23 Senegal 2018 - 2021 (3) 0.19 Serbia 1999 - 2021 (22) 0.22 Slovakia 1997 - 2021 (18) 0.09 Slovenia 1997 - 2021 (20) 0.10 South Africa 1997 - 2022 (21) 0.10 South Korea 1997 - 2020 (15) 0.04 Spain 1972 - 2021 (42) 0.06 Sweden 1974 - 2022 (41) 0.02 Switzerland 1993 - 2020 (24) 0.08 Tunisia 1997 - 2011 (12) 0.32 Turkey 2011 - 2015 (5) -0.16 USA 1972 - 2021 (44) 0.02 Ukraine 1997 - 2020 (23) 0.05 United Kingdom 1972 - 2021 (42) 0.02 Uruguay 2017 - 2017 (1) 0.56 Uzbekistan 2012 - 2021 (9) 0.05 Vanuatu 2003 - 2007 (3) 0.15 Venezuela 1999 - 2019 (20) 0.31
-Note: $\theta$ is defined as $\frac{\Delta s^{*}}{\Delta g(K)}$. Thrift theory predicts $\theta=1$.
+Note: $\theta$ is defined as $\frac{\Delta s^{*}}{\Delta g(K)}$. Thrift theory predicts $\theta=1$. See Section 3 for predictions of free growth theory.
 3 Interpretation of test results
@@
 Might these small but positive findings allow the possibility that some capital growth is explained by net saving?
-Any explanation must account for the chance that replacement saving will sometimes not be enough to offset depreciation.
-This shortfall is likeliest in prolonged downturns, where loss of income motivates households to forego net saving, and then to invade capital by putting off maintenance expense and replacement investment in order to protect consumption.
-Effects will include a positive effect on regression of $s^{*}$ on $g(K)$, and an upward effect on $\frac{\Delta s^{*}}{\Delta g(K)}$, as $s^{*}$ and $g(K)$ decline together in response to downward pressure on income.
-This interpretation could fit the appearance in Fig.
+Free growth theory predicts that net saving contributes nothing to capital growth, given that gross investment offsets depreciation.
+In years of recession, households may defer part or all of depreciation saving to support consumption.
+In these years, growth in the present value of better ideas will not realize equal net growth of capital, as some of it must offset any shortfall in depreciation saving first.
+Allowing for these years, we can say more compactly that free growth theory predicts that a gain in the present value of new ideas accounts for all capital creation, and thus for the sum of capital growth and restoration of shortfalls in depreciation investment.
+Thus free growth theory does not predict a zero value for $\theta$, but rather predicts $\theta$ to equal the share of capital creation applied to restoring those shortfalls.
+Insofar as 0.064 gives this share, free growth theory could fit the appearance in Fig.
 2 that $\theta$ found higher values during the period of the Iraq-Kuwait, dot com and subprime crises.
-In general, it might be expected that households stabilize consumption by building up reserves through net saving when earnings are higher, and deplete them when earnings are lower, as described in the life cycle theory of saving (Modigliani (1954)) or the permanent income hypothesis (Friedman (1957)).
-Free growth theory explains the rise in earnings as caused by a rise in productivity of current plant and implements and skills, and so predicts the same effects on regressions and $\theta$ as does thrift theory, but from an opposite causality.
-Although these results cannot exclude the possibility that some net saving has effect, data are easiest to reconcile with free growth theory, where all capital growth is explained by productivity gain.
-Pending further study, we will proceed on this interpretation.
+Pending clarification of that share, we will proceed on the simplifying assumption that free growth theory predicts correctly.
 4 Inferences for the creative destruction idea
@@
 Capital widening, however, is practical only so far before glut and diminishing returns set in.
-Further growth from that point must come from capital deepening, meaning improvements in the design of capital.
+Further growth must come from capital deepening, meaning improvements in the design of capital.
 Solow (1956) noted a kind of middle ground between capital widening and capital deepening in the disembodied growth mentioned earlier; ships carrying coals to Newcastle can raise prospective cash flows, and hence present value, by reversing the business plan.
 But Solow, who came to conclusions similar to ours from different evidence, puzzled as to how capital growth without net saving could be possible for capital deepening through “embodied” growth, where products of new design are made from plant of new design.
-Solow’s Nobel Prize acceptance speech in mentions the embodiment problem: [^4] Solow (1988). The terms capital deepening, capital widening, embodied growth and disembodied growth are all Solow’s.
+Solow received the Nobel Prize in 1987.
+His acceptance speech mentions the embodiment problem: [^5] Solow (1988).
+The terms capital deepening, capital widening, embodied growth and disembodied growth are all Solow’s.
 "… much technological progress, maybe most of it, could find its way into actual production only with the use of new and different capital equipment.
@@
 The innovator pays for these inputs at a market price determined by their value in established productive uses, but applies them innovatively to realize higher prospective cash flows, and hence higher present values, to the innovator (Marshall (1890), Schumpeter and Opie (1934)).
-This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.555i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns.
-The present value of additional cash flow enabled by this advantage in return quantifies the innovator’s reserve and equivalently the non-random component of free growth.
+This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.666i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns.
+The present value of additional cash flow enabled by this advantage in return quantifies the innovator’s reserve, and equivalently the non-random component of free growth.
 The innovator’s reserve quantifies the part of free growth explained by productivity gain as distinct from random market noise.
@@
 Our findings support those of Piketty and Zucman (2014) and Kurz (2023) as to the market power of innovators to explain capital growth beyond net saving.
+Those studies assumed that net saving is recovered in capital growth, so that this market power, which we call the innovator’s reserve, need explain only the part of capital growth remaining.
 Again, we go farther by questioning the assumption that net saving contributes even a part of capital growth.
@@
 Net output means value added.
-Appendix B will show an argument that to interpret value added as the sum of consumption and capital growth, and neither less nor more, would miss two components, one negative (human depreciation) and one positive (self-invested work), in value added to human capital, and that both of these components are unmeasurable by known means.
+Appendix B will show an argument that to interpret value added as the sum of consumption and capital growth, and neither less nor more, would miss two components, one negative (human depreciation) and one positive (self-invested work), in value added to human capital.
+Both of these flows are unmeasurable by known means.
 Since we work here from measurable data for net saving and market value capital, we did our best to reason from those two quantities alone.
@@
 We dispute only the efficacy of net investment in realizing capital growth after replacement investment has made up for depreciation.
+Evidence supports the proposal of Modigliani (1954) that the young save to support retirement, and that retirees dissave to support consumption.
+[^8] Friedman (1957) follows a different argument to a similar conclusion.
+Such age-based differences are expected to offset in the stationary state, where tradition and free growth theory agree that overall net saving is zero.
+We accept that individual net saving can realize equal individual capital growth, so long as collective saving and dissaving by all individuals offset to zero.
 16 Discussion and conclusions
@@
 Appendix B. Net output with human capital
-Human capital is impractical to measure, as it leaves little market record other than for its rental income in pay and investment cost in schooling.
-Thus national accounts leave it implicit, and allow us to infer what we can from data for pay and schooling.
+Human capital is impractical to measure, as it leaves little market record other than for its rental income in pay and investment cost in nurture and schooling.
+Thus national accounts leave it implicit, and allow us to infer what we can from data for pay and nurture and schooling.
 Those accounts are founded on the principle, sound when terms are appropriately redefined, that net output, or value added, is expressed in the sum of capital growth and net outflow from the value-added chain.
 In national accounts, then, where physical capital is the whole of capital while net outflow of the chain is the whole of consumption, the reasoning is
-$Y=\Delta K+C\ ,\quad\text{neglecting human capital.}$. (B.1)
+$Y=\Delta K+C\ ,\quad\text{neglecting human capital.}$ (B.1)
 It is possible in principle to model a value-added chain which includes human capital, and to compare findings with those shown in Eq.
@@
 Adapting the classic illustration of the value added principle, say that farms produce wheat, mills convert the wheat to flour, bakeries convert the flour into bread, and humans convert some of the bread, called invested consumption, into human capital.
-The net outflow from this extended value-added chain is not all of consumption, but only pure consumption, meaning the part remaining after the part invested in human capital (invested consumption) is subtracted [^7] The concepts of invested consumption, pure consumption, self-invested work and human depreciation were introduced in Schultz (1961)..
+The net outflow from this extended value-added chain is not all of consumption, but only pure consumption, meaning the part remaining after the part invested in human capital (invested consumption) is subtracted.
+[^9] The concepts of invested consumption, pure consumption, self-invested work and human depreciation were introduced in Schultz (1961).
 By this reasoning, the principle that net output is expressed in capital growth plus net outflow gives
-$Y=\Delta K+\Delta H+C_{p}\ ,\quad\text{allowing human capital,}$, (B.2)
+$Y=\Delta K+\Delta H+C_{p}\ ,\quad\text{allowing human capital,}$ (B.2)
 where $C_{p}$ gives pure consumption.
@@
 Thus the combined arguments of Schultz and Ben-Porath arrive at
-$C=C_{s}+C_{p}\quad\text{and}\quad\Delta H=C_{s}+W_{s}-D(H)\ ,\quad\text{allowing human capital.}$. (B.3)
+$C=C_{s}+C_{p}\quad\text{and}\quad\Delta H=C_{s}+W_{s}-D(H)\ ,\quad\text{allowing human capital.}$ (B.3)
 Substitution of these equations into Eq. (B.2) finds
 $Y=\Delta K+C_{s}+W_{s}-D(H)+C_{p}\quad\text{and consequently}$
-$Y=\Delta K+C+W_{s}-D(H)\ ,\quad\text{allowing human capital,}$, (B.4)
+$Y=\Delta K+C+W_{s}-D(H)\ ,\quad\text{allowing human capital,}$ (B.4)
 if Schultz and Ben-Porath and the reasoning here are right.
@@
 This passage may be the first clear statement of what we call free growth theory.
-His use of the words "accelerate" and "concurrently" suggest that his path of reasoning was something like ours, and that he meant what we call free growth rather than the alternating phases of higher and lower saving rates pictured in thrift theory.
+Mill’s use of the words "accelerate" and "concurrently" suggest that his path of reasoning was something like ours, and that he meant what we call free growth rather than the alternating phases of higher and lower saving rates pictured in thrift theory.
 […]

### v12

@@
 Abstract
-Data from national accounts show no effect of change in net saving, in ratio to market-value capital, on change in growth rate of market-value capital (capital acceleration).
-Thus it appears that capital growth and acceleration arrive without help from net saving.
+Data from national accounts support an argument that capital growth and acceleration arrive without help from net saving.
 We explore ways in which this is possible, and discuss implications for economic teaching and public policy.
@@
 1 Introduction and overview
-Many economists over the centuries have reasoned that net saving, or equivalently net investment [^1] As reported in national accounts; they differ only by statistical discrepancy.
-Also see Appendix A., should tend to give equal capital growth.
+Many economists over the centuries have reasoned that net saving [^1] As reported in national accounts; saving and investment differ only by statistical discrepancy.
+We view them as different, for reasons shown in Appendix A, but will equate them in this paper by using the word ”saving” in the Keynesian sense of invested saving.
+should tend to give equal capital growth.
 Economists since the early nineteenth century have added the proviso that net saving cannot safely outpace innovation; more capital must mean capital redesigned for greater productivity if economies are to escape risk of capital glut and diminishing returns (West (1815), Ricardo (1815), Malthus (1815)).
@@
 Robert Solow (1957) allowed that possibility for “disembodied” growth, where plant and products already existing are repurposed or redeployed in more productive ways.
-We test between those two explanations of capital growth, by net saving or by increase in productivity of capital and labor already in existence, by comparing net saving to concurrent change in market-value capital in 92 countries.
+We test between those two explanations of capital growth, by net saving or by increase in productivity of capital and labor already in existence, by comparing net saving to concurrent change in market-value capital in 92 countries, and also comparing changes in those two variables.
 All data are drawn from national accounts of those countries as collated on the free website World Inequality Database.
-Those test results from data for net saving/capital ratio are easiest to reconcile in an interpretation that net saving has no effect on capital growth.
-[^2] Assuming that gross saving is enough to offset depreciation (see Section 3).
-We will assume this coverage unless stated otherwise.
+Our findings are easiest to reconcile in an interpretation that net saving adds nothing to capital growth.
+[^2] We argue, however, that productivity gain will not give equal capital growth unless gross saving is enough to offset depreciation.
+We interpret any shortfall in this coverage as a negative component in net saving, and expect capital growth to equal productivity gain less that deficit (see Section 3).
 Net saving, if so, raises the physical quantity of capital, but not the aggregate value, and so reduces the value per unit.
+This condition describes capital glut.
 The data are most easily explained by the present value principle, and by production efficiencies enabled through innovation.
@@
 [^3] Present value is time-discounted, and affords innovators only enough, at first, for first steps in implementation.
-As implementation proceeds, present value rises and affords more.
+As implementation proceeds, present value rises with time and affords more.
 Early investors borrow from the future, in effect, and pay the interest from increments in present value as positive cash flows approach with time.
 Meanwhile the innovator acquires materials and plant capacity and labor skills at market prices determined by their uses in current technology, but applies them more productively until competition catches up.
-It is that temporary market advantage to the innovator which explains capital growth without net saving in a practical and mechanical sense, while the present value principle gives the source of investment.
-This idea will be called “free growth theory” for easy reference.
+It is that temporary market advantage to the innovator, which we call the innovator’s reserve, that explains capital growth without net saving in a practical and mechanical sense, while the present value principle gives the source of investment.
+This idea will be called free growth theory for easy reference.
 It predicts only at the largest scales, and only for the private sector.
@@
 2 Net saving, capital growth and capital acceleration
-Thrift theory will mean the combined ideas that all net saving $S_{net}$ is realized in equal growth of capital $K$, if $S_{net}$ holds within the warranted rate, and that there is no other source of capital growth except market noise regressing to zero.
+Thrift theory will mean the combined ideas that net saving $S_{net}$ is realized in equal growth of capital $K$, if $S_{net}$ holds within the warranted rate, and that there is no other source of capital growth except market noise regressing to zero.
 "Thrift assumptions" will mean those two plus the assumptions that current $S_{net}$ holds within the warranted rate.
@@
 $\frac{\Delta s^{*}}{\Delta g(K)}=\frac{\Delta g(K)}{\Delta g(K)}=1\ ,\quad\text{uta}.$ (4)
-The argument has continued from growth in Eq.
-(2) to acceleration in Eqs.
-(3) and (4) on the reasoning that causal relationships among variables, in this case between thrift and capital growth rate, tend to be revealed most reliably in their concurrent changes.
+Algorithms for correlation and covariance and regression are structured on the reasoning that causal relationships among variables tend to be revealed in concurrent changes.
+While Eqs.
+(1) and (2) enable measurement of the coexistence of net saving and capital growth, Eqs.
+(3) and (4) allow a better measure of the degree to which that coexistence is explained by causality.
 Eq.
 (4) was derived from Eq.
-(3) so that success of predictions from growth assumptions can be measured against a standard of unity.
+(3) so that success of predictions from thrift assumptions can be measured against a standard of unity.
 For notational convenience, then, we define the "thrift index" $\theta$ by $\theta=\frac{\Delta s^{*}}{\Delta g(K)}$, and restate Eq.
@@
 3 Interpretation of test results
-The finding $\frac{s^{*}}{g(K)}=0.481$ reveals average observed proportions between $s^{*}$ and $g(K)$, and does not of itself reveal causal relatedness between those two variables.
+The finding $\frac{s^{*}}{g(K)}=0.481$ reveals average observed proportions between $s^{*}$ and $g(K)$, we repeat, and does not of itself show causal relatedness between those two variables.
 Causal relatedness is shown rather by the degree of constancy in proportions, and thus on regression of $s^{*}$ on $g(K)$ found at 0.0771.
@@
 Might these small but positive findings allow the possibility that some capital growth is explained by net saving?
-Free growth theory predicts that net saving contributes nothing to capital growth, given that gross investment offsets depreciation.
+Free growth theory predicts that the present value of new ideas can pay for equal capital growth, but only after gross saving has made up for all depreciation.
 In years of recession, households may defer part or all of depreciation saving to support consumption.
 In these years, growth in the present value of better ideas will not realize equal net growth of capital, as some of it must offset any shortfall in depreciation saving first.
-Allowing for these years, we can say more compactly that free growth theory predicts that a gain in the present value of new ideas accounts for all capital creation, and thus for the sum of capital growth and restoration of shortfalls in depreciation investment.
+Allowing for these years, we can say more compactly that free growth theory predicts that a gain in the present value of new ideas accounts for all capital creation, and thus for the sum of capital growth and restoration of shortfalls in depreciation saving.
 Thus free growth theory does not predict a zero value for $\theta$, but rather predicts $\theta$ to equal the share of capital creation applied to restoring those shortfalls.
-Insofar as 0.064 gives this share, free growth theory could fit the appearance in Fig.
+With this understanding, free growth theory could fit the appearance in Fig.
 2 that $\theta$ found higher values during the period of the Iraq-Kuwait, dot com and subprime crises.
-Pending clarification of that share, we will proceed on the simplifying assumption that free growth theory predicts correctly.
+We will suppose that 0.064 gives a plausible estimate of the shortfall in depreciation saving, until more is known, and will proceed on the simplifying assumption that free growth theory predicts correctly.
 4 Inferences for the creative destruction idea
@@
 Granting the obsolescence of old ways by new, it does not follow that innovation reduces values of assets already present.
-Data cited in this paper appear to show that capital growth is financed wholly by depreciation investment.
+Data cited in this paper appear to show that capital growth is maximized when investment is enough to offset depreciation.
 A natural source of that investment is depreciation of assets the new technology will replace, and a natural market for the new technology may be owners of the assets to be replaced.
@@
 This analysis suggests that market forces favor a kind through orderly obsolescence through replacement without premature displacement as the new technology awaits its turn.
-It also shows that the new technology as a separate entity grows by net investment financed from disinvestment from the technology replaced.
-It is only at the scale of the closed economy, meaning of all assets together, that net saving and investment appear to be unproductive.
 5 Solow’s puzzle
@@
 The present value argument just given offers an accounting explanation of embodied growth in terms of sources of funds.
-The step-by-step mechanical explanation, we suggest, in terms of things rather than dollars, is that embodied growth is disembodied growth on a finer scale.
+The step-by-step mechanical explanation, we suggest, in terms of things rather than funds, is that embodied growth is disembodied growth on a finer scale.
 At each step toward realization of the new plant and products, raw materials and products and labor skills and plant capacity currently available on the market are adapted to new uses.
@@
 Again, we go farther by questioning the assumption that net saving contributes even a part of capital growth.
+7 Tech Externalities
+The innovator’s reserve, quantified by the present value of the innovator’s idea, gives the amount of free growth to the innovator.
+Others may benefit indirectly.
+Old factories and warehouses and residences gain in utility as they house higher-tech equipment and worker skills.
+Old highways do the same as they carry higher-tech freight and drivers and passengers to higher-tech destinations in higher-tech vehicles.
+[^7] Rae (1834) makes essentially the same argument in terms of the technology of his time.
+Most externalities, [^8] Externalities are defined as effects of events on nonparticipants.
+positive and negative, are part of random market noise expected to offset to zero over time and scale.
+The examples above, which we call tech externalities, can impose a prevailingly positive vector resisting the downward pull of depreciation.
+Free growth, in the full sense, means effects of the innovator’s reserve on capital of the innovator, plus effects of tech externalities on capital in general.
 8 Optimum investment policy
@@
 Effects of removing the double tax, and removing the difference between tax rates on ordinary income and on capital gains, could be revenue-neutral and non-partisan if the corporate tax were raised to match, if the tax rates on ordinary income and on capital gains met somewhere between, and if thoughtful grandfathering eased the transition.
+9 Optimum government policy
+Households dissave in consumption, while firms dissave through dividends and other distributions.
+Public policy, in many nations, has inhibited dissaving in the belief that such policies favor capital growth.
+Rather they inhibit consumption, add nothing to capital growth, and thus inhibit net output as a function of both together (see Appendix B).
+The effect of repealing such policies would be to address capital glut not by investing less, but by disinvesting more.
+[^10] Higher dividends and distributions buy firms, that is, would enable higher consumption by households.
+This strategy spares government the burden of guiding investment, at least until the effects of the repeal are known.
 10 Inferences for national accounts and macroeconomics
@@
 13 Net output
-The concept of net output is found in our discussions, but not in our equations.
+In Keynesian tradition, net output equals consumption plus net investment.
+We have disputed this reasoning in our argument that net investment adds nothing to capital growth, inhibits consumption, and thus tends to reduce the sum of both.
+We also critique it on another ground.
+Assume, for the sake of argument, that we are wrong, and that net investment realizes equal capital growth.
 Net output means value added.
@@
 Both of these flows are unmeasurable by known means.
-Since we work here from measurable data for net saving and market value capital, we did our best to reason from those two quantities alone.
+Thus we have not sought to measure net output or its growth, and have treated capital growth alone.
 14 Data sources
-All our data are drawn from Distributional National Accounts (DINA) from the free online database World Inequality Database (WID).
+All our data are drawn from the free online website World Inequality Database (WID).
 This source collates data from national accounts and tax data of 105 countries in constant currency units, and adjusts them where needed to conform to current standards of the System of National Accounts (SNA) published by the United Nations.
@@
 Such age-based differences are expected to offset in the stationary state, where tradition and free growth theory agree that overall net saving is zero.
-We accept that individual net saving can realize equal individual capital growth, so long as collective saving and dissaving by all individuals offset to zero.
+We accept that individual net saving can realize equal individual capital growth, so long as collective saving does not exceed collective dissaving.
 18 Discussion and conclusions
@@
 It is loosely defined as oversupply of capital at the current state of technology.
-We will not attempt a more exact definition here.
-Findings shown in our displays, anyhow, suggest that net saving raises the physical quantity of capital, say in number of shops, manufacturing plants or finished goods of similar design, without raising aggregate value of capital, and so contributes to capital glut.
+We take it to mean or include a state where more investment does not realize more capital value.
+Findings shown in our displays suggest that net saving raises the physical quantity of capital, say in number of shops, manufacturing plants or finished goods of similar design, without raising aggregate value of capital, and so contributes to capital glut by that understanding.
 These findings challenge the teachings that capital growth is effected by net saving enabled by consumption restraint, and that producer cost, including imputed interest as the opportunity cost of capital, converges to market realization.
@@
 Appendix A.  Equivalence of saving and investment
-Any usage which treats saving and investment as equal must deal with the fact that saving held in cash is not investment in the usual sense of spending on new productive capital, and contributes nothing to output.
-Keynes (1936) defined that unspent saving as intended saving, and actual saving as the part so spent.
-We suggest that investors seek to maximize return, within risk tolerance, and will sometimes hold saving in cash or equivalents at zero return in recessions and depressions when positive returns cannot be found, and so when investing in the usual sense would tend to reduce capital and output rather than increase them.
-Thus investment could have the usual meaning in usual times, when prospective returns bring animal spirits, and could include saving in cash when not.
-It is in this sense, differing from Keynes by equating intended (cash) investment to actual investment when firms and households see nothing better, that we equate saving to investment.
+Keynes (1936) argued correctly that saving, if not invested in currently produced producer goods and services, adds nothing to capital, and thus nothing to output.
+He inferred that uninvested saving does not achieve actual saving, as it accomplishes no growth, and described it as "intended saving." True saving, by that reasoning, equalled investment.
+We suggest that investors seek to maximize return, within risk tolerance, and will sometimes hold saving in cash or equivalents at zero return in recessions and depressions when positive returns cannot be found, and so when investing would tend to reduce capital and output rather than increase them.
+Thus we do not share the Keynesian view that saving need be invested in order to maximize its value.
+Saving will tend to equal investment in normal times, when prospective returns bring animal spirits, and could include saving in cash when not.
+It is only for these profitable times that we accept the equality of saving and investment, and only as a simplifying assumption.
+[^12] The weakness in the Keynes argument, if we are right, was his inference that saving does not accomplish its purpose unless invested to realize capital growth.
+Its purpose, we argue, is to maximize capital growth according to investment opportunities, and to maximize it by withholding investment when all investments promise loss.
 Appendix B. Net output with human capital

### v13

@@
 Abstract
-Data from national accounts support an argument that capital growth and acceleration arrive without help from net saving.
+Data from national accounts suggest that net saving, at the collective scale, contributes less than 8% to growth in capital value.
 We explore ways in which this is possible, and discuss implications for economic teaching and public policy.
@@
 Economists since the early nineteenth century have added the proviso that net saving cannot safely outpace innovation; more capital must mean capital redesigned for greater productivity if economies are to escape risk of capital glut and diminishing returns (West (1815), Ricardo (1815), Malthus (1815)).
-Roy Harrod (1939)) described that limit for safe net saving, meaning the rate of imagining and developing new ideas for more productive forms of capital, as the “warranted rate”.
+Roy Harrod (1939) described that limit for safe net saving, meaning the rate of imagining and developing new ideas for more productive forms of capital, as the “warranted rate”.
 Harrod, and many other economists of his time and since, have focused on growth of output rather than of capital, but have modeled growth of output by first assuming the equivalence of net saving and capital growth, within the warranted rate, and then looking for effects of that capital growth on later output growth.
@@
 All data are drawn from national accounts of those countries as collated on the free website World Inequality Database.
-Our findings are easiest to reconcile in an interpretation that net saving adds nothing to capital growth.
-[^2] We argue, however, that productivity gain will not give equal capital growth unless gross saving is enough to offset depreciation.
-We interpret any shortfall in this coverage as a negative component in net saving, and expect capital growth to equal productivity gain less that deficit (see Section 3).
-Net saving, if so, raises the physical quantity of capital, but not the aggregate value, and so reduces the value per unit.
+Our findings imply that the efficiency of net saving in realizing capital growth has been less than 8%.
+We will argue that it may be substantially less, and possibly zero.
+[^2] We will find three measures, of which the highest is 7.71%, of the causal relatedness between net saving and capital growth.
+These measures do not give the direction of causality.
+We will argue that exogenous disruptions such as wars or epidemics or supply or demand shocks can bring market value and income down as a direct effect, and that households, in response, may be motivated to trim saving in order to maintain consumption.
+In these cases, the drop in capital growth $g(K)$ would precede and explain the drop in net saving $S_{net}$.
+The same direction of causality would appear in recoveries from these setbacks.
+As capital values rise again, that is, income would rise with them, and households would be motivated to restore old saving rates as insurance against the next downturn.
+(See Appendix E).
+Net saving, allowing for that margin of possible exception, has raised the physical quantity (volume) of capital, but not the aggregate value, and so has reduced the value per unit.
 This condition describes capital glut.
-The data are most easily explained by the present value principle, and by production efficiencies enabled through innovation.
-Value is created in the mind of the market at the moment when prospective cash flows are discounted.
-It is created only if the market sees a path, step by step, from the start, to practical realization of those prospective cash flows.
-Then capital growth arrives when the market first evaluates prospective cash flows, and is realized eventually in physical outcomes insofar as the market has predicted correctly.
-That is to say that capital growth, in the mind of the market, precedes and pays for any needed investment in new plant and equipment.
-[^3] Present value is time-discounted, and affords innovators only enough, at first, for first steps in implementation.
-As implementation proceeds, present value rises with time and affords more.
-Early investors borrow from the future, in effect, and pay the interest from increments in present value as positive cash flows approach with time.
-Meanwhile the innovator acquires materials and plant capacity and labor skills at market prices determined by their uses in current technology, but applies them more productively until competition catches up.
-It is that temporary market advantage to the innovator, which we call the innovator’s reserve, that explains capital growth without net saving in a practical and mechanical sense, while the present value principle gives the source of investment.
-This idea will be called free growth theory for easy reference.
-It predicts only at the largest scales, and only for the private sector.
-Individuals and groups and even small economies can grow in capital through saving if others dissave equally.
-That possibility is foreclosed only at the scale of all capital and all economies together.
+The explanation we suggest for capital growth without net saving is in production efficiencies of innovation.
+The innovator acquires materials and plant capacity and labor skills at market prices determined by their uses in current technology, but applies them more productively until competition catches up.
+That temporary advantage to the innovator in market value of output per unit cost of input, which we call the innovator’s reserve, creates new income from which to pay that cost, and leaves original income free to pay for consumption.
+This new income expands the economy, meanwhile, and justifies an equal expansion of the money supply to maintain price stability.
+Thus growth by a gain in productivity of existing resources costs no sacrifice in consumption, and no diversion of funds from other uses.
+An innovator in need of outside funding may approach investors.
+Investors calculate present value, just as with projects relying on established rather than innovative technology, and invest up to that amount.
+[^3] Present value is the sum of time-discounted cash flows expected.
+These flows tend to be negative over any period of investment, and then positive as products are finished and sold.
+An innovative project may need more investment than others, and a longer period of investment, and will not reward investment unless positive cash flows at the end, discounted to present value, are large enough to offset., [^4] An innovator not in need of outside funding would still require equal present value as a minimum condition for investment., [^5] Competition among investors should tend to keep investment close to present value.
+Innovative projects may need more investment over a longer period than others, but will still justify funding if expected return at the end is seen as enough to bring present value equal at least to investment.
+Any excess of present value over investment quantifies the investor’s reserve.
+[^6] Appendix D will show that an innovative investment will produce cash flows forward at the rate of return $r_{I}$ realized by the innovation, while those cash flows will be discounted backward to present value at the market time-discount rate $r_{m}$ (see Appendix D).
+We further suggest, without evidence, that innovation should tend to bring positive externalities to capital of third parties not participating in the innovation.
+Old structures and infrastructure, for example, should tend to gain value as they house or support higher-tech workers and equipment.
+We call this effect "tech externalities." "Free growth" will mean the combined growth in capital of innovators through the innovator’s reserve, and in capital generally through tech externalities.
+Free growth theory will mean the prediction that the rate of productivity gain, which Harrod called the warranted rate, tends to bring an equal rate of capital growth through the innovator’s reserve and tech externalities, without help from net saving, and so brings most or all of capital growth not explained by market noise.
+Thus the theory allows for some contribution by net saving, but expects any such contribution to be small.
+We will use the term "thrift residue" to mean any part of capital growth that might be explained by net saving, less than 8% by our findings, with the understanding that the part actually explained by net saving may be zero.
+Although net saving is expected to be largely or wholly ineffective, the theory does not question the value of gross saving to make up for losses to depreciation.
+If these losses are not made up first, value created by productivity gain will go first to make up this shortfall, and only the remainder will apply to capital growth.
+The theory predicts only at the largest scales, where the data are clear, and only for the private sector.
+We accept that households commonly grow or decline in net worth through net saving or dissaving (see Section 16), and that the same is possible for larger groups and even small nations.
+Free growth theory predicts only at the scale of all capital and all economies together.
 The public sector, meanwhile, responds to political rather than market choices, and grows or shrinks accordingly.
@@
 One of the central doctrines of the marginalist revolution has held that market realization converges to producer cost, when that cost includes imputed interest on assets owned.
-Net saving gives producer cost, and falls short of market realization in the presence of technological growth from new ideas.
-Meanwhile the doctrine that net income equals consumption plus net saving is put into question by evidence offered here suggesting that net saving increases the physical quantity of capital, but not the aggregate value.
+A gain in productive efficiency can bring market value above producer cost, and by most or all of capital growth if free growth theory holds.
+Meanwhile the doctrine that net income equals consumption plus net saving is put into question by evidence offered here suggesting that net saving increases the physical quantity of capital, but adds less than 8% to aggregate value.
 2 Net saving, capital growth and capital acceleration
 Thrift theory will mean the combined ideas that net saving $S_{net}$ is realized in equal growth of capital $K$, if $S_{net}$ holds within the warranted rate, and that there is no other source of capital growth except market noise regressing to zero.
-"Thrift assumptions" will mean those two plus the assumptions that current $S_{net}$ holds within the warranted rate.
+"Thrift assumptions" will mean those two plus the assumption that current $S_{net}$ holds within the warranted rate.
 The acronym uta will mean "under thrift assumptions." Then
@@
 $g(K)=s^{*}\ ,\quad\text{uta,}\quad\text{where }s^{*}=\frac{S_{net}}{K}.\quad\text{ Then}$ (2)
-$\Delta g(K)=\Delta s^{*}\ ,\quad\text{uta.}$ (3)
+[…]
 $\Delta g(K)$ and $\Delta s^{*}$ respectively will be called "capital acceleration" and "thrift change," either of which may be positive or negative or zero.
@@
 $\theta=1\ ,\quad\text{uta,}\quad\text{where}\ \theta=\frac{\Delta s^{*}}{\Delta g(K)}\ .$ (5)
-Test results for the predictions $\frac{s^{*}}{g(K)}=\theta=1$, under thrift assumptions, appear in Fig.
+Test results for the predictions $\frac{s^{*}}{g(K)}=1\ \text{and}\ \theta=1$, under thrift assumptions, appear in Fig.
 1 and in Tables 1, 2 and 3.
@@
 Table 1 shows regression of $s^{*}$ on $g(K)$ as 0.0771, and regression of $\Delta s^{*}$ on $\Delta g(K)$ as 0.0559.
-Table 1: Regression of $s^{*}$ on $g(K)$ and $\Delta s^{*}$ on $\Delta g(K)$, GDP-weighted. Screen = 0.01. $H_{0}$ per thrift theory: $\frac{s^{*}}{g(K)}=\frac{\Delta s^{*}}{\Delta g(K)}=1$.
+3 Interpretation of test results
+The finding $\frac{s^{*}}{g(K)}=0.481$ reveals average observed proportions between $s^{*}$ and $g(K)$, we repeat, and does not of itself show causal relatedness between those two variables.
+Causal relatedness is shown rather by the degree of constancy in proportions, and thus on regression of $s^{*}$ on $g(K)$ found at 0.0771.
+$\theta$ or $\frac{\Delta s^{*}}{\Delta g(K)}$, measured at $0.064$, gives a separate measure of causal relatedness between $s^{*}$ and $g(K)$ as explained.
+Regression of $\Delta s^{*}$ on $\Delta g(K)$, found at 0.0559, gives a measure of causal relatedness between $\Delta s^{*}$ and $\Delta g(K)$.
+Might these small but positive findings allow the possibility that some capital growth is explained by net saving?
+Appendix E will develop our argument in Section 1 (footnote 2) that the causal relatedness revealed in $\theta$ and its associated regressions might also be explained as joint effects of outside events affecting $g(K)$ directly, and $s^{*}$ indirectly in adjustment to those effects on $g(K)$.
+We also note that investment in government-owned property such as infrastructure is excluded from predictions of free growth theory, but included in net investment reported in the national accounts from which we take our data.
+Much government-owned property is unsold, and impractical to appraise, and accordingly is valued in national accounts at depreciated cost.
+Thus net investment in such government assets shows in national accounts as equal to capital growth, and contributes to $\theta$.
+Free growth theory takes no position on the actual role of net saving in growth of government-owned property, and consequently neglects this contribution to $\theta$ as outside the scope of its prediction.
+4 Inferences for the creative destruction idea
+Creative destruction (Schumpeter (1912) [^8] Published by Schumpeter in 1912, and translated by Opie in 1934.) is obsolescence of existing capital through innovation of new and preferred technology, as with horse-drawn carriages and automobiles.
+Granting the obsolescence of old ways by new, it does not follow that innovation reduces values of assets already present.
+We have argued that capital growth is maximized when investment is enough to offset depreciation first.
+A natural source of that investment is depreciation of assets the new technology will replace, and a natural market for the new technology may be owners of the assets to be replaced.
+Carriage owners were better able to afford cars when the carriages in which they had invested before lived out their economic lives as originally expected, so that the depreciation component in the cash flows originally expected can pay for cars in replacement just as it might have paid for new carriages.
+This analysis suggests that market forces favor a kind of orderly obsolescence through replacement, without premature displacement, as the new technology awaits its turn.
+Table 1: Regression of $s^{*}$ on $g(K)$ and $\Delta s^{*}$ on $\Delta g(K)$, GDP-weighted. Screen = 0.01. $H_{0}$ per thrift theory: $\frac{s^{*}}{g(K)}\cong\frac{\Delta s^{*}}{\Delta g(K)}\cong 1$.
 Regression of $s^{*}$ on $g(K)$ 0.0771∗∗∗
@@
 Fig. 2: Average $\theta_{s}^{*}$ with LOESS smoothing over all countries, GDP-weighted, 1980-2022
-Table 2: Average $\frac{s^{*}}{g(K)}$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()
+Table 2: Average $\frac{s^{*}}{g(K)}$ in 92 countries over periods shown. Screen = 0.01. Number of years clearing screen shown in ()
 Armenia 1996 - 2020 (23) -0.63 Aruba 1996 - 2001 (6) 1.25 Australia 1980 - 2018 (33) 0.24 Austria 1996 - 2021 (23) 0.83 Azerbaijan 1996 - 2020 (23) 1.57 Bahrain 2009 - 2013 (5) -2.43 Belgium 1996 - 2021 (21) 0.61 Bolivia 1998 - 2015 (17) 0.40 Botswana 1996 - 1999 (4) 3.54 Brazil 1996 - 2019 (21) 0.25 British Virgin Islands 1996 - 1999 (4) 2.45 Bulgaria 1996 - 2016 (16) 0.14 Burkina Faso 2000 - 2019 (20) 0.23 Cameroon 1996 - 2019 (24) 1.10 Canada 1980 - 2022 (38) 0.37 Cape Verde 2008 - 2017 (9) 0.53 Chile 1997 - 2021 (19) 0.84 China 1992 - 2020 (29) 0.81 Colombia 1996 - 2022 (27) 0.60 Costa Rica 2013 - 2020 (6) 0.21 Croatia 1996 - 2021 (24) 0.43 Curaçao 2001 - 2016 (14) 1.26 Cyprus 1996 - 2021 (25) 0.07 Czechia 1994 - 2020 (20) 0.17 Côte d’Ivoire 1996 - 2000 (5) 0.29 Denmark 1996 - 2022 (26) 0.22 Dominican Republic 1996 - 2016 (10) 1.52 Ecuador 2008 - 2020 (13) 0.53 Egypt 1997 - 2015 (19) 0.92 El Salvador 2015 - 2019 (5) -0.62 Estonia 1996 - 2021 (24) 0.45 Finland 1996 - 2021 (23) 0.17 France 1980 - 2021 (35) 0.26 Germany 1980 - 2022 (38) 0.70 Greece 1996 - 2021 (26) 0.17 Guatemala 2002 - 2021 (20) -0.72 Guinea 2004 - 2010 (5) 0.72 Honduras 2001 - 2015 (14) -0.57 Hong Kong 1997 - 2011 (15) 0.90 Hungary 1996 - 2021 (21) 0.14 Iceland 2001 - 2014 (14) 0.22 India 1999 - 2019 (21) 0.57 Indonesia 2017 - 2019 (3) 0.94 Iran 1997 - 2017 (20) 4.96 Ireland 1996 - 2021 (22) 0.26 Israel 1996 - 2019 (24) 0.46 Italy 1980 - 2022 (36) 0.13 Japan 1980 - 2021 (29) 0.19 Kazakhstan 1997 - 2022 (25) 0.60 Kuwait 2003 - 2017 (14) 1.11 Kyrgyzstan 1996 - 2021 (24) 0.46 Latvia 1996 - 2021 (24) -0.23 Lithuania 1996 - 2021 (24) 0.13 Luxembourg 1996 - 2021 (22) 0.03 Malaysia 2007 - 2015 (9) 2.52 Malta 1996 - 2021 (23) 0.16 Mauritius 2014 - 2018 (5) 0.04 Mexico 1996 - 2021 (22) 0.15 Moldova 1996 - 2019 (22) -0.91 Mongolia 2006 - 2020 (15) 0.10 Morocco 1999 - 2021 (23) 1.13 Netherlands 1996 - 2021 (23) 0.51 New Zealand 1996 - 2019 (21) 0.78 Nicaragua 2007 - 2018 (12) 0.08 Niger 1996 - 2019 (24) 0.90 Norway 1981 - 2021 (38) 0.95 Peru 2008 - 2021 (14) 0.76 Philippines 1996 - 2022 (27) 1.76 Poland 1996 - 2021 (25) 0.71 Portugal 1996 - 2022 (24) 0.00 Qatar 2002 - 2017 (14) 1.89 Romania 1996 - 2020 (22) 0.03 Russia 1996 - 2017 (13) -0.19 Saudi Arabia 2003 - 2009 (7) 2.65 Senegal 2015 - 2021 (7) 0.24 Serbia 1998 - 2021 (23) 0.61 Slovakia 1996 - 2022 (24) 0.15 Slovenia 1996 - 2021 (20) 0.28 South Africa 1996 - 2022 (24) 0.28 South Korea 1996 - 2020 (24) 0.63 Spain 1995 - 2021 (25) 0.32 Sweden 1980 - 2021 (37) 0.41 Switzerland 1993 - 2021 (25) 0.69 Tunisia 1996 - 2011 (16) 0.29 Turkey 2010 - 2017 (8) 1.32 USA 1980 - 2021 (38) 0.31 Ukraine 1996 - 2019 (21) 0.17 United Kingdom 1981 - 2021 (36) 0.16 Uruguay 2016 - 2017 (2) 1.16 Uzbekistan 2011 - 2021 (10) 0.21 Vanuatu 2002 - 2007 (6) 0.29 Venezuela 1998 - 2019 (20) 0.55
-Note: $s^{*}$ is defined as $\frac{S_{net}}{K}$. Thrift theory predicts $\frac{s^{*}}{g(K)}=1$. Free growth theory makes no prediction for these data.
-Table 3: Average $\theta$ in 92 countries (screen = 0.01). Number of years clearing screen shown in ()
+Note: $s^{*}$ is defined as $\frac{S_{net}}{K}$. Thrift theory predicts $\frac{s^{*}}{g(K)}\cong 1$. Free growth theory makes no prediction for these data.
+Table 3: Average $\theta$ in 92 countries over periods shown. Screen = 0.01. Number of years clearing screen shown in ()
 Armenia 1997 - 2020 (20) 0.12 Aruba 1997 - 2001 (5) 1.22 Australia 1974 - 2019 (35) 0.03 Austria 1997 - 2021 (16) 0.10 Azerbaijan 1997 - 2020 (23) 0.75 Bahrain 2010 - 2013 (4) 0.41 Belgium 1997 - 2021 (19) 0.05 Bolivia 1999 - 2014 (12) 0.04 Botswana 1997 - 2000 (4) 0.24 Brazil 1998 - 2020 (22) 0.10 British Virgin Islands 1997 - 1998 (2) 2.24 Bulgaria 1997 - 2017 (15) -0.03 Burkina Faso 2001 - 2019 (16) 0.04 Cameroon 1997 - 2019 (19) 0.18 Canada 1974 - 2022 (41) 0.09 Cape Verde 2009 - 2016 (8) 0.27 Chile 1998 - 2021 (22) 0.06 China 1981 - 2020 (33) 0.07 Colombia 1997 - 2022 (18) -0.03 Costa Rica 2014 - 2020 (6) 0.04 Croatia 1997 - 2021 (22) 0.19 Curaçao 2002 - 2016 (14) 0.41 Cyprus 1997 - 2021 (23) -0.06 Czechia 1995 - 2021 (19) -0.03 Côte d’Ivoire 1997 - 2000 (4) -0.03 Denmark 1997 - 2022 (22) 0.05 Dominican Republic 2007 - 2015 (9) 0.15 Ecuador 2009 - 2020 (12) 0.17 Egypt 1998 - 2015 (16) 0.13 El Salvador 2016 - 2019 (4) 0.98 Estonia 1997 - 2021 (19) 0.06 Finland 1997 - 2022 (21) -0.02 France 1972 - 2021 (33) 0.09 Germany 1972 - 2022 (31) 0.05 Greece 1997 - 2021 (24) 0.24 Guatemala 2003 - 2021 (11) 0.15 Guinea 2005 - 2010 (6) 0.08 Honduras 2003 - 2015 (10) 0.40 Hong Kong 1997 - 2011 (13) 0.04 Hungary 1997 - 2021 (21) 0.07 Iceland 2002 - 2014 (12) 0.10 India 1997 - 2019 (16) 0.14 Indonesia 2019 - 2019 (1) -0.41 Iran 1998 - 2018 (19) 1.09 Ireland 1997 - 2021 (23) 0.02 Israel 1997 - 2017 (18) 0.07 Italy 1972 - 2021 (29) 0.05 Japan 1972 - 2021 (38) 0.05 Kazakhstan 1997 - 2022 (23) 0.08 Kuwait 2005 - 2017 (11) 0.04 Kyrgyzstan 1998 - 2021 (22) 0.23 Latvia 1998 - 2021 (20) 0.01 Lithuania 1997 - 2021 (18) 0.10 Luxembourg 1997 - 2021 (25) 0.00 Malaysia 2008 - 2015 (7) 0.01 Malta 1997 - 2021 (23) 0.08 Mauritius 2015 - 2018 (4) 0.02 Mexico 1997 - 2021 (22) 0.06 Moldova 1997 - 2019 (21) 0.11 Mongolia 2008 - 2021 (11) 0.10 Morocco 2000 - 2021 (17) 0.19 Netherlands 1997 - 2021 (20) 0.03 New Zealand 1997 - 2019 (18) 0.14 Nicaragua 2008 - 2019 (12) 0.06 Niger 1997 - 2019 (19) 0.22 Norway 1982 - 2021 (34) 0.12 Peru 2009 - 2021 (11) 0.04 Philippines 1997 - 2022 (20) 0.37 Poland 1997 - 2021 (15) 0.20 Portugal 1997 - 2022 (17) 0.02 Qatar 2004 - 2018 (12) 0.29 Romania 1997 - 2020 (14) 0.02 Russia 1997 - 2018 (14) 0.09 Saudi Arabia 2005 - 2009 (3) 2.23 Senegal 2018 - 2021 (3) 0.19 Serbia 1999 - 2021 (22) 0.22 Slovakia 1997 - 2021 (18) 0.09 Slovenia 1997 - 2021 (20) 0.10 South Africa 1997 - 2022 (21) 0.10 South Korea 1997 - 2020 (15) 0.04 Spain 1972 - 2021 (42) 0.06 Sweden 1974 - 2022 (41) 0.02 Switzerland 1993 - 2020 (24) 0.08 Tunisia 1997 - 2011 (12) 0.32 Turkey 2011 - 2015 (5) -0.16 USA 1972 - 2021 (44) 0.02 Ukraine 1997 - 2020 (23) 0.05 United Kingdom 1972 - 2021 (42) 0.02 Uruguay 2017 - 2017 (1) 0.56 Uzbekistan 2012 - 2021 (9) 0.05 Vanuatu 2003 - 2007 (3) 0.15 Venezuela 1999 - 2019 (20) 0.31
-Note: $\theta$ is defined as $\frac{\Delta s^{*}}{\Delta g(K)}$. Thrift theory predicts $\theta=1$. See Section 3 for predictions of free growth theory.
-3 Interpretation of test results
-The finding $\frac{s^{*}}{g(K)}=0.481$ reveals average observed proportions between $s^{*}$ and $g(K)$, we repeat, and does not of itself show causal relatedness between those two variables.
-Causal relatedness is shown rather by the degree of constancy in proportions, and thus on regression of $s^{*}$ on $g(K)$ found at 0.0771.
-$\theta$ or $\frac{\Delta s^{*}}{\Delta g(K)}$, measured at $0.064$, gives a separate measure of causal relatedness between $s^{*}$ and $g(K)$ as explained.
-Regression of $\Delta s^{*}$ on $\Delta g(K)$, found at 0.0559, gives a measure of causal relatedness between $\Delta s^{*}$ and $\Delta g(K)$.
-Might these small but positive findings allow the possibility that some capital growth is explained by net saving?
-Free growth theory predicts that the present value of new ideas can pay for equal capital growth, but only after gross saving has made up for all depreciation.
-In years of recession, households may defer part or all of depreciation saving to support consumption.
-In these years, growth in the present value of better ideas will not realize equal net growth of capital, as some of it must offset any shortfall in depreciation saving first.
-Allowing for these years, we can say more compactly that free growth theory predicts that a gain in the present value of new ideas accounts for all capital creation, and thus for the sum of capital growth and restoration of shortfalls in depreciation saving.
-Thus free growth theory does not predict a zero value for $\theta$, but rather predicts $\theta$ to equal the share of capital creation applied to restoring those shortfalls.
-With this understanding, free growth theory could fit the appearance in Fig.
-2 that $\theta$ found higher values during the period of the Iraq-Kuwait, dot com and subprime crises.
-We will suppose that 0.064 gives a plausible estimate of the shortfall in depreciation saving, until more is known, and will proceed on the simplifying assumption that free growth theory predicts correctly.
-4 Inferences for the creative destruction idea
-Creative destruction (Schumpeter (1912) [^4] Published by Schumpeter in 1912, and Translated by Opie in 1934.) is obsolescence of existing capital through innovation of new and preferred technology, as with horse-drawn carriages and automobiles.
-Granting the obsolescence of old ways by new, it does not follow that innovation reduces values of assets already present.
-Data cited in this paper appear to show that capital growth is maximized when investment is enough to offset depreciation.
-A natural source of that investment is depreciation of assets the new technology will replace, and a natural market for the new technology may be owners of the assets to be replaced.
-Carriage owners were better able to afford cars when the carriages in which they had invested before lived out their economic lives as originally expected, so that the depreciation component in the cash flows originally expected can pay for cars in replacement just as it might have paid for new carriages.
-This analysis suggests that market forces favor a kind through orderly obsolescence through replacement without premature displacement as the new technology awaits its turn.
+Note: $\theta$ is defined as $\frac{\Delta s^{*}}{\Delta g(K)}$. Thrift theory predicts $\theta\cong 1$. See Section 3 for predictions of free growth theory.
 5 Solow’s puzzle
@@
 But it did not… I do not know if that finding should be described as a paradox, but it was at least a puzzle."
-Solow was right in that much innovation requires months or years of gross investment to pay for negative cash flows in building new plant before the new products roll out from the factory doors.
-What he left unsaid is that this investment would not have been forthcoming unless positive cash flows at the end of the tunnel were predicted to be sufficient to create positive present value from the start.
-Data in our displays suggest that this initial creation of value, plus depreciation investment, are enough.
+Solow was right in that much innovation requires months or years of negative cash flow, in the form of gross investment, in building new plant before the new products roll out from the factory doors.
+From arguments in Section 1, however, this investment would not have been forthcoming unless positive cash flows at the end of the tunnel were predicted to be sufficient to create equal or greater present value from the start.(See Appendix D.) That initial creation of present value, plus replacement investment, [^10] Also called depreciation investment.
+are enough.
 6 Mechanics of free growth
-The present value argument just given offers an accounting explanation of embodied growth in terms of sources of funds.
-The step-by-step mechanical explanation, we suggest, in terms of things rather than funds, is that embodied growth is disembodied growth on a finer scale.
+In terms of Solow’s puzzle, we suggest that embodied growth is disembodied growth on a finer scale.
 At each step toward realization of the new plant and products, raw materials and products and labor skills and plant capacity currently available on the market are adapted to new uses.
 The innovator pays for these inputs at a market price determined by their value in established productive uses, but applies them innovatively to realize higher prospective cash flows, and hence higher present values, to the innovator (Marshall (1890), Schumpeter and Opie (1934)).
-This difference in present value realized less price paid will here be called the “innovator’s reserve”, meaning reserve price for inputs of capital and labor.666i.e., capital and labor inputs are worth more to the innovator in that the innovator applies them in ways to realize greater returns.
-The present value of additional cash flow enabled by this advantage in return quantifies the innovator’s reserve, and equivalently the non-random component of free growth.
-The innovator’s reserve quantifies the part of free growth explained by productivity gain as distinct from random market noise.
-As such, it is the quantity added to depreciation saving to enable embodied growth, so that net saving is never needed.
+That new use of materials and skills applied could give the solution to Solow’s puzzle with respect to practical feasibility, as argued in Section 1, while the economic advantage to the investor, here called innovator’s reserve,111111i.e., reserve price which the innovator, who realizes more value from capital and labor inputs, might have been willing to pay for those resources.
+could give the solution with respect to source of payment.
+Appendix D will clarify.
 Our findings support those of Piketty and Zucman (2014) and Kurz (2023) as to the market power of innovators to explain capital growth beyond net saving.
-Those studies assumed that net saving is recovered in capital growth, so that this market power, which we call the innovator’s reserve, need explain only the part of capital growth remaining.
+Those studies assumed that net saving is recovered in capital growth, so that this market power, meaning the innovator’s reserve, would explain only the part of capital growth remaining.
 Again, we go farther by questioning the assumption that net saving contributes even a part of capital growth.
 7 Tech Externalities
-The innovator’s reserve, quantified by the present value of the innovator’s idea, gives the amount of free growth to the innovator.
-Others may benefit indirectly.
+We argued in Section 1 that innovation can add value eslewhere.
 Old factories and warehouses and residences gain in utility as they house higher-tech equipment and worker skills.
@@
 positive and negative, are part of random market noise expected to offset to zero over time and scale.
-The examples above, which we call tech externalities, can impose a prevailingly positive vector resisting the downward pull of depreciation.
+The examples above, which we call tech externalities, are exceptions in that they can impose a prevailingly positive vector on capital value resisting the downward pull of depreciation.
 Free growth, in the full sense, means effects of the innovator’s reserve on capital of the innovator, plus effects of tech externalities on capital in general.
+Note that Solow’s capital deepening, or growth through more efficient use of resources, fits the definition of productivity gain affording the innovator’s reserve.
+The distinction between capital deepening and widening is otherwise complex.
+Early investors in technological gain, say venture capitalists, may be rewarded with returns above the current norm.
+Such early backers form a kind of second-order innovators, and that advantage in return contributes to the innovator’s reserve.
+Meanwhile other producers may learn and adopt the same technology in a form of externality called spillover, and thus earn above-average returns before the market catches up.
+Products made by those producers may tend to duplicate those made by the original innovator, but are innovative by comparison to products not yet adopting the innovation.
+Thus we do not seek to draw a clear line between capital deepening and capital widening, or between tech externalities and the innovator’s reserve.
+8 Market value and book value
+It is natural to reason that depreciation gives the only change in capital value, once investment has ended, apart from market noise converging to zero.
+Insofar as depreciation can be predicted, then, depreciated cost would show the norm (expectation) toward which current market value converges.
+It would follow that investors might anticipate the market by selling when prices are above depreciated cost, and buying when below.
+This reasoning would justify the reliance of modern national accounts, and the equations of macroeconomics (macro) which they were designed to inform, since the inception of both in Keynes’s time, on depreciated cost as the preferred measure of capital value.
+[^14] As argued by James Tobin in 1968 and 1969.
+”Tobin’s qq” is the ratio of market to replacement value (RV).
+RV means depreciated cost under replacement cost accounting, where past investments and subsequent depreciation are re-expressed in the cheaper dollars of today.
+Investors were advised to invest when Tobin’s $q$ is less than one, and divest or withhold investment conversely.
+We have argued that capital grows and declines with effects of the innovator’s reserve and tech externalities, as well as investment and depreciation and market noise.
+Consider, however, a period when innovation is slight, so that depreciation and market noise give substantially the only changes in value of an investment once booked.
+It would follow that depreciated cost (of investment) could find the norm described, even if free growth were absent and depreciation were predicted perfectly, only insofar as did original cost.
+An argument that optimal accounting can give a sounder measure of worth than market value would seem to need a supporting argument that price and worth were closer together at the time assets were acquired or created, and depreciation bases set at their cost, than over the depreciation period following.
+Garbage in, garbage out.
+We do not question the value of depreciation accounting, or the importance of the depreciation concept in economic understanding.
+But we suggest that current market value gives a more reliable measure of capital, when both are known, than an assumption-dependent projection from past market value subject to the same criticisms as current market value.
+These reflections help explain our own reliance on market-value capital.
+They might also suggest rethinking of the logic by which macro and national accounts have preferred to reason from replacement cost when data for market-value capital are available.
 9 Optimum investment policy
 Data and arguments suggest that the optimum amount of saving, at the global scale, is replacement saving to offset depreciation, and nothing more.
-That would not mean book depreciation, as this study has stressed differences between book and market values.
-Up to a point, it should be possible to analyze the composition of market capital, and to model depreciation of the whole.
-A better plan, as Solow (1956) wrote in response to Harrod’s knife edge argument (Harrod, 1939), is to trust the market to maximize rate of return, and to sense the point where glut begins and returns fall.
+That would not mean depreciation found by the depreciated cost method, which equates net investment to capital growth at all scales, and which cannot measure the contributions of the innovator’s reserve or tech externalities.
+Up to a point, it should be possible to analyze the composition of market-value capital, and to model depreciation of the whole, just as national balance sheet accounting does with replacement value capital today.
+A better plan, as Solow (1956) wrote in response to Harrod’s knife edge argument (Harrod, 1939), is to expect the market to maximize rate of return, and to sense the point where glut begins and returns fall.
 [^15] Harrod had argued that saving must hit the warranted rate exactly or risk positive feedback through the operation of the output/capital ratio (accelerator).
 Markets do so imperfectly when tax and other public policy reward saving over distributions and consumption.
-Findings in this paper suggest review of such policies.
+Findings in this paper suggest reconsideration of such policies.
 These include the double tax on dividends, and the greater tax rate on ordinary income than on capital gains.
@@
 Households dissave in consumption, while firms dissave through dividends and other distributions.
-Public policy, in many nations, has inhibited dissaving in the belief that such policies favor capital growth.
-Rather they inhibit consumption, add nothing to capital growth, and thus inhibit net output as a function of both together (see Appendix B).
+Public policy, in many nations, has inhibited dissaving in the belief that policies to promote net saving favor capital growth.
+Rather they inhibit consumption, add less than 8% to capital growth, and thus inhibit net output as a function of both together (see Appendix B).
 The effect of repealing such policies would be to address capital glut not by investing less, but by disinvesting more.
-[^10] Higher dividends and distributions buy firms, that is, would enable higher consumption by households.
-This strategy spares government the burden of guiding investment, at least until the effects of the repeal are known.
-10 Inferences for national accounts and macroeconomics
-We have attempted to show that net saving does not tend to equate to capital growth, and indeed contributes no part of capital growth.
-If our argument proves correct, then national accounts, and the teachings of macroeconomics (macro) which they express, are constructed on a fundamental misperception.
-We would therefore welcome any critiques of our findings as a matter of academic and public interest.
+[^16] Higher dividends and distributions by firms, that is, would enable higher consumption by households.
+Demand for dividends should rise in response as shareholders are freed from the burden of the double tax, and firms would be incentivized to compete to meet that demand.
+As firms did so, and raised dividends, firms would find a lower share of income left to invest, and would tend to invest more selectively.
+Lower-return investment opportunities would tend to find fewer bidders, and average returns, in consequence, should tend to rise.
+Government might watch and wait, in such a scenario, and weigh other policy as outcomes appear.
 11 Inferences for microeconomics
@@
 It teaches that producer cost, including interest as the cost of time, tends to equilibrate to market value realized.
-We have proposed an exception in the innovator’s reserve; the innovator pays the going rates for skills and materials needed, but applies them to make products of greater market value.
-12 Inferences for inflation
-If free growth theory is correct, then net investment adds a superfluous cost to capital growth, while decreasing market value per unit through capital glut.
-Both effects would tend to contribute directly to capital inflation, and indirectly to inflation of consumer goods by adding depreciation costs.
-This consideration adds to the case for urgent attention to our argument, and for more study to confirm or refute it.
-13 Net output
-In Keynesian tradition, net output equals consumption plus net investment.
-We have disputed this reasoning in our argument that net investment adds nothing to capital growth, inhibits consumption, and thus tends to reduce the sum of both.
+We have proposed an exception in the innovator’s reserve; the innovator pays the going rates for skills and materials needed, and the going interest rate as the price of time, but applies them to make products of greater market value (See Appendix D).
+12 Inferences for interpretation of national accounts
+National accounts define the sum of consumption and investment as product.
+Common teaching identifies this sum, in its version net of depreciation, as net output in the sense of value added.
+We have disputed this reasoning in our argument that net investment realizes less than 8% of its cost in capital growth, inhibits consumption by its full cost, and thus tends to reduce the sum of both.
 We also critique it on another ground.
 Assume, for the sake of argument, that we are wrong, and that net investment realizes equal capital growth.
-Net output means value added.
 Appendix B will show an argument that to interpret value added as the sum of consumption and capital growth, and neither less nor more, would miss two components, one negative (human depreciation) and one positive (self-invested work), in value added to human capital.
@@
 Thus we have not sought to measure net output or its growth, and have treated capital growth alone.
+Common teaching also equates the sum of investment and consumption to expenditure.
+We agree that spending on new goods and services can find no other dispositions, and that investment and consumption can find no other sources, if we recognize imputed payment for voluntary consumption services as in parenting.
+By the argument in Appendix B noted above, consumption and market-value capital growth give the measurable components of true net output in the sense of value added.
+Then their sum, which might be thought of as market-value net product, would give a measurable proxy for net output, although not net output per se.
+It appears from the above that gross product and what we call market-value net product are both informative.
+The first gives expenditure, although not value added, and the second gives at least a first approximation of value added.
 13 Data sources
-All our data are drawn from the free online website World Inequality Database (WID).
+All our data are drawn from the free website World Inequality Database wid.world.com (WID).
 This source collates data from national accounts and tax data of 105 countries in constant currency units, and adjusts them where needed to conform to current standards of the System of National Accounts (SNA) published by the United Nations.
@@
 Tables and other displays of our findings for each country, and showing our methods of calculation, can be accessed at the web appendix (https://web-appendix.shinyapps.io/Sources_of_Capital_Growth).
-16 Displays
-The web appendix includes displays of $\theta$ for each year in each country over the report period.
+15 Screening out small denominators
+Our displays include findings for $\theta$ for each year in each country over the report period.
 These tend to show upward and downward spikes in values of $\theta$ in some years.
 Those spikes tend to be associated with small absolute values of denominators, in these cases $\Delta g(K)$, in those countries and years.
-Small denominators magnify errors in measurements of numerators.
+Small denominators, even if measured perfectly, magnify errors in measurements of numerators.
 Worse, when denominators are small, small mismeasurements in them might reverse the term in sign.
@@
 $\theta$ is plotted for each country unscreened and at each of the seven successive levels of screening.
-All displays in the main body of this paper apply a screen of 0.01.
+The screen chosen in displays in this paper is 0.01.
 The denominator whose absolute value is screened is capital acceleration $\Delta g(K)$ or capital growth rate $g(K)$ in all displays.
@@
 We dispute only the efficacy of net investment in realizing capital growth after replacement investment has made up for depreciation.
-Evidence supports the proposal of Modigliani (1954) that the young save to support retirement, and that retirees dissave to support consumption.
+Evidence supports the proposal of Modigliani (1954) that the young save to afford families and retirement, and that retirees dissave to pay for both as they arrive.
 [^17] Friedman (1957) follows a different argument to a similar conclusion.
@@
 17 Discussion and conclusions
-Capital glut is the condition warned against by West, Ricardo, Malthus and Harrod.
-It is loosely defined as oversupply of capital at the current state of technology.
-We take it to mean or include a state where more investment does not realize more capital value.
-Findings shown in our displays suggest that net saving raises the physical quantity of capital, say in number of shops, manufacturing plants or finished goods of similar design, without raising aggregate value of capital, and so contributes to capital glut by that understanding.
-These findings challenge the teachings that capital growth is effected by net saving enabled by consumption restraint, and that producer cost, including imputed interest as the opportunity cost of capital, converges to market realization.
-Evidence suggests that all capital growth is free, and consequently that market realization, in the presence of innovation, exceeds producer cost by the entirety of capital growth.
-Meanwhile the same evidence, which indicates that net saving adds no capital value, suggests a review of the teachings that consumption plus net saving gives net income, and that consumption plus net investment gives net output.
+With disclaimers noted, and with allowance for uncertainties of the thrift residue, we have argued that the innovator’s reserve and tech externalities pay for capital growth as a whole, and that net saving by some can realize equal capital growth only insofar as others dissave as much.
+Market appraisals raise present value at the warranted rate with no need for net saving at the overall scale.
+We adopted the Modigliani model, where young households save to afford families and retirement while older households dissave to pay for both as they arrive, as an example of this saving/dissaving balance.
+These findings challenge the teachings that capital growth at all scales is effected by net saving enabled by consumption restraint, and that producer cost, including imputed interest as the cost of time, converges to market realization.
+Evidence suggests that much or all capital growth at the collective scale is free, and consequently that market value, in the presence of innovation, exceeds producer cost by at least the majority of capital growth.
+Meanwhile the same evidence suggests a review of the teachings that consumption plus net saving gives net income, and that consumption plus net investment gives net output.
 Appendix B will also question the latter teachings.
@@
 It redeploys or repurposes existing labor skills, raw materials, and plant capacity, as well as existing finished goods, to achieve higher returns than available from the customary uses which determine their prices.
-The present value of yields from this advantage in return, or equivalently the innovator’s reserve, defines the non-random component in free growth.
+The present value of yields from this advantage in return, or equivalently the innovator’s reserve, plus tech externalities, defines the non-random component in free growth.
 Solow’s puzzle runs deep.
 How can thoughts be things?
-How can assessments of present value pay, at least in part, for their own realization?
+How can assessments of cash flows over the future pay for their realization in the present, even when plant of new design must come first?
 Science and philosophy push our minds into areas where they are not at home.
@@
 Our findings, with that allowance, are awkward for much of economics as generally taught.
-We encourage others to check our data and methods and inferences, and improve on them wherever they can.
+The data are compelling, even so, and our explanatory hypothesis in Section 1 fits common observation.
+We can see that producing more from the same inputs can spare a need for cutting back elsewhere.
+And we know that markets invest according to forecast future earnings, even when present earnings are nil.
+Free growth theory adds only that these well-understood effects, described in Piketty and Zucman (2014) and Kurz (2023), explain all or substantially all of capital growth.
 Appendix A.  Equivalence of saving and investment
 Keynes (1936) argued correctly that saving, if not invested in currently produced producer goods and services, adds nothing to capital, and thus nothing to output.
-He inferred that uninvested saving does not achieve actual saving, as it accomplishes no growth, and described it as "intended saving." True saving, by that reasoning, equalled investment.
+He inferred that uninvested saving has no value, as it accomplishes no growth, and described it as "intended saving." True saving, by that reasoning, equalled saving actually invested.
 We suggest that investors seek to maximize return, within risk tolerance, and will sometimes hold saving in cash or equivalents at zero return in recessions and depressions when positive returns cannot be found, and so when investing would tend to reduce capital and output rather than increase them.
@@
 It is only for these profitable times that we accept the equality of saving and investment, and only as a simplifying assumption.
-[^12] The weakness in the Keynes argument, if we are right, was his inference that saving does not accomplish its purpose unless invested to realize capital growth.
-Its purpose, we argue, is to maximize capital growth according to investment opportunities, and to maximize it by withholding investment when all investments promise loss.
+[^18] The weakness in the Keynes argument, if we are right, was his inference that net saving finds no utility, and hence no economic value, except as invested to realize capital growth.
+Its utility in a more general sense, we suggest, is to maximize capital growth according to investment opportunities, and to maximize it by withholding investment when all investments promise loss.
 Appendix B. Net output with human capital
@@
 Let human capital $H$, in that new model, stand as the last link in the value-added chain.
-Adapting the classic illustration of the value added principle, say that farms produce wheat, mills convert the wheat to flour, bakeries convert the flour into bread, and humans convert some of the bread, called invested consumption, into human capital.
+Adapting the classic illustration of the value added principle, that is, where farms produce wheat, mills convert the wheat to flour, and bakeries convert the flour into bread, say that humans convert some of the bread, called invested consumption, into human capital.
 The net outflow from this extended value-added chain is not all of consumption, but only pure consumption, meaning the part remaining after the part invested in human capital (invested consumption) is subtracted.
@@
 Yoram Ben-Porath (1967) reasoned that growth in human capital equals invested consumption plus self-invested work less human depreciation [^20] Equation 4 in Ben-Porath’s paper, summarizing his first three equations.
-His terms and notation differ from ours..
+His terms and notation differ from ours.
+He expressed his idea in Cobb-Douglas functions, whose values vary with scalars and exponents chosen, as opposed to our simpler algebra.
+We believe, however, that meanings agree..
 Let $C_{s},\ W_{s}$ and $D(H)$ show these flows respectively.
@@
 But it is obvious that whatever increases the productive power of labour creates an additional fund to make savings from, and enables capital to be enlarged not only without additional privation, but concurrently with an increase of personal consumption.
-This passage may be the first clear statement of what we call free growth theory.
-Mill’s use of the words "accelerate" and "concurrently" suggest that his path of reasoning was something like ours, and that he meant what we call free growth rather than the alternating phases of higher and lower saving rates pictured in thrift theory.
+This passage may be the first clear statement of what we call free growth theory. Mill’s use of the words "accelerate" and "concurrently" suggest that his path of reasoning was something like ours.
+Appendix D. Present value and cost
+We use the term cash flow to mean value transferred from an asset to its owner, in cash or in kind, less value transferred from the owner to the asset.
+Thus cash flow will generally be negative over any period of investment by owners in buying or building the asset, and then positive when investment ends and net yields begin.
+Present value of the asset, at any moment, is the sum of future cash flows expected from that moment on, each discounted at the market time discount rate (see Appendix F).
+Thus the present value of a differential amount of future cash flow expected at time $t$, shown as $dF(t)$, if discounted at the market rate $r_{m}$, finds a present value $dPV$ given by
+$dF(t)e^{-r_{m}t}=dPV\ .$ (D.1)
+Present value of all cash flows together, then, is
+$PV=\int^{\infty}F(t)e^{-r_{m}t}dt\ ,$ (D.2)
+where the implicit lower limit of integration is the present time zero.
+The actual investment needed to realize this present value will vary inversely with the productivity of the investment.
+Productivity of a given investment is measured in the rate of total return $r_{I}$ it generates, composed of growth rate $g(K)$ and cash flow rate.
+Consider a differential amount of investment $dI$ invested now, and yielded out wholly, including both principal and accumulated growth, in a differential amount of cash flow $dF$ at a single later time tt.212121dIdI and $dF$ need not represent the whole of investment or cash flow at times $t_{1}$ and $t_{2}$.
+It is always possible, in principle, to divide a given investment, including appreciation, into infinitesimal (differential) parts each yielded out at a single later moment.
+As no part of principal or accumulated growth from this particular investment was yielded during the interim $t$, by assumption, growth rate $g(K)$, over that interim, equalled the whole of rate of return $r_{I}$.
+Hence
+$dF(t)=dIe^{r_{I}t}\ ,\quad\text{or equivalently}\quad dI=dF(t)e^{-r_{I}t}\ .$ (D.3)
+Investment $I$ is the sum of these differential components $dI$. That is,
+$I=\int^{\infty}dI=\int^{\infty}dF(t)e^{-r_{I}t}dt\ .$ (D.4)
+Consider the case where $r_{I}$ and $r_{m}$ are equal. By Eqs. (D.1) through (D.4),
+$dPV=dFe^{-r_{m}t}=dFe^{-r_{I}t}=dI\quad\text{and}\quad PV=I\ ,\quad\text{if}\quad r_{I}=r_{m}\ .$ (D.5)
+Now relax that assumption, and allow $r_{m}$ and $r_{I}$ to differ.
+The innovator’s reserve is defined as present value less cost invested.
+Thus the component in each differential cash flow due to the innovator’s reserve is
+$dPV-dI=dF(t)e^{-r_{m}t}-dF(t)e^{-r_{I}t}=dF(t)(e^{-r_{m}t}-e^{-r_{I}t})\ .$
+The innovator’s reserve as a stock, which we show as $RSV$, is accordingly found at
+$RSV=PV-I=\int^{\infty}F(t)e^{-r_{m}t}dt-\int^{\infty}F(t)e^{-r_{I}t}dt=\int^{\infty}F(t)(e^{-r_{m}t}-e^{-r_{I}t})dt\ .$ (D.6)
+These equations confirm that present value equals investment needed when $r_{I}$ equals $r_{m}$, and exceeds investment needed when $r_{I}$ exceeds $r_{m}$.
+Appendix E. Analysis of the thrift residue
+The data shown in our displays find the thrift residue by three measures ranging from 7.71% to 5.59%.
+This residue gives the extent to which $s^{*}$ and $g(K)$ are causally related, but does not give the direction of causality.
+It shows the extent to which change in $s^{*}$ brought about a change in $g(K)$, or conversely, or both were caused primarily by a common third factor.
+In the last case, the third factor might bring about a direct effect on $s^{*}$ or $g(K)$ alone, at first, which might then cause a reaction by the other.
+Thrift theory predicts that a change in $s^{*}$ will bring an equal change in $g(K)$ before allowance for market noise, whether the change in $s^{*}$ was itself the primary cause, or whether it was a proximate cause driven by an exogenous shock, say war or epidemic or other disruption in supply or demand, as the primary one.
+Internal misjudgement can contribute to these setbacks by weakening defences, or can bring them about without need for significant outside causes.
+Some economists attribute at least some recessions to underspending on currently produced consumer and producer goods, made worse by stock market skittishness.
+This interpretation would fit thrift theory in picturing a shortfall in net investment as causing the drop in market value of capital.
+Another view pictures recessions as caused or facilitated by accumulated overinvestment through overestimation of growth prospects, and pictures recession at the end as a reality check.
+That scenario would reverse the direction of causality expected in thrift theory by putting the drop in $g(K)$ and income first, and by seeing the drop in $s^{*}$, the ratio of net saving to capital, as driven by a need to maintain consumption as income falls.
+The same chain of causality would repeat in the opposite direction as recovery from these crises brings capital value and income up as a direct effect, and so enables households to save against the next crisis.
+This issue was debated famously by Keynes and the Austrian school during the great depression, by Malthus, Say and Ricardo two centuries ago, and Smith and Mandeville before.
+They did not have the data reported in our displays.
+We suggest that the thrift theory explanation might have been urged less confidently in the light of information that net investment has accounted for less than 8% of capital growth.
+Other points can be added.
+Why should investors trade less efficiently in market equities than in the goods bought and sold by the underlying firms?
+This is the question we posed in our discussion of Tobin’s $q$ in Section 8.
+Appendix F. Clarifying cash flow and $r_{m}$
+The term "cash flow" is not much used in economics outside the specialty of finance, where it is taken as the flow discounted to find present value, or added to capital appreciation to find total return.
+We extend this concept to non-financial as well as financial assets by interpreting the services or utilities for which we buy a product as cash flows, or net withdrawals of value from assets to owners.
+[^22] We buy a non-financial asset, say a book or chair or car or house, for the services (taste satisfactions) we expect from it, each discounted at the risk-adjusted time discount rate we each apply.
+As each expected satisfaction eventuates, we deduct its value from the pool of remaining expected satisfactions, and so from remaining value to us of the asset.
+Although we do these calculations intuitively rather than formally, they fit the essential criteria for of finding present value as the sum of time-discounted projected cash flows.
+These cash flows are discounted to present value at the risk-adjusted time preference rate which the market applies to each product.
+Free growth theory predicts only at the universal scale, and is interested primarily in cash flows and time-discount rates at the scale of all products and all capital together.
+Thus we define $r_{m}$, except where stated otherwise, as the market time-discount rate, or equivalently the cap-weighted average of risk-adjusted rates, for the economy overall.
+Appendix D argued that $r_{m}$ also gives the market-wide rate of return in the absence of innovation, but that productivity gain through innovation can bring rate of return above the time discount rate.
+Thus $r_{m}$ can also be understood as the current standard rate of return before allowance for innovation.
+Note also that diminishing returns brought by capital glut should tend to drive $r_{m}$ down, as the measure of overall return before effects of current innovation, and so make further investment easier to justify by enhancing present value under Eq.
+(D.2).
+In this sense, capital glut may tend to be self-perpetuating.
 […]